### v1

# Abelian congruences and similarity in varieties with a weak difference term

Abstract.
This is the first of three papers motivated by the author’s desire to understand and explain “algebraically” one aspect of Dmitriy Zhuk’s proof of the CSP Dichotomy Theorem. In this paper we study abelian congruences in varieties having a weak difference term. Each class of the congruence supports an abelian group structure; if the congruence is minimal, each class supports the structure of a vector space over a division ring determined by the congruence. A construction due to J. Hagemann, C. Herrmann and R. Freese in the congruence modular setting extends to varieties with a weak difference term, and provides a “universal domain” for the abelian groups or vector spaces that arise from the classes of the congruence within a single class of the annihilator of the congruence. The construction also supports an extension of Freese’s similarity relation (between subdirectly irreducible algebras) from the congruence modular setting to varieties with a weak difference term.
Key words and phrases:
universal algebra, abelian congruence, weak difference term, similarity
2010 Mathematics Subject Classification:
08A05, 08B99 (Primary).
The support of the Natural Sciences and Engineering Research Council (NSERC) of Canada is gratefully acknowledged.
1. Introduction
Arguably the most important result in universal algebra in the last ten years is the positive resolution to the Constraint Satisfaction Problem (CSP) Dichotomy Conjecture, announced independently in 2017 by Andrei Bulatov [2] and Dmitriy Zhuk [36, 37]. One particular feature of Zhuk’s proof is his analysis of “rectangular critical” subdirect products $R\leq_{sd}{\mathbf{A}}_{1}\times\cdots\times{\mathbf{A}}_{n}$ of finite algebras ${\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{n}$ in certain locally finite idempotent Taylor varieties. Zhuk showed that, in the varieties he considers, such subdirect products generate derived relations, which he called “bridges,” between certain meet-irreducible congruences $\rho_{i}\in\operatorname{Con}({\mathbf{A}}_{i})$ which $R$ determines. Zhuk established a number of useful properties of his bridges and used them to tease out implicit linear equations in the CSP instances he considered.
Strikingly, Zhuk’s definition of “bridge” formally resembles some aspects of Hagemann and Herrmann’s formulation [11] from 1979 of the centrality relation on congruences of a single algebra. Moreover, the “there-exists-a-bridge” relation, under suitable restrictions, has an uncanny resemblance to the “similarity” relation of Freese [5] from 1983; this latter relation is between subdirectly irreducible algebras in a congruence modular variety, and can be viewed as an amplification of a certain correspondence involving the centrality relation. As well, Zhuk’s bridges arising in rectangular critical subdirect products in locally finite idempotent Taylor varieties resemble the “graphs of similarity” which Kearnes and Szendrei proved arise in rectangular critical relations of the form $R\leq{\mathbf{A}}^{n}$ in congruence modular varieties [20, Theorem 2.5(6)].
Our main goal in this paper and two companion papers [33, 32] is to establish precise connections between Zhuk’s bridges, centrality, and similarity, and then to apply these connections to the study of rectangular critical relations. In order to connect Zhuk’s bridges to similarity, we need to first generalize the similarity relation from its original setting (congruence modular varieties) to a broader class also containing all locally finite idempotent Taylor varieties. For this purpose, we focus on the class of varieties “having a weak difference term.” Here a “weak difference term” is an idempotent ternary term that specializes to a Maltsev operation whenever it is restricted to a block of an abelian congruence. Such terms are weak versions of terms, known as “difference terms,” that exist and play an important role in congruence modular varieties [9]. Weak difference terms arose organically in tame congruence theory in the 1980s, when Hobby and McKenzie [13] showed that every locally finite Taylor has such a term. The property of a (general) variety having a weak difference term was first explicitly identified in the 1990s in work of Lipparini [23], Kearnes [15], and Kearnes and Szendrei [19], who explored the extent to which centralizer properties in congruence modular varieties extend to varieties with a weak difference term. Recent work concerning varieties with a weak difference term is represented by [17, 35, 21, 16].
In this paper we show that Freese’s similarity relation generalizes nicely to varieties with a weak difference term. It turns out that in the abelian monolith case, Freese’s original definition of similarity, as well as the reformulation in [6], does not generalize in a straightforward way. Instead, we take as our starting point two equivalent (in the congruence modular setting) characterizations of similarity due to Freese – one based on the construction from Hagemann and Herrmann’s seminal work, the other a congruence lattice-theoretic formulation – and show that both characterizations generalize to, and are provably equivalent in, varieties with a weak difference term (Theorems 4.6 and 7.8). We also give a third equivalent characterization using a variant of Zhuk’s bridges, which we call “proper bridges” (Definition 7.3).
Along the way, we explore the abelian groups arising from the congruence classes of an abelian congruence $\theta$ in an algebra in a variety having a weak difference term. While the situation is not as well-behaved as in the congruence modular case (see Example 8.4), we discern some structure in Section 5. In particular, we show that the construction of Hagemann and Herrmann mentioned above provides a sort of “universal domain” for the abelian groups arising from the $\theta$-classes of ${\mathbf{A}}$ within each $(0:\theta)$-class, where $(0:\theta)$ is the largest congruence centralizing $\theta$ (Lemmas 5.2 and 5.4(3) and Corollary 5.8).
We also expend some energy explaining a folklore result: that each class of an abelian congruence $\theta$ (in an algebra ${\mathbf{A}}$ with a weak difference term) naturally supports the structure of a left vector space over a division ring $\mathbb{F}$, provided $\theta$ is minimal. Here $\mathbb{F}$ depends on ${\mathbf{A}}$ and $\theta$ but not on the particular $\theta$-class. In particular, we show that if two subdirectly irreducible algebras in a variety with a weak difference term have abelian monoliths and are similar, then the division rings determined by their monoliths are isomorphic.
In the companion paper [33], we will develop Zhuk’s theory of bridges (in a slightly more general setting than Zhuk did), and describe the precise relationships between the objects in Zhuk’s theory, the centrality relation, similarity, and proper bridges as defined here. In a third paper [32], we will study rectangular critical relations in locally finite Taylor varieties, extending the work of Zhuk [37] and connecting these relations to similarity (as was done by Kearnes and Szendrei [20] in the congruence modular setting). In particular, we will show explicitly how each such relation can be understood as encoding a family of “parallel” linear equations over a finite field.
Acknowledgment. We thank Keith Kearnes for several helpful conversations, including one which led us to a proof of an early version of Theorem 4.6.
2. Definitions
We assume that the reader is familiar with the fundamentals of universal algebra as given in [4], [26] or [1]. Our notation generally follows that in [26], [1] and [17]. We follow [17] and refer to a set of operation symbols with assigned arities as a signature.
If ${\mathbf{A}}$ is an algebra and $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\leq\beta$, then $\beta/\alpha$ denotes the congruence of ${\mathbf{A}}/\alpha$ corresponding to $\beta$ via the Correspondence Theorem ([26, Theorem 4.12] or [1, Theorem 3.6]). If in addition $\gamma,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\gamma\leq\delta$, then we write $(\alpha,\beta)\nearrow(\gamma,\delta)$ and say that $(\alpha,\beta)$ is perspective up to $(\gamma,\delta)$, or transposes up to $(\gamma,\delta)$, if $\beta\wedge\gamma=\alpha$ and $\beta\vee\gamma=\delta$. The notation $(\gamma,\delta)\searrow(\alpha,\beta)$ means the same thing. We write $\alpha\prec\beta$ and say that $\beta$ covers $\alpha$, or is an upper cover of $\alpha$, if $\alpha<\beta$ and there does not exist a congruence $\gamma$ satisfying $\alpha<\gamma<\beta$. A congruence $\alpha$ is minimal if it covers 00, and is completely meet-irreducible if $\alpha\neq 1$ and there exists $\alpha^{+}$ with $\alpha\prec\alpha^{+}$ and such that $\alpha<\beta\implies\alpha^{+}\leq\beta$ for all $\beta\in\operatorname{Con}({\mathbf{A}})$. A subset $T\subseteq A$ is a transversal for a congruence $\alpha$ if it contains exactly one element from each $\alpha$-class.
$[n]$ denotes $\{1,2,\ldots,n\}$. If ${\mathbf{A}}$ is an algebra, an $n$-ary term $t(x_{1},\ldots,x_{n})$ of ${\mathbf{A}}$ is idempotent if ${\mathbf{A}}$ satisfies the identity $t(x,\ldots,x)\approx x$, and is a Taylor term if it is idempotent, $n>1$, and for each $i\in[n]$, ${\mathbf{A}}$ satisfies an identity of the form
$t(u_{1},\ldots,u_{n})\approx t(v_{1},\ldots,v_{n})$
where each $u_{j}$ and $v_{k}$ is the variable $x$ or $y$, and $\{u_{i},v_{i}\}=\{x,y\}$. A particularly important example of a Taylor term is a weak near-unanimity term (WNU), which is an $n$-ary idempotent term $w(x_{1},\ldots,x_{n})$ with $n>1$ which satisfies
$w(y,x,x,\ldots,x)\approx w(x,y,x,\ldots,x)\approx w(x,x,y,\ldots,x)\approx\cdots\approx w(x,\ldots,x,y).$
Another important example of a Taylor term is a Maltsev term; this is a ternary term $p(x,y,z)$ satisfying the identities
(2.1) $p(x,x,y)\approx y\approx p(y,x,x).$
The identities (2.1) are called the Maltsev identities. Any ternary operation (whether a term or not) satisfying them is called a Maltsev operation.
Before defining “weak difference term,” we recall the ternary centralizer relation on congruences and the notion of abelian congruences. Given a non-empty set $A$, let $A^{2\times 2}$ denote the set of all $2\times 2$ matrics over $A$. If ${\mathbf{A}}$ is an algebra, let ${\mathbf{A}}^{2\times 2}$ denote the algebra with universe $A^{2\times 2}$ which is isomorphic to ${\mathbf{A}}^{4}$ via the bijection
$\begin{pmatrix}a_{1}&a_{3}\\ a_{2}&a_{4}\end{pmatrix}\mapsto(a_{1},a_{2},a_{3},a_{4}).$
Definition 2.1.
Suppose ${\mathbf{A}}$ is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    ${\mathbf{A}}(\theta,\varphi)$ is the subalgebra of ${\mathbf{A}}^{2\times 2}$ consisting of the matrices whose columns belong to $\theta$ and rows belong to $\varphi$.
  2. (2)
    $M(\theta,\varphi)$ is the subuniverse of ${\mathbf{A}}(\theta,\varphi)$ generated by the set
    $X(\theta,\varphi):=\left\{\begin{pmatrix}c&c\\ d&d\end{pmatrix}:(c,d)\in\theta\right\}\cup\left\{\begin{pmatrix}a&b\\ a&b\end{pmatrix}:(a,b)\in\varphi\right\}.$
    The matrices in $M(\theta,\varphi)$ are called $(\theta,\varphi)$-matrices.
Definition 2.2.
Suppose $\theta,\varphi,\delta\in\operatorname{Con}({\mathbf{A}})$. We say that $\varphi$ centralizes $\theta$ modulo $\delta$, and write $C(\varphi,\theta;\delta)$, if any of the following equivalent conditions holds:
  1. (1)
    For every matrix in $M(\varphi,\theta)$, if one row is in $\delta$, then so is the other row.
  2. (2)
    For every matrix in $M(\theta,\varphi)$, if one column is in $\delta$, then so is the other column.
  3. (3)
    For every term $t(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})$ and all pairs $(a_{i},b_{i})\in\varphi$ ($i\in[m]$) and $(c_{j},d_{j})\in\theta$ ($j\in[n]$),
    if $t(\mathbf{a},\mathbf{c})\stackrel{{\scriptstyle\delta}}{{\equiv}}t(\mathbf{a},\mathbf{d})$, then $t(\mathbf{b},\mathbf{c})\stackrel{{\scriptstyle\delta}}{{\equiv}}t(\mathbf{b},\mathbf{d})$.
More generally, if $S,T$ are tolerances of ${\mathbf{A}}$ (i.e., reflexive symmetric subuniverses of ${\mathbf{A}}^{2}$), then we can define ${\mathbf{A}}(S,T)$, $X(S,T)$, $M(S,T)$, and the relation $C(S,T;\delta)$ in the same way; see [17, §2.5].
Let $\theta,\delta$ be congruences of an algebra ${\mathbf{A}}$. The centralizer (or annihilator) of $\theta$ modulo $\delta$, denoted $(\delta:\theta)$, is the unique largest congruence $\varphi$ for which $C(\varphi,\theta;\delta)$ holds. We say that $\theta$ is abelian if $C(\theta,\theta;0)$ holds; equivalently, if $\theta\leq(0:\theta)$. More generally, if $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\theta$, then we say that $\theta$ is abelian modulo $\delta$ if any of the following equivalent conditions hold: (i) $\theta/\delta$ is an abelian congruence of ${\mathbf{A}}/\delta$; (ii) $C(\theta,\theta;\delta)$ holds; (iii) $\theta\leq(\delta:\theta)$. The following lemma expresses that centralizers behave nicely under certain quotients and surjective homomorphisms.
Lemma 2.3.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are algebras in a common signature.
  1. (1)
    If $\delta^{\prime},\delta,\theta\in\operatorname{Con}({\mathbf{A}})$ with $\delta^{\prime}\leq\delta\leq\theta$, then $(\delta/\delta^{\prime}:\theta/\delta^{\prime})=(\delta:\theta)/\delta^{\prime}$.
  2. (2)
    Suppose $f:{\mathbf{A}}\rightarrow{\mathbf{B}}$ is a surjective homomorphism and $\lambda,\mu\in\operatorname{Con}({\mathbf{B}})$ with $\lambda\leq\mu$. Then $f^{-1}((\lambda:\mu))=(f^{-1}(\lambda):f^{-1}(\mu))$.
Proof.
[…]
$\overline{f}((\delta:\theta)/\delta^{\prime})$ $=\overline{f}((\delta/\delta^{\prime}:\theta/\delta^{\prime}))$ by (1)
$=(\overline{f}(\delta/\delta^{\prime}):\overline{f}(\theta/\delta^{\prime}))$ as $\overline{f}$ is an isomorphism
$=(\lambda:\mu),$
which proves $(\delta:\theta)=f^{-1}((\lambda:\mu))$. ∎
Definition 2.4.
Let $\mathcal{V}$ be a variety, ${\mathbf{A}}\in\mathcal{V}$, and $d(x,y,z)$ a ternary term in the signature of $\mathcal{V}$.
  1. (1)
    $d$ is a weak difference term for ${\mathbf{A}}$ if $d$ is idempotent and for every pair $\delta,\theta$ of congruences with $\delta\leq\theta$ and $\theta/\delta$ abelian, we have
    (2.2) $d(a,a,b)\stackrel{{\scriptstyle\delta}}{{\equiv}}b\stackrel{{\scriptstyle\delta}}{{\equiv}}d(b,a,a)$ for all $(a,b)\in\theta$.
  2. (2)
    $d$ is a weak difference term for $\mathcal{V}$ if it is a weak difference term for every algebra in $\mathcal{V}$.
Note in particular that if $d$ is a weak difference term for ${\mathbf{A}}$ and $\theta$ is an abelian congruence, then setting $\delta=0$ in (2.2) gives that the restriction of $d$ to any $\theta$-class is a Maltsev operation on that class.
Because the definition of “weak difference term” is not given in terms of identities, it can happen that an algebra has a weak difference term, but does not belong to any variety having a weak difference term. Therefore, we will take care to distinguish “having a weak difference term” from the stronger “belonging to a variety with a weak difference term.”
At the level of varieties, the property of having a weak difference term can be characterized by the existence of idempotent terms satisfying a certain pattern of identities, i.e., by an idempotent Maltsev condition, as shown by Kearnes and Szendrei [19, Theorem 4.8]. (See Theorem 8.1 for a slight improvement.) In the context of locally finite varieties, the situation is simpler.
Theorem 2.5.
For a locally finite variety $\mathcal{V}$, the following are equivalent:
  1. (1)
    $\mathcal{V}$ has a Taylor term.
  2. (2)
    $\mathcal{V}$ has a WNU term.
  3. (3)
    $\mathcal{V}$ has a weak difference term.
Proof.
This follows by combining [13, Theorem 9.6], [30, Corollary 5.3], [24, Theorem 2.2], and [19, Theorem 4.8]. ∎
Lemma 5.2 and Corollary 5.3 mention difference terms; these are weak difference terms which moreover satisfy the identity $d(x,x,y)\approx y$. By an old theorem of Gumm [9], every congruence modular variety has a difference term.
The proof of Proposition 3.4 will make use of the two-term condition. This is the requirement on a congruence $\theta\in\operatorname{Con}({\mathbf{A}})$ that for all $(\theta,\theta)$-matrices
$\begin{pmatrix}a_{1}&a_{3}\\ a_{2}&a_{4}\end{pmatrix},\begin{pmatrix}b_{1}&b_{3}\\ b_{2}&b_{4}\end{pmatrix}\in M(\theta,\theta),$
if $a_{i}=b_{i}$ for $i=1,2,3$, then $a_{4}=b_{4}$. It is easy an easy exercise to show that if $\theta$ satisfies the two-term condition, then $\theta$ is abelian. The converse implication is also true if ${\mathbf{A}}$ belongs to a variety having a weak difference term, by [23, Theorem 3.1] or [19, Corollary 4.5]. (See Lemma 3.3 for a slight improvement.)
3. Tools
In this section we gather some technical results which will be used in later sections. We start with results about algebras having a weak difference term.
The first result is an easily proved generalization of the well-known fact (see e.g. [1, Theorem 4.65(1)]) that if an algebra ${\mathbf{A}}$ has a Maltsev term, then every reflexive subuniverse of ${\mathbf{A}}^{2}$ is a congruence of ${\mathbf{A}}$. (A proof can be easily extracted from the proof of [17, Corollary 6.22].)
Lemma 3.1.
Suppose ${\mathbf{A}}$ is an algebra and $\rho$ is a reflexive subuniverse of ${\mathbf{A}}^{2}$. Suppose ${\mathbf{A}}$ has a ternary term $d(x,y,z)$ such that for all $(a,b)\in\rho$ we have $d(a,a,b)=b$ and $d(a,b,b)=a$. Then $\rho\in\operatorname{Con}({\mathbf{A}})$.
Corollary 3.2.
Suppose ${\mathbf{A}}$ is an algebra with a weak difference term, $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian, and $\theta$ is minimal. Then for all $(a,b),(a^{\prime},b^{\prime})\in\theta$, if $a\neq b$ then there exists a unary polynomial $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ satisfying $f(a)=a^{\prime}$ and $f(b)=b^{\prime}$.
Proof.
Let $\rho=\{(f(a),f(b))\,:\,\text{$f\in\operatorname{Pol}_{1}({\mathbf{A}})$}\}$. Then $\rho$ is a reflexive subuniverse of ${\mathbf{A}}^{2}$ satisfying $0_{A}\neq\rho\subseteq\theta$. Hence $\rho\in\operatorname{Con}({\mathbf{A}})$ by Lemma 3.1, and thus $\rho=\theta$ as $\theta$ is minimal. ∎
Lemma 3.3.
Suppose the algebra ${\mathbf{A}}$ has a weak difference term and $\theta\in\operatorname{Con}({\mathbf{A}})$. If $\theta$ is abelian, then $\theta$ satisfies the two-term condition.
Proof.
See Lipparini’s proof of [23, Theorem 3.1(a)]. His Theorem 3.1(a) assumes that ${\mathbf{A}}$ belongs to a variety with a difference term, but in fact all the argument requires is that ${\mathbf{A}}$ itself have a weak difference term. ∎
As a consequence, a well-known property of difference terms in congruence modular varieties due to Gumm [9] partially extends to algebras with weak difference terms.
Proposition 3.4.
Suppose ${\mathbf{A}}$ is an algebra with a weak difference term $d(x,y,z)$, and $\theta\in\operatorname{Con}({\mathbf{A}})$ is an abelian congruence of ${\mathbf{A}}$. Then for every polynomial operation $t(x_{1},\ldots,x_{n})$ and all $a_{i},b_{i},c_{i}\in A$ $(1\leq i\leq n)$ with $a_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}b_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}c_{i}$,
$d(t(\mathbf{a}),t(\mathbf{b}),t(\mathbf{c}))=t(d(a_{1},b_{1},c_{1}),\ldots,d(a_{n},b_{n},c_{n})).$
Proof.
Obviously $t(\mathbf{a}),t(\mathbf{b}),t(\mathbf{c})$ belong to a common $\theta$-class. Define polynomial operations $r(\mathbf{x},\mathbf{y})=d(t(\mathbf{a}),t(\mathbf{x}),t(\mathbf{y}))$ and $s(\mathbf{x},\mathbf{y})=t(d(a_{1},x_{1},y_{1}),\ldots,d(a_{n},x_{n},y_{n}))$. Then
$\begin{pmatrix}r(\mathbf{a},\mathbf{b})&r(\mathbf{a},\mathbf{c})\\ r(\mathbf{b},\mathbf{b})&r(\mathbf{b},\mathbf{c})\end{pmatrix}=\begin{pmatrix}t(\mathbf{b})&t(\mathbf{c})\\ t(\mathbf{a})&d(t(\mathbf{a}),t(\mathbf{b}),t(\mathbf{c}))\end{pmatrix}=:M_{1}$
while
$\begin{pmatrix}s(\mathbf{a},\mathbf{b})&s(\mathbf{a},\mathbf{c})\\ s(\mathbf{b},\mathbf{b})&s(\mathbf{b},\mathbf{c})\end{pmatrix}=\begin{pmatrix}t(\mathbf{b})&t(\mathbf{c})\\ t(\mathbf{a})&t(d(a_{1},b_{1},c_{1}),\ldots,d(a_{n},b_{n},c_{n}))\end{pmatrix}=:M_{2}.$
Obviously $M_{1},M_{2}$ are $(\theta,\theta)$-matrices, and since they agree at all positions except one, they form an instance of the two-term condition for $\theta$. As $\theta$ is abelian and ${\mathbf{A}}$ has a weak difference term, $\theta$ satisfies the two-term condition by Lemma 3.3. Thus the lower-right entries of $M_{1}$ and $M_{2}$ are equal, as desired. ∎
The next result is an immediate consequence of the definition of weak difference term and a classical theorem of Gumm [8] and Herrmann [12].
Lemma 3.5.
Suppose ${\mathbf{A}}$ is an algebra having a weak difference term $d(x,y,z)$, and $\theta$ is an abelian congruence of ${\mathbf{A}}$. Then for all $e\in A$, the structure $(e/\theta,+,e)$ is an abelian group where $x+y:=d(x,e,y)$ for $x,y\in e/\theta$. Moreover, in this group we have $-x=d(e,x,e)$ and $d(x,y,z)=x-y+z$ for all $x,y,z\in e/\theta$.
Proof.
Let $B=e/\theta$. Then $(B,d{\upharpoonright}_{B^{3}})$ is an abelian algebra with Maltsev operation $d{\upharpoonright}_{B^{3}}$. Now apply Gumm [8, Theorem 4.7] or Herrmann’s main theorem from [12] (or see [31, Corollary 2] for a short proof). ∎
While an algebra may have many different weak difference terms, they all agree whenever restricted to a class of an abelian congruence.
Lemma 3.6.
Suppose ${\mathbf{A}}$ is an algebra and $\theta$ is an abelian congruence. If $d(x,y,z)$ and $t(x,y,z)$ are two weak difference terms for ${\mathbf{A}}$, then for every $\theta$-class $B$ we have that $d(a,b,c)=t(a,b,c)$ whenever $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b\stackrel{{\scriptstyle\theta}}{{\equiv}}c$.
Proof.
Again this is a well-known consequence of $(B,d{\upharpoonright}_{B^{3}})$ being an abelian Maltsev operation. For $a,b,c\in B$, Proposition 3.4 yields
$d(t(a,b,b),t(b,b,b),t(b,b,c))=t(d(a,b,b),d(b,b,b),d(b,b,c))$
which simplifies to $d(a,b,c)=t(a,b,c)$. ∎
Definition 3.7.
Suppose ${\mathbf{A}}$ has a weak difference term, $\theta$ is an abelian congruence, and $e\in A$. We let $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ denote the abelian group with universe $e/\theta$, zero element $e$, and group operation $x+y:=d(x,e,y)$, where $d$ is a weak difference term.
The operation $+$ in the previous definition does not depend on the choice of the weak difference term, by Lemma 3.6. The notation $M(\theta,e)$ has been used in [6] and elsewhere for $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. In case ${\mathbf{A}}$ is finite and $\theta$ is a minimal congruence, more can be said about the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. The following result is possibly folklore; we include it here for completeness.
Theorem 3.8.
Suppose ${\mathbf{A}}$ is a finite algebra with a weak difference term, $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian, and $0\prec\theta$. There exists a prime $p$ such that for all $e\in A$, $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ is an elementary abelian $p$-group. Hence every $\theta$-class has cardinality $p^{k}$ for some $k\geq 0$ (where $k$ can depend on the $\theta$-class).
Proof sketch.
This was proved by Freese [5, §4] in the congruence modular case, and essentially the same proof works here. We recast his argument using tools from a recent paper of Mayr and Szendrei [25]. (See also the arguments in Section 6, where a generalization is proved.)
First note that we can assume that ${\mathbf{A}}$ has no nullary operations (by replacing each nullary operation with a constant unary operation). Let ${\mathbf{I}}={\mathbf{A}}/\theta$, and let $\nu_{\theta}:{\mathbf{A}}\rightarrow{\mathbf{I}}$ be the natural map. The pair $({\mathbf{A}},\nu_{\theta})$ is called an algebra over ${\mathbf{I}}$ in [25, Definition 2.1]. By an old construction of Novotný [28], one can obtain a multisorted algebra ${\mathbf{M}}=\mathfrak{M}({\mathbf{A}},\nu_{\theta})$ whose sort domains are the $\theta$-classes of ${\mathbf{A}}$; then applying another old construction of Gardner [7], one can convert ${\mathbf{M}}$ to a one-sorted algebra whose universe is the Cartesian product of the $\theta$-classes of ${\mathbf{A}}$. Call this final algebra $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$. For precise definitions, see [25, Propositions 2.2 and 2.5]. Because $\theta$ is minimal and abelian, it follows from [25, Corollary 3.5 and Theorem 3.6] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple and abelian. And because ${\mathbf{A}}$ has a term $d(x,y,z)$ whose restriction to each $\theta$-class is a Maltsev operation, it follows from the definitions and [25, Remark 2.8] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ has a Maltsev term. In fact, writing the universe of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ as $B_{1}\times\cdots\times B_{k}$ where $B_{1},\ldots,B_{k}$ is an enumeration of the $\theta$-classes, then the Maltsev operation $m(x,y,z)$ of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is given coordinatewise by $d$.
Since $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple, abelian, and has a Maltsev term, it follows that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is polynomially equivalent to a (finite) simple module by a theorem of Gumm [8] and Herrmann [12]. If we fix a $k$-tuple $\mathbf{e}=(e_{1},\ldots,e_{k})\in B_{1}\times\cdots\times B_{k}$, then we can assume that $\mathbf{e}$ is the zero element of this module and addition is given by $x+y=m(x,\mathbf{e},y)$. By Schur’s Lemma and finiteness, the endomorphism ring of this module is a finite field $\mathbb{F}$, say of cardinality $p^{m}$. It follows that the abelian group $(B_{1}\times\cdots\times B_{k},+,\mathbf{e})$ inherits the structure of a vector space over $\mathbb{F}$, which forces $px=\mathbf{e}$ for all $x\in B_{1}\times\cdots\times B_{k}$. Interpreting this equation coordinatewise gives that each group $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ satisfies $px=e_{i}$ for each $i=1,\ldots,k$ as required. ∎
Note that if ${\mathbf{A}},\theta,p$ are as in the statement of Theorem 3.8, then every $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ can be viewed as a vector space over $\mathbb{Z}_{p}$. In fact, Freese’s proof of Theorem 3.8 (in the congruence modular case) actually showed that each $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ carries the structure of a vector space over $\mathbb{F}$. Freese proved something similar when ${\mathbf{A}}$ is infinite: there is a division ring $\mathbb{F}$ determined by ${\mathbf{A}}$ and the abelian minimal congruence $\theta$ such that every $\theta$-class inherits the structure of a left $\mathbb{F}$-vector space. We will show in Section 6 that Freese’s proof carries through assuming only that ${\mathbf{A}}$ has a weak difference term.
The next lemma will be needed in Section 7. Its proof, in the congruence modular case, essentially can be extracted from the proof of [6, Theorem 10.11].
Lemma 3.9.
Suppose ${\mathbf{A}}$ is an with a weak difference term, $\theta\in\operatorname{Con}({\mathbf{A}})$, and $\theta$ is abelian. Let $D_{1},D_{2}$ be subuniverses of ${\mathbf{A}}$ which are transversals for $\theta$. Then there exists an automorphism $\sigma$ of ${\mathbf{A}}$ satisfying $\sigma(D_{1})=D_{2}$ and $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}x$ for all $x\in A$.
Proof.
For $i=1,2$ let $\pi_{i}:A\rightarrow D_{i}$ be the unique map given by $\pi_{i}(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}x$ for all $x\in A$, and note that $\pi_{1}$ and $\pi_{2}$ are retractions of ${\mathbf{A}}$ with kernel $\theta$. Let $d(x,y,z)$ be a weak difference term for $\mathcal{V}$ and define $\sigma:A\rightarrow A$ and $\tau:A\rightarrow A$ by
$\sigma(x)$ $=d(x,\pi_{1}(x),\pi_{2}(x))$
$\tau(x)$ $=d(x,\pi_{2}(x),\pi_{1}(x)).$
Observe first that $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}\tau(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}d(x,x,x)=x$ for all $x\in A$. Fix a $\theta$-class $B$; by construction, there exist $b,c\in B$ such that $\pi_{1}(x)=b$ and $\pi_{2}(x)=c$ for all $x\in B$. Pick an abelian group operation $+$ on $B$ such that $d(x,y,z)=x-y+z$ for all $x,y,z\in B$. Then $\sigma(x)=x-b+c$ and $\tau(x)=x-c+b$ for all $x\in B$, and since this is true for every $\theta$-class, we get $\sigma\circ\tau=\tau\circ\sigma=\mathrm{id}_{A}$, proving $\sigma$ is a bijection. Our aim is to prove that $\sigma$ is the desired automorphism.
Let $f$ be an $n$-ary basic operation. Let $a_{1},\ldots,a_{n}\in A$ and put $a=f(a_{1},\ldots,a_{n})$; we need to prove $\sigma(a)=f(\sigma(a_{1}),\ldots,\sigma(a_{n}))$. Let $b_{i}=\pi_{1}(a_{i})$ and $c_{i}=\pi_{2}(a_{i})$, and note that $a_{i},b_{i},c_{i}$ belong to a common $\theta$-class $B_{i}$ for each $i=1,\ldots,n$. Let $b=f(b_{1},\ldots,b_{n})$ and $c=f(c_{1},\ldots,c_{n})$, and note that $b\in D_{1}$ and $c\in D_{2}$ (since $D_{1},D_{2}\leq{\mathbf{A}}$), so $\pi_{1}(a)=b$ and $\pi_{2}(a)=c$. Thus what must be shown is
$d(f(a_{1},\ldots,a_{n}),f(b_{1},\ldots,b_{n}),f(c_{1}\ldots,c_{n}))=f(d(a_{1},b_{1},c_{1}),\ldots,d(a_{n},b_{n},c_{n}))$
given that $a_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}b_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}c_{i}$ for all $i=1,\ldots,n$. This follows from Proposition 3.4. Hence $\sigma$ is an automorphism of ${\mathbf{A}}$.
Finally, suppose $a\in D_{1}$. Then $\pi_{1}(a)=a$, so $\sigma(a)=d(a,a,\pi_{2}(a))=\pi_{2}(a)\in D_{2}$. This proves $\sigma(D_{1})\subseteq D_{2}$. A similar argument shows $\tau(D_{2})\subseteq D_{1}$, so $\sigma(D_{1})=D_{2}$. ∎
Next we give some results concerning varieties with a weak difference term. The first three Propositions are extracted from [17, 18].
Proposition 3.10.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    If $C(\theta,\theta;\delta)$, then $\theta\vee\delta=\delta\circ\theta\circ\delta$ and $(\theta\vee\delta)/\delta$ is abelian.
  2. (2)
    If $\theta$ is abelian, then $C(\theta,\theta;\delta)$.
Proof.
These are special cases of Lemmas 6.7 and 6.8 respectively from [17]. ∎
Proposition 3.11.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\leq\beta$. If $\beta/\alpha$ is abelian, then the interval $I[\alpha,\beta]$ is a modular lattice of permuting equivalence relations.
Proof.
This is a special case of [17, Theorem 6.16]. ∎
Proposition 3.12.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\sigma,\tau,\delta,\varepsilon\in\operatorname{Con}({\mathbf{A}})$ with $\sigma\leq\tau$ and $\delta\leq\varepsilon$ and $(\sigma,\tau)\nearrow(\delta,\varepsilon)$.
  1. (1)
    $\tau/\sigma$ is abelian if and only if $\varepsilon/\delta$ is abelian.
  2. (2)
    If $\tau/\sigma$ is abelian, then the map $J:I[\sigma,\tau]\rightarrow I[\delta,\varepsilon]$ given by $J(x)=x\vee\delta$ is surjective.
Proof.
(1) follows from [17, Theorem 3.27] as corrected in [18]. (2) is a special case of [17, Theorem 6.24]. ∎
Our last result in this section generalizes [14, Lemma 2.6].
Proposition 3.13.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\delta,\varepsilon,\sigma,\tau\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\varepsilon$ and $\sigma\leq\tau$ where $(\sigma,\tau)\nearrow(\delta,\varepsilon)$.
  1. (1)
    If $\varepsilon/\delta$ is abelian, then $\varepsilon=\delta\circ\tau\circ\delta$ and $(\delta:\varepsilon)=(\sigma:\tau)$.
  2. (2)
    If $\delta\prec\varepsilon$, then $(\delta:\varepsilon)=(\sigma:\tau)$.
Proof.
(1) We have $C(\varepsilon,\varepsilon;\delta)$ by hypothesis, and hence $C(\tau,\tau;\delta)$ by monotonicity. Then $\varepsilon=\tau\vee\delta=\delta\circ\tau\circ\delta$ by Proposition 3.10(1).
$\delta\wedge\tau=\sigma$ implies $C(\delta,\tau;\sigma)$ by e.g. [17, Theorem 2.19(8)], proving $(\sigma:\tau)\geq\delta$. Thus to prove $(\delta:\varepsilon)=(\sigma:\tau)$, it is enough to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\sigma)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. We have $C(\alpha,\tau;\sigma)\iff C(\alpha,\tau;\delta)$ by [17, Theorem 2.19(4)], since $\tau\wedge\sigma=\tau\wedge\delta$. Thus we only need to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\delta)$. The forward implication follows immediately by monotonicity. Conversely, suppose $C(\alpha,\tau;\delta)$. Then $C(\alpha,\delta\circ\tau\circ\delta;\delta)$ by [17, Theorem 2.19(3)], i.e., $C(\alpha,\varepsilon;\delta)$.
(2) If $\varepsilon/\delta$ is abelian, then we can apply item (1). Assume $\varepsilon/\delta$ is nonabelian. Then $\tau/\sigma$ is also nonabelian by Proposition 3.12(1). As in the proof of (1), we only need to prove $C(\alpha,\tau;\delta)\implies C(\alpha,\varepsilon;\delta)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. Assume $C(\alpha,\tau;\delta)$. If $\alpha\geq\varepsilon$, then $C(\tau,\tau;\delta)$ by monotonicity, which contradicts nonabelianness of $\tau/\sigma$. Thus $\alpha\ngeq\varepsilon$. Since $\alpha\geq\delta$ and $\delta\prec\varepsilon$, we get $\alpha\wedge\varepsilon=\delta$ and hence $C(\alpha,\varepsilon;\delta)$ as required. ∎
4. Generalizing an old construction
In this section we revisit an old construction of Hagemann and Herrmann [11] which they used to characterize the centrality relation in congruence modular varieties. Freese [5] adapted the construction to give a clean characterization of his similarity relation in congruence modular varieties. We will show that the construction adapts nicely to varieties with a weak difference term, and we will use it in Section 7 to define similarity in that setting.
In the following definition, we follow the notation of [6].
Definition 4.1.
Suppose ${\mathbf{A}}$ is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$ with $\theta\leq\alpha$.
  1. (1)
    ${\mathbf{A}}(\theta)$ denotes $\theta$ viewed as a subalgebra of ${\mathbf{A}}^{2}$.
  2. (2)
    $\Delta_{\theta,\varphi}$ denotes the congruence of ${\mathbf{A}}(\theta)$ generated by $\{((a,a),(b,b))\,:\,\text{$(a,b)\in\varphi$}\}$.
The next notations depart from those in [6].
Definition 4.2.
Let ${\mathbf{A}},\theta,\varphi$ be as in Definition 4.1.
  1. (1)
    $\operatorname{pr}_{1},\operatorname{pr}_{2}:{\mathbf{A}}(\theta)\rightarrow{\mathbf{A}}$ denote the two projection homomorphisms.
  2. (2)
    $\eta_{i}=\ker(\operatorname{pr}_{i})\in\operatorname{Con}({\mathbf{A}}(\theta))$ for $i=1,2$.
  3. (3)
    If $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\theta$, then $\overline{\alpha}$ denotes
    $\overline{\alpha}=\{((a_{1},a_{2}),(b_{1},b_{2}))\in A(\theta)\times A(\theta)\,:\,\text{$(a_{1},b_{1})\in\alpha$}\}.$
Note that in the context of Definition 4.2 we have $\overline{\alpha}\in\operatorname{Con}({\mathbf{A}}(\theta))$ and $\overline{\alpha}$ is also given by $\overline{\alpha}=\{((a_{1},a_{2}),(b_{1},b_{2}))\in A(\theta)\times A(\theta)\,:\,\text{$(a_{2},b_{2})\in\alpha$}\}$. Also observe that for each $i=1,2$, $\operatorname{Con}({\mathbf{A}})$ is naturally isomorphic to the interval $I[\eta_{i},1]$ in $\operatorname{Con}({\mathbf{A}}(\theta))$ via $\operatorname{pr}_{i}^{-1}$, and $\alpha$ maps to $\overline{\alpha}$ under either isomorphism; see Figure 1.
It will be useful to view $\Delta_{\theta,\varphi}$ as a set of $(\theta,\varphi)$-matrices. Conversely, it will be useful to view $M(\theta,\varphi)$ as a tolerance of ${\mathbf{A}}(\theta)$.
Definition 4.3.
Suppose ${\mathbf{A}}$ is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    Let $\Delta_{\mathsf{h}}(\theta,\varphi)$ denote the subuniverse of ${\mathbf{A}}(\theta,\varphi)$ “corresponding to” $\Delta_{\theta,\varphi}$:
    $\Delta_{\mathsf{h}}(\theta,\varphi)=\left\{\begin{pmatrix}a&c\\ b&d\end{pmatrix}:((a,b),(c,d))\in\Delta_{\theta,\varphi}\right\}.$
  2. (2)
    Let $M_{\theta,\varphi}$ denote the subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$ “corresponding to” $M(\theta,\varphi)$:
    $M_{\theta,\varphi}=\left\{((a,b),(c,d)):\begin{pmatrix}a&c\\ b&d\end{pmatrix}\in M(\theta,\varphi)\right\}.$
Lemma 4.4.
Suppose ${\mathbf{A}}$ is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    $\Delta_{\theta,\varphi}$ is the transitive closure of of $M_{\theta,\varphi}$. Thus $\Delta_{\mathsf{h}}(\theta,\varphi)$ is the “horizontal transitive closure of $M(\theta,\varphi)$.”
  2. (2)
    $\Delta_{\mathsf{h}}(\theta,\varphi)$ is “vertically symmetric.” I.e.,
    $\mbox{If~{}}\begin{pmatrix}a&c\\ b&d\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi),~{}\mbox{then}~{}\begin{pmatrix}b&d\\ a&c\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi).\$
  3. (3)
    If ${\mathbf{A}}$ has a weak difference term and $\varphi$ is abelian, then $\Delta_{\mathsf{h}}(\theta,\varphi)=M(\theta,\varphi)$.
  4. (4)
    If ${\mathbf{A}}$ has a weak difference term and $\theta$ is abelian, then $\Delta_{\mathsf{h}}(\theta,\varphi)$ is “vertically transitive.” I.e.,
    If $\begin{pmatrix}a&c\\ b&d\end{pmatrix},\begin{pmatrix}b&d\\ r&s\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi),~{}\mbox{then}~{}\begin{pmatrix}a&c\\ r&s\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi).$
Proof.
(1) $M_{\theta,\varphi}$ is a reflexive, symmetric subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$.
(2) This follows from (1) and the fact that the generating set $X(\theta,\varphi)$ for $M(\theta,\varphi)$ is vertically symmetric.
(3) Let $\rho=M_{\theta,\varphi}$; then $\rho$ is a reflexive subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$. If $\varphi$ is abelian and ${\mathbf{A}}$ has a weak difference term $d$, then $d$ interpreted coordinatewise in ${\mathbf{A}}(\theta)$ satisfies the hypothesis of Lemma 3.1 for $\rho$. Thus $\rho$ is transitive by Lemma 3.1, implying $\Delta_{\theta,\varphi}=M_{\theta,\varphi}$ and thus $\Delta_{\mathsf{h}}(\theta,\varphi)=M(\theta,\varphi)$.
(4) We can consider $\Delta_{\mathsf{h}}(\theta,\varphi)$ as a binary relation on the set of its rows. As such, it is a reflexive subuniverse of ${\mathbf{A}}(\varphi)\times{\mathbf{A}}(\varphi)$, and if $\theta$ is abelian, a weak difference term for ${\mathbf{A}}$, interpreted coordinatewise in ${\mathbf{A}}(\varphi)$, will satisfy the hypothesis of Lemma 3.1 for this relation. ∎
Remark 4.5.
Assuming the hypotheses of Lemma 4.4(4), $\Delta_{\mathsf{h}}(\theta,\varphi)$ is an example of what A. Moorhead calls a “2-dimensional congruence” [27]. In fact, $\Delta_{\mathsf{h}}(\theta,\varphi)$ is the smallest 2-dimensional congruence generated by $M(\theta,\varphi)$, which Moorhead denotes by $\Delta(\theta,\varphi)$.
The next theorem has its roots in [11, Proposition 1.1].
Theorem 4.6.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is a nonzero abelian congruence of ${\mathbf{A}}$, and $\alpha=(0:\theta)$. Let $\Delta=\Delta_{\theta,\alpha}$, let $\eta_{1},\eta_{2}$ be the kernels of the two projection maps $\operatorname{pr}_{1},\operatorname{pr}_{2}:{\mathbf{A}}(\theta)\rightarrow{\mathbf{A}}$, and let $\varepsilon=\overline{\theta}\wedge\Delta$. Then:
  1. (1)
    For any $a\in A$, $(a,a)/\Delta=\{(b,b)\,:\,\text{$(a,b)\in\alpha$}\}$.
  2. (2)
    $\Delta<\overline{\alpha}$ and $(\Delta:\overline{\alpha})=\overline{\alpha}$.
  3. (3)
    $\overline{\theta}$ is abelian, and the interval $I[0,\overline{\theta}]$ in $\operatorname{Con}({\mathbf{A}}(\theta))$ is a modular lattice consisting of permuting congruences.
  4. (4)
    $\{0,\eta_{1},\eta_{2},\varepsilon,\overline{\theta}\}$ is a sublattice of $I[0,\overline{\theta}]$ isomorphic to ${\mathbf{M}}_{3}$.
  5. (5)
    Each of $(\varepsilon,\overline{\theta}),(0,\eta_{1}),(0,\eta_{2})$ transposes up to $(\Delta,\overline{\alpha})$.
  6. (6)
    If $\theta$ is minimal, i.e., $0\prec\theta$, then the interval $I[0,\overline{\theta}]$ has height $2$, $\Delta\prec\overline{\alpha}$, and $\Delta$ is completely meet-irreducible.
Proof.
(1) follows from the fact that $\alpha$ centralizes $\theta$. In $\operatorname{Con}({\mathbf{A}}(\theta))$ we have $\eta_{1}\wedge\eta_{2}=0$ and $\eta_{i}<\overline{\theta}$ for $i=1,2$. Clearly $\Delta\leq\overline{\alpha}$. Item (1) and $\theta\neq 0$ imply $\overline{\theta}\nleq\Delta$ and hence $\Delta<\overline{\alpha}$ and $\varepsilon<\overline{\theta}$. If $\varphi$ is any congruence of ${\mathbf{A}}$ satisfying $\theta\leq\varphi\leq\alpha$, then we can show that $(\Delta\wedge\overline{\varphi})\vee\eta_{1}=(\Delta\wedge\overline{\varphi})\vee\eta_{2}=\overline{\varphi}$. Indeed, if $((a_{1},a_{2}),(b_{1},b_{2}))\in\overline{\varphi}$ then
$(a_{1},a_{2})\stackrel{{\scriptstyle\eta_{i}}}{{\equiv}}(a_{i},a_{i})\stackrel{{\scriptstyle\Delta\wedge\overline{\varphi}}}{{\equiv}}(b_{i},b_{i})\stackrel{{\scriptstyle\eta_{i}}}{{\equiv}}(b_{1},b_{2}).$
In particular, this proves $\Delta\vee\eta_{i}=\overline{\alpha}$ and $\varepsilon\vee\eta_{i}=\overline{\theta}$ for $i=1,2$; hence $\varepsilon\neq 0$. The first equality also proves $\Delta\vee\overline{\theta}=\overline{\alpha}$. We also get $\Delta\wedge\eta_{1}=\Delta\wedge\eta_{2}=0$ by (1); hence $\varepsilon\wedge\eta_{i}=0$ for $i=1,2$. And we can easily verify $\eta_{1}\vee\eta_{2}=\overline{\theta}$. This proves (4) and (5). See Figure 1 showing the placement of these congruences in $\operatorname{Con}({\mathbf{A}}(\theta))$.
[…]
Figure 1. $\operatorname{Con}({\mathbf{A}})$ and $\operatorname{Con}({\mathbf{A}}(\theta))$
Since $\theta$ is abelian, we have $C(\overline{\theta},\overline{\theta};\eta_{i})$ for $i=1,2$. As $\eta_{1}\wedge\eta_{2}=0$, we get $C(\overline{\theta},\overline{\theta};0)$, i.e., $\overline{\theta}$ is abelian. Hence by Proposition 3.11, the interval $I[0,\overline{\theta}]$ is a modular lattice of permuting equivalence relations. This proves (3).
As $\overline{\theta}$ is abelian, we use Proposition 3.10(2) to get $C(\overline{\theta},\overline{\theta};\varepsilon)$, i.e., $\overline{\theta}/\varepsilon$ is abelian. As $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get that $\overline{\alpha}/\Delta$ is abelian by Proposition 3.12(1). Since also $\overline{\theta}/\eta_{1}$ is abelian and $(\eta_{1},\overline{\theta})\searrow(0,\eta_{2})\nearrow(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get $(\eta_{1}:\overline{\theta})=(0:\eta_{2})=(\varepsilon:\overline{\theta})=(\Delta:\overline{\alpha})$ by Proposition 3.13(1). Finally, $(\eta_{1}:\overline{\theta})=\overline{\alpha}$ by Lemma 2.3(2) and our choice of $\alpha$, so $(\Delta:\overline{\alpha})=\alpha$. This completes the proof of (2).
Now assume $0\prec\theta$. Then $\eta_{i}\prec\overline{\theta}$ for $i=1,2$. Since $\eta_{1}\wedge\eta_{2}=0$, we get $0\prec\eta_{1},\eta_{2}$ (see e.g. [26, Corollary 2.29(ii)]), and every maximal chain in the interval $I[0,\overline{\theta}]$ has length 2 [26, Theorem 2.37]. Thus $I[0,\overline{\theta}]$ has height 2. Note that this also implies $0\prec\varepsilon\prec\overline{\theta}$. Because $\overline{\theta}/\varepsilon$ is abelian and $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, the map $J:I[\varepsilon,\overline{\theta}]\rightarrow I[\Delta,\overline{\alpha}]$ given by $J(\gamma)=\gamma\vee\varepsilon$ is surjective by Proposition 3.12(2). Thus $\Delta\prec\overline{\alpha}$. Now let $\varphi\in\operatorname{Con}({\mathbf{A}}(\mu))$ with $\Delta\leq\varphi$ but $\overline{\alpha}\nleq\varphi$. Then $\overline{\alpha}\wedge\varphi=\Delta$, so $C(\varphi,\overline{\alpha};\Delta)$, so $\varphi\leq(\Delta:\overline{\alpha})=\overline{\alpha}$, so $\varphi=\overline{\alpha}\wedge\varphi=\Delta$. This proves (6). ∎
The next definition is from [5, 6] in the congruence modular setting.
Definition 4.7.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\theta\in\operatorname{Con}({\mathbf{A}})$ where $\theta$ is abelian. Also let $\alpha=(0:\theta)$. Then $D({\mathbf{A}},\theta)$ denotes the algebra ${\mathbf{A}}(\theta)/\Delta_{\theta,\alpha}$.
Corollary 4.8.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is an abelian congruence of ${\mathbf{A}}$, and $\alpha=(0:\theta)$. Let $\varphi:=\overline{\alpha}/\Delta_{\theta,\alpha}$ and $D^{o}=\{(a,a)/\Delta_{\theta,\alpha}\,:\,\text{$a\in A$}\}$. Then $\varphi$ is an abelian congruence of $D({\mathbf{A}},\theta)$, $(0:\varphi)=\varphi$, and there exist a surjective homomorphism $h:{\mathbf{A}}(\theta)\rightarrow D({\mathbf{A}},\theta)$ and an isomorphism $h^{\ast}:{\mathbf{A}}/\alpha\cong D({\mathbf{A}},\theta)/\varphi$ such that:
  1. (1)
    $D^{o}$ is a transversal for $\varphi$ and $D^{o}=\{0_{C}\,:\,\text{$C\in A/\alpha$}\}\leq D({\mathbf{A}},\theta)$.
  2. (2)
    $h^{-1}(D^{o})=0_{A}$.
  3. (3)
    $h$ and $h^{\ast}$ are compatible in the following sense: for all $(a,b)\in\theta$, $h(a,b)/\varphi=h^{\ast}(a/\alpha)$.
If in addition $\theta$ is a minimal congruence of ${\mathbf{A}}$, then $D({\mathbf{A}},\theta)$ is subdirectly irreducible with monolith $\varphi$.
Proof.
That $\varphi$ is abelian and $(0:\varphi)=\varphi$ follows from Theorem 4.6(2). Let $\Delta=\Delta_{\theta,\alpha}$. For each $\alpha$-class $C$ observe that $0_{C}=\{(c,c)\,:\,\text{$c\in C$}\}$ is a $\Delta$-class by Theorem 4.6(1). Hence $D^{o}=\{0_{C}\,:\,\text{$C\in A/\alpha$}\}$. It is easy to check that $D^{o}\leq D({\mathbf{A}},\theta)$. Given $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(a,a)$ and so $(a,b)/\Delta\stackrel{{\scriptstyle\varphi}}{{\equiv}}(a,a)/\Delta\in D^{o}$. Thus $D^{o}$ meets every $\varphi$-class. Suppose $0_{C},0_{C^{\prime}}\in D^{o}$ (so $C,C^{\prime}\in A/\alpha$) and $0_{C}\stackrel{{\scriptstyle\varphi}}{{\equiv}}0_{C^{\prime}}$. Pick $c\in C$ and $c^{\prime}\in C^{\prime}$. Then $(c,c)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(c^{\prime},c^{\prime})$, so $(c,c^{\prime})\in\alpha$, so $C=C^{\prime}$. Thus $D^{o}$ is a transversal for $\varphi$.
Let $h:{\mathbf{A}}(\theta)\rightarrow D({\mathbf{A}},\theta)$ be the natural homomorphism $(a,b)\mapsto(a,b)/\Delta$. To prove item (2), we must show that for all $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(c,c)$ for some $c\in A$ if and only if $a=b$; and this follows from Theorem 4.6(1). Finally, we have obvious isomorphisms ${\mathbf{A}}/\alpha\cong{\mathbf{A}}(\theta)/\overline{\alpha}$ and ${\mathbf{A}}(\theta)/\overline{\alpha}\cong({\mathbf{A}}(\theta)/\Delta)/(\overline{\alpha}/\Delta)=D({\mathbf{A}},\theta)/\varphi$, and we take $h^{\ast}$ to be their composition: $h^{\ast}(a/\alpha)=((a,a)/\Delta)/\varphi$. It is easy to check that item (3) is satisfied.
Finally, if $\theta$ is a minimal congruence, then Theorem 4.6(6) implies that $D({\mathbf{A}},\theta)$ is subdirectly irreducible with monolith $\varphi$. ∎
5. The classes of an abelian congruence
Suppose ${\mathbf{A}}$ is an algebra in a variety $\mathcal{V}$ with a weak difference term, $\theta$ is an abelian congruence of ${\mathbf{A}}$, and $\alpha=(0:\theta)$. Assuming $\mathcal{V}$ is congruence modular, a classical result of Gumm [9, see 3.3] states that if $B,B^{\prime}$ are two $\theta$-classes contained in a common $\alpha$-class, then $|B|=|B^{\prime}|$ and in fact $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e^{\prime})$ are isomorphic for any $e\in B$ and $e^{\prime}\in B^{\prime}$. This result extends to varieties with a difference term, as we shall see in Corollary 5.3 below. However it fails in general for varieties with a weak difference term, and in Section 8 we will show just how badly it can fail there.
Nonetheless, we shall show in this section that the algebra $D({\mathbf{A}},\theta)$ defined in the previous section, with its abelian congruence $\varphi=\overline{\alpha}/\Delta$ satisfying $(0:\varphi)=\varphi$, has the following property: for every $\alpha$-class $C$, the abelian group $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ forms a sort of “universal domain” for all the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ with $e\in C$.
Definition 5.1.
Suppose ${\mathbf{A}}$ is an algebra in a variety $\mathcal{V}$ with a weak difference term, $\theta$ is an abelian congruence, $\alpha=(0:\theta)$, $\Delta=\Delta_{\theta,\alpha}$, and $\varphi=\overline{\alpha}/\Delta$. Given $e\in A$, the map $\chi_{e}:e/\theta\rightarrow D({\mathbf{A}},\theta)$ is defined by $\chi_{e}(x)=(x,e)/\Delta$.
Lemma 5.2.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1.
  1. (1)
    For all $e\in A$, the map $\chi_{e}$ is an embedding of $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ into $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{e/\alpha})$.
  2. (2)
    If $\mathcal{V}$ has a difference term, then $\chi_{e}$ is an isomorphism.
Proof.
(1) Let $C=e/\alpha$. Clearly $\chi_{e}(e)=0_{C}$. If $(x,e)/\Delta=(y,e)/\Delta$, then $x,y,e$ all belong to a common $\theta$-class. So
$(x,e)$ $\stackrel{{\scriptstyle\Delta}}{{\equiv}}(y,e),\quad(y,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(y,e),\quad(y,y)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(y,y).$
Applying the weak difference term $d$ to these pairs and using $(x,y),(y,e)\in\theta$ gives
$(x,y)=(d(x,y,y),d(e,e,y))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(d(y,y,y),d(e,e,y))=(y,y).$
By Theorem 4.6(1), we get $x=y$. Thus $\chi_{e}$ is injective. It is easily checked that $\chi_{e}$ preserves $+$.
(2) Suppose that $d(x,y,z)$ is a difference term, i.e., $\mathcal{V}\models d(x,x,y)\approx y$. Given $(r,s)/\Delta\in 0_{C}/\varphi$, so $(r,s)\in\theta\cap C^{2}$, define $a:=d(r,s,e)$. Then $a\stackrel{{\scriptstyle\theta}}{{\equiv}}d(r,r,e)=e$, so $a\in e/\theta$. Observe that
$\begin{pmatrix}r&r\\ s&s\end{pmatrix},\ \begin{pmatrix}s&s\\ s&s\end{pmatrix},\begin{pmatrix}s&e\\ s&e\end{pmatrix}\in M(\theta,\alpha).$
Also observe that $d(r,s,s)=r$ as $(r,s)\in\theta$. Thus applying $d(x,y,z)$ to the above matrices gives
$\begin{pmatrix}r&a\\ s&e\end{pmatrix}=\begin{pmatrix}d(r,s,s)&d(r,s,e)\\ d(s,s,s)&d(s,s,e)\end{pmatrix}\in M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha).$
Hence $(r,s)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,e)$, so $\chi_{e}(a)=(r,s)/\Delta$, proving $\chi_{e}$ is surjective. ∎
Corollary 5.3.
Suppose ${\mathbf{A}}$ is an algebra in a variety with a difference term, $\theta$ is an abelian congruence, $\alpha=(0:\theta)$, and $(e,e^{\prime})\in\alpha$. Then $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)\cong\operatorname{Grp}_{{\mathbf{A}}}(\theta,e^{\prime})$.
Proof.
By Lemma 5.2, $\chi_{e^{\prime}}^{-1}\circ\chi_{e}$ is an isomorphism. ∎
Lemma 5.4.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1.
  1. (1)
    If $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b\stackrel{{\scriptstyle\theta}}{{\equiv}}e$, then $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(d(a,b,e),e)$.
  2. (2)
    If $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b\stackrel{{\scriptstyle\theta}}{{\equiv}}c$ and $(c,e)\in\alpha$, then $(d(a,e,c),d(b,e,c))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$.
  3. (3)
    For all $(e,e^{\prime})\in\theta$ we have $\operatorname{ran}(\chi_{e})=\operatorname{ran}(\chi_{e^{\prime}})$.
Proof.
(1) Let $x=d(a,b,c)$. Observe that
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\ \begin{pmatrix}b&b\\ b&b\end{pmatrix},\begin{pmatrix}c&b\\ c&b\end{pmatrix}\in M(\theta,\theta).$
Applying $d$ to these matrices gives
$\begin{pmatrix}x&a\\ c&b\end{pmatrix}=\begin{pmatrix}d(a,b,c)&d(a,b,b)\\ d(b,b,c)&d(b,b,b)\end{pmatrix}\in M(\theta,\theta)\subseteq M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha).$
Hence $(x,c)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ as required.
(2) This is proved similarly to (1), by applying $d$ to the matrices
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\quad\begin{pmatrix}c&e\\ c&e\end{pmatrix},\quad\begin{pmatrix}c&c\\ c&c\end{pmatrix}.$
(3) By (1), every $(a,e)/\Delta\in\operatorname{ran}(\chi_{e})$ can be written as $(d(a,e,e^{\prime}),e^{\prime})/\Delta$, proving $(a,e)/\Delta\in\operatorname{ran}(\chi_{e^{\prime}})$. Thus $\operatorname{ran}(\chi_{e})\subseteq\operatorname{ran}(\chi_{e^{\prime}})$, which by symmetry suffices. ∎
The previous lemma justifies the next definition.
Definition 5.5.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. If $B$ is a $\theta$-class and $C$ is the $\alpha$-class containing $B$, then let $\operatorname{ran}(B)$ denote the subgroup of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ with universe $\operatorname{ran}(\chi_{e})$ for any $e\in B$.
The next definition and Lemma contain an important tool that will help us understand the arrangement of ranges of $\theta$-classes in their corresponding $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$.
Definition 5.6.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. Let $B,B^{\prime}$ be $\theta$-classes contained in a common $\alpha$-class. We write $B\rightarrow B^{\prime}$ to mean there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(B)\subseteq B^{\prime}$ and such that $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ for all $a,b\in B$. We call such $f$ with these properties a witness to $B\rightarrow B^{\prime}$.
Lemma 5.7.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1.
  1. (1)
    $\rightarrow$ is reflexive and transitive.
  2. (2)
    If $B\rightarrow B^{\prime}$, then $\operatorname{ran}(B)\subseteq\operatorname{ran}(B^{\prime})$.
  3. (3)
    For all $\theta$-classes $B_{1},B_{2}$ contained in a common $\alpha$-class $C$, there exists a $\theta$-class $B^{\prime}\subseteq C$ with $B_{1}\rightarrow B^{\prime}$ and $B_{2}\rightarrow B^{\prime}$.
  4. (4)
    If $\mathcal{V}$ has a difference term, then $B_{1}\rightarrow B_{2}$ for all $\theta$-classes $B_{1},B_{2}$ contained in a common $\alpha$-class.
  5. (5)
    If $B\rightarrow B^{\prime}$, then for all $e\in B$ and $e^{\prime}\in B$ there exists a witness $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ to $B\rightarrow B^{\prime}$ satisfying $f(e)=e^{\prime}$.
Proof.
(1) is obvious. To prove (2), choose a witness $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ to $B\rightarrow B^{\prime}$, fix $e\in B$, and let $e^{\prime}=f(e)$. It suffices to show $\operatorname{ran}(\chi_{e})\subseteq\operatorname{ran}(\chi_{e^{\prime}})$. Let $a\in B$. Then $\chi_{e}(a)=(a,e)/\Delta=(f(a),e^{\prime})/\Delta\in\operatorname{ran}(\chi_{e^{\prime}})$.
(3) Given $B_{1},B_{2}$ contained in the common $\alpha$-class $C$, choose $e_{i}\in B_{i}$ for $i=1,2$ and let $B^{\prime}$ be the $\theta$-class containing $d(e_{1},e_{1},e_{2})$. $B^{\prime}\subseteq C$ by idempotence of $d$. Define $f_{1}(x)=d(x,e_{1},e_{2})$ and $f_{2}(x)=d(e_{1},e_{1},x)$. Then $f_{1},f_{2}\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $f_{i}(B_{i})\subseteq B^{\prime}$ for $i=1,2$. To see that $f_{1}$ witnesses $B_{1}\rightarrow B^{\prime}$, let $a,b\in B_{1}$ and note that
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\begin{pmatrix}e_{1}&e_{1}\\ e_{1}&e_{1}\end{pmatrix},\begin{pmatrix}e_{2}&e_{1}\\ e_{2}&e_{1}\end{pmatrix}$
are $(\theta,\alpha)$-matrices. Applying $d$ to these matrices gives
$\begin{pmatrix}f_{1}(a)&a\\ f_{1}(b)&b\end{pmatrix}\in M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha),$
which proves $(f_{1}(a),f_{1}(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ (see Lemma 4.4(1)). That $f_{2}$ witnesses $B_{2}\rightarrow B^{\prime}$ is proved similarly; if $a,b\in B_{2}$, then apply $d$ to the matrices
$\begin{pmatrix}e_{1}&e_{2}\\ e_{1}&e_{2}\end{pmatrix},\begin{pmatrix}e_{1}&e_{2}\\ e_{1}&e_{2}\end{pmatrix},\begin{pmatrix}a&a\\ b&b\end{pmatrix}.$
(4) If $d$ is a difference term, then in the proof of (3) we will have $d(e_{1},e_{1},e_{2})=e_{2}$ and so $B^{\prime}=B_{2}$.
(5) Suppose $f$ witnesses $B\rightarrow B^{\prime}$, $e\in B$, and $e^{\prime}\in B^{\prime}$. Define $g(x)=d(f(x),f(e),e^{\prime})$. Then $g\in\operatorname{Pol}_{1}({\mathbf{A}})$, $g(e)=e^{\prime}$ (so $g(B)\subseteq B^{\prime}$), and for all $a,b\in B$,
$(g(a),g(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$
where the first equality is by Lemma 5.4(2). ∎
Corollary 5.8.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1. Let $C\in A/\alpha$.
  1. (1)
    The set $S_{C}:=\{\operatorname{ran}(B)\,:\,\text{$B\in A/\theta$ and $B\subseteq C$}\}$ is a directed set of subgroups of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ whose union is $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$.
  2. (2)
    Hence if $C$ is finite, then there exists $e\in C$ with $\operatorname{ran}(\chi_{e})=\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$.
Proof.
For all $B_{1},B_{2}\in A/\theta$ with $B_{1},B_{2}\subseteq C$, there exists $B\in A/\theta$ with $\operatorname{ran}(B_{1})\cup\operatorname{ran}(B_{2})\subseteq\operatorname{ran}(B)$ by Lemma 5.7; this means $S_{C}$ is directed. To show that the union of $S_{C}$ is all of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$, let $(a,b)/\Delta\in\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$. Let $B=b/\theta$; then $a\in B\subseteq C$ so $(a,b)/\Delta=\chi_{b}(a)\in\operatorname{ran}(\chi_{b})=\operatorname{ran}(B)\subseteq\bigcup S_{C}$. ∎
If ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1 and $\mathcal{V}$ is congruence modular, then $\operatorname{ran}(\chi_{e})=\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ for all $e\in C\in A/\alpha$ by Lemma 5.2(2). Thus all $\theta$-classes in $C$ have the same cardinality. As suggested by Corollary 5.8, this uniformity is no longer to be expected in varieties with a weak difference term. (In Section 8, we will show that essentially nothing more about the sizes of $\theta$-classes can be proved than what is implied by Corollary 5.8 and Theorem 3.8, even when ${\mathbf{A}}$ is finite and $\theta$ is a minimal congruence.) However, we do know of one nontrivial restriction on the sizes of the classes of an abelian minimal congruence $\theta$ when the algebra is idempotent and $(0:\theta)=1$.
Lemma 5.9.
Suppose ${\mathbf{A}}$ is a finite idempotent algebra in a variety with a Taylor term, $\theta$ is an abelian minimal congruence, and $(0:\theta)=1$. Let $q$ be the maximal size of a $\theta$-class. Then every $\theta$-class has size $1$ or $q$.
Proof.
$\mathsf{HSP}({\mathbf{A}})$ has a weak difference term by finiteness of ${\mathbf{A}}$ and Theorem 2.5. Pick a $\theta$-class $B_{\mathsf{max}}$ of size $q$. Suppose there exists a $\theta$-class $B$ with $1<|B|<q$. Let $\Delta=\Delta_{\theta,1}$. Fix $e\in B_{\mathsf{max}}$ and $0\in B$. We have $\operatorname{ran}(\chi_{0})\subset A(\theta)/\Delta=\operatorname{ran}(\chi_{e})$ by Lemma 5.2 and Corollary 5.8. Fix $b\in B\setminus\{0\}$ and let $b^{\prime}\in B_{\mathsf{max}}\setminus\{e\}$ be such that $\chi_{0}(b)=\chi_{e}(b^{\prime})$. Also pick $u\in B_{\mathsf{max}}$ with $\chi_{e}(u)\not\in\operatorname{ran}(\chi_{0})$.
By Corollary 3.2, there exists a unary polynomial $f\in\operatorname{Pol}_{1}({\mathbf{B}})$ with $f(b^{\prime})=u$ and $f(e)=e$. Choose a $(1+n)$-ary term $t(x,\mathbf{y})$ and $\mathbf{c}\in A^{n}$ so that $f(x)=t(x,\mathbf{c})$. Let $u^{\prime}=t(b,0,\ldots,0)$. We have the following:
$(b^{\prime},e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,0),\quad(c_{1},c_{1})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(0,0),\quad\ldots,\quad(c_{n},c_{n})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(0,0)$
where the first congruence holds because $\chi_{e}(b^{\prime})=\chi_{0}(b)$ while the remaining congruences are because $(0:\theta)=1$. Now applying $t$ coordinatewise and using idempotency gives
$(u,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(u^{\prime},0)$
which contradicts our choice of $u$. ∎
6. Abelian minimal congruences
In this section we will explore further the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ where ${\mathbf{A}}$ is an algebra with a weak difference term, $\theta$ is an abelian congruence, and $\theta$ is minimal, i.e., $0\prec\theta$. Under this added assumption of minimality, we explain in detail how to construct a division ring $\mathbb{F}$ and equip each $\theta$-class with the structure of a left vector space over $\mathbb{F}$. The construction of $\mathbb{F}$ was sketched by Freese in [5, Section 4] assuming ${\mathbf{A}}$ belongs to a congruence modular variety. We describe the construction in somewhat more detail here, so that it will be apparent that congruence modularity is not needed. In case ${\mathbf{A}}$ belongs to a variety with a weak difference term, we compare $\mathbb{F}$ to the division ring similarly obtained for $D({\mathbf{A}},\theta)$ and its monolith, and show that they are isomorphic. We will not need these results in Section 7 on similarity; however, they will be used in an essential way in our subsequent paper on rectangular critical relations [32].
Notation 6.1.
Throughout this section, let ${\mathbf{A}}$ be an algebra with a weak difference term, and let $\theta\in\operatorname{Con}({\mathbf{A}})$ be minimal and abelian. Let the $\theta$-classes be indexed by a set $I$ via the notation $A/\theta=\{B_{i}\,:\,\text{$i\in I$}\}$. Let $T$ be a transversal for $\theta$. For each $i\in I$ let $0_{i}$ denote the unique element in $B_{i}\cap T$.
  1. (1)
    For all $i,j\in I$, $R_{ij}$ denotes the set of restrictions to $B_{j}$ of unary polynomials of ${\mathbf{A}}$ which send $0_{j}$ to $0_{i}$. (Thus $R_{ij}$ is a set of functions from $B_{j}$ to $B_{i}$.)
  2. (2)
    $\mathbb{F}$ denotes the set of all indexed families $\lambda=(\lambda_{i})_{i\in I}$ such that
    1. (a)
      $\lambda_{i}\in\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$ for each $i\in I$; and
    2. (b)
      for all $i,j\in I$ and $r\in R_{ij}$ we have $\lambda_{i}\circ r=r\circ\lambda_{j}$.
Proposition 6.2.
Let ${\mathbf{A}},\theta,T,(B_{i})_{i\in I},(0_{i})_{i\in I},(R_{ij})_{i,j\in I}$ and $\mathbb{F}$ be as in Notation 6.1.
  1. (1)
    For all $i,j\in I$, $R_{ij}$ is a subgroup of $\operatorname{Hom}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j}),\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$.
  2. (2)
    For all $i,j,k\in I$, if $r\in R_{ij}$ and $s\in R_{jk}$, then $r\circ s\in R_{ik}$.
  3. (3)
    For all $i\in I$, $R_{ii}$ is a unital subring of $\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$.
  4. (4)
    $\mathbb{F}$ is a unital subring of $\prod_{i\in I}\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$ and is a division ring.
Proof sketch.
(1) Let $d$ be a weak difference term for ${\mathbf{A}}$. Fix $i,j\in I$. Let $+_{i}$ and $+_{j}$ denote the group operations of $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ and $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j})$ respectively. Let $G$ be the group $\operatorname{Hom}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j}),\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$. Let $O_{ij}$ be the constant map $B_{j}\mapsto\{0_{i}\}$. Clearly $O_{ij}\in R_{ij}$ so $R_{ij}\neq\varnothing$. Fix $r\in R_{ij}$. Then for all $x,y\in B_{j}$,
$r(x+_{j}y)=r(d(x,0_{j},y))\stackrel{{\scriptstyle\ast}}{{=}}d(r(x),r(0_{j}),r(y))=r(x)+_{i}r(y)$
where $\stackrel{{\scriptstyle\ast}}{{=}}$ is by Proposition 3.4. This proves $R_{ij}\subseteq G$. Now suppose $r,s\in R_{ij}$ and let $t=r+s$ in the group $G$. Then for all $x\in B_{j}$, $t(x)=d(r(x),0_{i},s(x))$, so $t$ is the restriction to $B_{j}$ of a unary polynomial of ${\mathbf{A}}$, proving $t\in R_{ij}$.
(2) is obvious. (3) follows from (1) and (2) and the fact that $\mathrm{id}_{B_{i}}\in R_{ii}$.
(4) Let $\mathcal{R}$ be the set of $I\times I$ matrices $(r_{ij})_{i,j\in I}$ where $r_{ij}\in R_{ij}$ for all $i,j\in I$, and for each $j\in I$, $r_{ij}=O_{ij}$ for all but finitely many $i\in I$. Because of items (1)–(3), we can turn $\mathcal{R}$ into a unital ring in the usual “matrix” way: addition is defined coordinatewise in the groups $(R_{ij},+)$, and multiplication is defined by $(r_{ij})_{i,j\in I}\cdot(s_{jk})_{j,k\in I}:=(t_{ik})_{i,k\in I}$ where
$t_{ik}:=\sum_{j\in I,\,s_{jk}\neq O_{jk}}r_{ij}\circ s_{jk}\quad\mbox{computedin $(R_{ik},+)$}.$
Let $V$ denote the direct sum of the abelian groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$, $i\in I$, realized as the subgroup of $\prod_{i\in I}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ consisting of those $(a_{i})_{i\in I}\in\prod_{i\in I}B_{i}$ for which $a_{j}=0_{j}$ for all but finitely many $j\in I$. $\mathcal{R}$ acts in the usual “matrix” way on $V$: given $M=(r_{ij})_{i,j\in I}\in\mathcal{R}$ and $\mathbf{a}=(a_{i})_{i\in I}\in V$, define $M\mathbf{a}:=(b_{i})_{i\in I}$ where
$b_{i}=\sum_{j\in I,\,a_{j}\neq 0_{j}}r_{ij}(a_{j})\quad\mbox{computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$}.$
Via this action we obtain a faithful left $\mathcal{R}$-module ${}_{\mathcal{R}}V$. It is well known (see e.g. [5, p. 150]) that minimality of $\theta$ implies that ${}_{\mathcal{R}}V$ is a simple module. Here is a direct proof of this. Let $\overline{0}=(0_{i})_{i\in I}\in V$. The claim of simplicity reduces to proving the following: if $\mathbf{a}=(a_{i})_{i\in I}\in V$ with $\mathbf{a}\neq\overline{0}$, $\mathbf{b}=(b_{i})_{i\in I}\in V$, $k\in I$, and $b_{i}=0_{i}$ for all $i\in I$ with $i\neq k$, then $\mathbf{b}\in\mathcal{R}\mathbf{a}$. To prove it, pick $\ell\in I$ with $a_{\ell}\neq 0_{\ell}$. By minimality of $\theta$ we have $(b_{k},0_{k})\in\operatorname{Cg}(a_{\ell},0_{\ell})$. Thus by Corollary 3.2, there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(a_{\ell})=b_{k}$ and $f(0_{\ell})=0_{k}$. Let $M=(r_{ij})_{i,j\in I}\in\mathcal{R}$ be the matrix given by $r_{k\ell}=f{\upharpoonright}_{B_{\ell}}$ and $r_{ij}=O_{ij}$ for $(i,j)\neq(k,\ell)$. Then $M\mathbf{a}=\mathbf{b}$ as required.
Since ${}_{\mathcal{R}}V$ is a simple faithful module, it follows by Schur’s Lemma for modules that the endomorphism ring $\operatorname{End}_{\mathcal{R}}(V)$ of ${}_{\mathcal{R}}V$ is a division ring. For clarity, we consider both the endomorphisms of ${}_{\mathcal{R}}V$ and the endomorphisms of $V$ to act on the left. Thus $\operatorname{End}_{\mathcal{R}}(V)$ is the unital subring of $\operatorname{End}(V)$ consisting of those $f\in\operatorname{End}(V)$ for which $f(M\mathbf{a})=Mf(\mathbf{a})$ for all $M\in\mathcal{R}$ and all $\mathbf{a}\in V$.
Now let ${\mathbf{R}}$ be the unital ring $\prod_{i\in I}\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$. Define a unital ring embedding $\mathsf{L}:{\mathbf{R}}\hookrightarrow\operatorname{End}(V)$ in the obvious way: given $\lambda=(\lambda_{i})_{i\in I}\in R$ and $\mathbf{a}=(a_{i})_{i\in I}\in V$, define $\mathsf{L}_{\lambda}(\mathbf{a}):=(\lambda_{i}(a_{i}))_{i\in I}$. We claim that $\operatorname{End}_{\mathcal{R}}(V)\subseteq\operatorname{ran}(\mathsf{L})$ and $\mathbb{F}=\mathsf{L}^{-1}(\operatorname{End}_{\mathcal{R}}(V))$, which will prove (4). The first claim can be deduced from the equation $f(M\mathbf{a})=Mf(\mathbf{a})$ specialized to matrices $M$ which, for some $k\in I$, have $r_{kk}=\mathrm{id}_{B_{k}}$ and $r_{ij}=O_{ij}$ for all $(i,j)\neq(k,k)$. To prove the second claim, for each $k,\ell\in I$ and $r\in R_{k\ell}$ let $E_{k\ell}^{r}\in\mathcal{R}$ be the matrix $(r_{ij})_{i,j\in I}$ where $r_{k\ell}=r$ and $r_{ij}=O_{ij}$ for all $(i,j)\neq(k,\ell)$. For each $k\in I$ let $\varepsilon_{k}:\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{k})\hookrightarrow V$ be the obvious group embedding: $\varepsilon_{k}(a)=(a_{i})_{i\in I}$ where $a_{k}=a$ and $a_{i}=0_{i}$ for all $i\neq k$. Now fix $\lambda=(\lambda_{i})_{i\in I}\in R$ and let $f=\mathsf{L}_{\lambda}$. If $f\in\operatorname{End}_{\mathcal{R}}(V)$, then specializing the condition $f(M\mathbf{a})=Mf(\mathbf{a})$ to $M=E_{k\ell}^{r}$ and $\mathbf{a}\in\operatorname{ran}(\varepsilon_{\ell})$, we get $\lambda_{k}(r(a))=r(\lambda_{\ell}(a))$ for all $a\in B_{\ell}$. This proves $\lambda\in\mathbb{F}$. Conversely, if $\lambda\in\mathbb{F}$ and $M=(r_{ij})_{i,j\in I}\in\mathcal{R}$, then for all $j\in I$ and $\mathbf{a}\in\operatorname{ran}(\varepsilon_{j})$ we get $\mathsf{L}_{\lambda}(M\mathbf{a})=(\lambda_{i}(r_{ij}(a_{j})))_{i\in I}=(r_{ij}(\lambda_{j}(a_{j})))_{i\in I}=M\mathsf{L}_{\lambda}(\mathbf{a})$. By linearity, we get $\mathsf{L}_{\lambda}(M\mathbf{a})=M\mathsf{L}_{\lambda}(\mathbf{a})$ for all $\mathbf{a}\in V$, proving $\mathsf{L}_{\lambda}\in\operatorname{End}_{\mathcal{R}}(V)$. ∎
Definition 6.3.
Let ${\mathbf{A}},\theta,T,(0_{i})_{i\in I}$, and $\mathbb{F}$ be as in Notation 6.1.
  1. (1)
    We call $\mathbb{F}$ the division ring associated to ${\mathbf{A}}$, $\theta$ and $T$.
  2. (2)
    Given $i\in I$, the canonical action of $\mathbb{F}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ is defined as follows: for $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$ and $a\in\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$, put $\lambda\cdot a:=\lambda_{i}(a)$.
The canonical action of $\mathbb{F}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ turns $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ into a left $\mathbb{F}$-vector space, which we denote by ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$. Observe that Proposition 6.2(1) and condition 6.1(2b) imply that each $r\in R_{ij}$ is an $\mathbb{F}$-linear transformation from ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j})$ to ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$.
If $i\in I$ and $e\in B_{i}$ with $e\neq 0_{i}$, then we can obtain an action of $\mathbb{F}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ “by translation.” Choose a weak difference term $d$ for ${\mathbf{A}}$. Given any $(a,b)\in\theta$, let $\tau_{a}^{b}:\operatorname{Grp}_{{\mathbf{A}}}(\theta,a)\cong\operatorname{Grp}_{{\mathbf{A}}}(\theta,b)$ be the isomorphism given by $\tau_{a}^{b}(x)=d(x,a,b)$. Then for $\lambda\in\mathbb{F}$ and $x\in B_{i}$ define $\lambda\cdot^{e}x:=\tau_{0_{i}}^{e}(\lambda\cdot\tau_{e}^{0_{i}}(x))$. This expands $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ to a left $\mathbb{F}$-vector space which we may denote by ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$.
The construction of $\mathbb{F}$ and its actions on the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ depended on our choice of the transversal $T$ for $\theta$. Though we will not prove it here, it is not hard to show that if $T^{\prime}$ is any other choice of transversal, the division ring $\mathbb{F}^{\prime}$ associated to ${\mathbf{A}},\theta,T^{\prime}$ is isomorphic to $\mathbb{F}$ by the obvious “conjugation” isomorphism: letting $0_{i}^{\prime}$ be the unique element of $B_{i}\cap T^{\prime}$ and defining $\tau_{i}:B_{i}\rightarrow B_{i}$ by $\tau_{i}(x)=d(x,0_{i},0_{i}^{\prime})$, the isomorphism $\mathbb{F}\rightarrow\mathbb{F}^{\prime}$ sends $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$ to $(\tau_{i}\circ\lambda_{i}\circ\tau_{i}^{-1})_{i\in I}\in\mathbb{F}^{\prime}$. Modulo this isomorphism, $\mathbb{F}$ and $\mathbb{F}^{\prime}$ induce the same actions on the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. (In the special case when ${\mathbf{A}}$ belongs to a variety with a weak difference term, these claims follow from Lemma 6.15.)
In addition to providing a vector space structure to each $\theta$-class, the construction of the simple module ${}_{\mathcal{R}}V$ in the proof of Proposition 6.2 carries some useful information about the polynomials of ${\mathbf{A}}$ when restricted to $\theta$-classes. This information is sketched in [5, p. 151] in the congruence modular case; we develop it fully here.
Definition 6.4.
Let ${\mathbf{A}},\theta,T,(B_{i})_{i\in I},(0_{i})_{i\in I}$ be as in Notation 6.1. Let $\mathbb{F}$ be the division ring associated to ${\mathbf{A}},\theta,T$. Let $j_{1},\ldots,j_{n},i\in I$ and $f:B_{j_{1}}\times\cdots\times B_{j_{n}}\rightarrow B_{i}$. We say that $f$ is $\mathbb{F}$-affine if there exists an $\mathbb{F}$-linear transformation
$g:{}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j_{1}})\times\cdots\times{}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j_{n}}))\rightarrow{}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$
and an element $b\in B_{i}$ such that $f(\mathbf{a})=g(\mathbf{a})+b$ for all $\mathbf{a}\in B_{j_{1}}\times\cdots\times B_{j_{n}}$ (where the sum is in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$).
Proposition 6.5.
Let ${\mathbf{A}},\theta,T,(B_{i})_{i\in I}$ be as in Notation 6.1. Let $\mathbb{F}$ be the division ring associated to ${\mathbf{A}},\theta,T$. Fix $j_{1},\ldots,j_{n}\in I$.
  1. (1)
    Every restriction to $B_{i_{1}}\times\cdots\times B_{i_{n}}$ of an $n$-ary polynomial of ${\mathbf{A}}$ is $\mathbb{F}$-affine.
  2. (2)
    Conversely, every $\mathbb{F}$-affine function $f$ with domain $B_{j_{1}}\times\cdots\times B_{j_{n}}$ is finitely approximated by $n$-ary polynomials of ${\mathbf{A}}$. That is, for every finite subset $X\subseteq B_{j_{1}}\times\cdots\times B_{j_{n}}$ there exists $f_{X}\in\operatorname{Pol}_{n}({\mathbf{A}})$ such that $f(\mathbf{a})=f_{X}(\mathbf{a})$ for all $\mathbf{a}\in X$.
In particular, if $B_{j_{1}},\ldots,B_{j_{n}}$ are finite, then $\{f{\upharpoonright}_{B_{j_{1}}\times\cdots\times B_{j_{n}}}\,:\,\text{$f\in\operatorname{Pol}_{n}({\mathbf{A}})$}\}$ is the set of all $\mathbb{F}$-affine functions with domain $B_{j_{1}}\times\cdots\times B_{j_{n}}$.
Proof sketch.
(1) This can be shown by an easy adaptation of known arguments such as those found in [10, p. 32]. Given $f\in\operatorname{Pol}_{n}({\mathbf{A}})$ and $i\in I$ with $f(B_{j_{1}}\times\cdots\times B_{j_{n}})\subseteq B_{i}$, define $g(x_{1},\ldots,x_{n})=d(f(x_{1},\ldots,x_{n}),f(0_{j_{1}},\ldots,0_{j_{n}}),0_{i})$. Then $g(0_{j_{1}},\ldots,0_{j_{n}})=0_{i}$ and it will suffice to show that $g{\upharpoonright}_{B_{j_{1}}\times\cdots\times B_{j_{n}}}$ is $\mathbb{F}$-linear. For each $t=1,\ldots,n$ define $r_{t}:B_{j_{t}}\rightarrow B_{i}$ by $r_{t}(x)=g(0_{j_{1}},\ldots,0_{j_{t-1}},x,0_{j_{t+1}},\ldots,0_{j_{n}})$. Then for each $t$ we have that $r_{t}$ is in $R_{ij_{t}}$ and so is $\mathbb{F}$-linear by the comments following Definition 6.3. Furthermore, for all $(x_{1},\ldots,x_{n})\in B_{j_{1}}\times\cdots\times B_{j_{n}}$ we have
$g(x_{1},\ldots,x_{n})$ $=g(d(x_{1},0_{j_{1}},0_{j_{1}}),d(0_{j_{2}},0_{j_{2}},x_{2}),\ldots,d(0_{j_{n}},0_{j_{n}},x_{n}))$
$=d(g(x_{1},0_{j_{2}},\ldots,0_{j_{n}}),g(0_{j_{1}},0_{j_{2}},\ldots,0_{j_{n}}),g(0_{j_{1}},x_{2},\ldots,x_{n}))$
by Proposition 3.4, which simplifies to
$g(x_{1},\ldots,x_{n})=r_{1}(x_{1})+g(0_{j_{1}},x_{2},\ldots,x_{n})\quad\mbox{computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$}.$
Iterating this argument eventually gives
$g(x_{1},\ldots,x_{n})=\sum_{t=1}^{n}r_{i}(x_{i})\quad\mbox{computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$}$
which implies that $g{\upharpoonright}_{B_{j_{1}}\times\cdots\times B_{j_{n}}}$ is $\mathbb{F}$-linear.
(2) Recall the construction of the module ${}_{\mathcal{R}}V$ from the proof of Proposition 6.2(4). As we saw there, ${}_{\mathcal{R}}V$ is simple and $\mathsf{L}:\mathbb{F}\cong\operatorname{End}_{\mathcal{R}}(V)$. $V$ naturally inherits the structure of a left $\operatorname{End}_{\mathcal{R}}(V)$-vector space and thus also the structure of a left $\mathbb{F}$-vector space via $\mathsf{L}$, the latter which we denote by ${}_{\mathbb{F}}V$. Also for each $M\in\mathcal{R}$ let $\phi(M):V\rightarrow V$ be the group endomorphism given by $M$’s action: $\phi(M)(\mathbf{a}):=M\mathbf{a}$. Then $\phi(\mathcal{R})$ is a subring of $\operatorname{End}_{\mathbb{F}}(V)$, and by Jacobson’s Density Theorem, $\phi(\mathcal{R})$ is dense in $\operatorname{End}_{\mathbb{F}}(V)$.
We will first prove the claim in the following special case: if $i,j\in I$ and $f:{}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j})\rightarrow{}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ is $\mathbb{F}$-linear, then $f$ is finitely approximated by members of $R_{ij}$. Fix such an $f$. Let $X$ be a finite subset of $B_{j}$. Recall the embeddings $\varepsilon_{k}:\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{k})\hookrightarrow V$ from the proof of Proposition 6.2(4) and note that each $\varepsilon_{k}$ is actually an $\mathbb{F}$-linear map ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{k})\hookrightarrow{}_{\mathbb{F}}V$, and ${}_{\mathbb{F}}V$ is the direct sum of the spaces $\operatorname{ran}(\varepsilon_{k})$ ($k\in I$). Let $h:V\rightarrow V$ be the $\mathbb{F}$-linear map determined by $h{\upharpoonright}_{\operatorname{ran}(\varepsilon_{k})}=\overline{0}$ for $k\neq j$ and $h{\upharpoonright}_{\operatorname{ran}(\varepsilon_{j})}=\varepsilon_{i}\circ f\circ\varepsilon_{j}^{-1}$. Because $\phi(\mathcal{R})$ is dense in $\operatorname{End}_{\mathbb{F}}(V)$, there exists $M=(r_{k\ell})_{k,\ell\in I}\in\mathcal{R}$ with $h{\upharpoonright}_{\varepsilon_{j}(X)}=\phi(M){\upharpoonright}_{\varepsilon_{j}(X)}$. Then $r_{ij}\in R_{ij}$ and $f{\upharpoonright}_{X}=r_{ij}{\upharpoonright}_{X}$ as required.
Now we prove the claim in general. Suppose that $f:B_{j_{1}}\times\cdots\times B_{j_{n}}\rightarrow B_{i}$ is $\mathbb{F}$-affine. Then we can write
$f(x_{1},\ldots,x_{n})=\sum_{t=1}^{n}f_{t}(x_{t})+b\quad\mbox{computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$}$
for some $\mathbb{F}$-linear transformations $f_{t}:{}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{j_{t}})\rightarrow{}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ and some $b\in B_{i}$. To prove that $f$ is finitely approximated by members of $\operatorname{Pol}_{n}({\mathbf{A}})$, it is enough to show that each $f_{t}$ is finitely approximated by members of $R_{ij_{t}}$, which we did in the special case. ∎
We now focus on the special case where ${\mathbf{A}}$ belongs to a variety $\mathcal{V}$ with a weak difference term. Let $\alpha=(0:\theta)$, $\Delta=\Delta_{\theta,\alpha}$, and $\varphi=\overline{\alpha}/\Delta$, and recall that $D({\mathbf{A}},\theta)={\mathbf{A}}(\theta)/\Delta$ and $D^{o}=\{0_{C}/\Delta\,:\,\text{$C\in A/\alpha$}\}$. Then $D({\mathbf{A}},\theta)$ has a difference term, $\varphi$ is an abelian minimal congruence of $D({\mathbf{A}},\theta)$, and $D^{o}$ is a transversal for $\varphi$, by Theorem 4.6. Thus our previous construction associates a division ring $\overline{\mathbb{F}}$ to $D({\mathbf{A}},\theta)$, $\varphi$ and $D^{o}$. In the remainder of this section, we will give a relatively simple characterization of $\overline{\mathbb{F}}$, prove that $\overline{\mathbb{F}}$ is isomorphic to $\mathbb{F}$, describe a natural action of $\overline{\mathbb{F}}$ on the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$, and show that this action agrees with the action of $\mathbb{F}$ on these groups, modulo the isomorphism between $\overline{\mathbb{F}}$ and $\mathbb{F}$.
Notation 6.6.
Let ${\mathbf{A}},\theta,T,(B_{i})_{i\in I},(0_{i})_{i\in I},(R_{ij})_{i,j\in I}$, and $\mathbb{F}$ be as in Notation 6.1. Assume that ${\mathbf{A}}$ belongs to a variety with a weak difference term and let $\alpha,\Delta,\varphi$ and $D^{o}$ be as described in the previous paragraph. For the remainder of this section:
  1. (1)
    Let ${\mathbf{D}}$ denote $D({\mathbf{A}},\theta)$.
  2. (2)
    For $i,j\in I$, define $i\sim j\iff$ $B_{i}$ and $B_{j}$ are contained in a common $\alpha$-class.
  3. (3)
    For $i\in I$, define $[i]=\{j\in I\,:\,\text{$i\sim j$}\}$. Let $[I]=\{[i]\,:\,\text{$i\in I$}\}$.
  4. (4)
    For $J\in[I]$, let $C_{J}=\bigcup\{B_{j}\,:\,\text{$j\in J$}\}\subseteq A$ and $E_{J}=(\theta\cap(C_{J})^{2})/\Delta\subseteq D$.
Thus $\{C_{J}\,:\,\text{$J\in[I]$}\}$ is an enumeration of the $\alpha$-classes in ${\mathbf{A}}$, and $\{E_{J}\,:\,\text{$J\in[I]$}\}$ is an enumeration of the $\varphi$-classes in ${\mathbf{D}}$.
Notation 6.6 (continued).
  1. (5)
    For each $J\in[I]$, let $\overline{0}_{J}$ denote $0_{C_{J}}/\Delta$, which is the unique element of $E_{J}\cap D^{o}$.
  2. (6)
    For $J,K\in[I]$, let $\overline{R}_{JK}$ be defined as in Notation 6.1 but with ${\mathbf{D}},\varphi,D^{o}$, and $(E_{J})_{J\in[I]}$ in place of ${\mathbf{A}},\theta,T$, and $(B_{i})_{i\in I}$. That is, $\overline{R}_{JK}$ is the set of restrictions to $E_{K}$ of unary polynomials of ${\mathbf{D}}$ which send $\overline{0}_{K}$ to $\overline{0}_{J}$.
  3. (7)
    Let $\overline{\mathbb{F}}$ be the division ring associated to ${\mathbf{D}},\varphi$ and $D^{o}$. Thus $\overline{\mathbb{F}}$ is the subring of $\prod_{J\in[I]}\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J}))$ whose universe is the set of functions $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}$ satisfying
    1. (a)
      $\overline{\lambda}_{J}\in\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J}))$ for all $J\in[I]$.
    2. (b)
      For all $J,K\in[I]$ and $\overline{r}\in\overline{R}_{JK}$ we have $\overline{\lambda}_{J}\circ\overline{r}=\overline{r}\circ\overline{\lambda}_{K}$.
The next lemma gives a slightly simpler description of the sets $\overline{R}_{JK}$.
Definition 6.7.
For each $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ let $f^{{\mathbf{D}}}:D\rightarrow D$ be the function given by $f^{{\mathbf{D}}}((a,b)/\Delta)=(f(a),f(b))/\Delta$.
Lemma 6.8.
  1. (1)
    Let $f\in\operatorname{Pol}_{1}({\mathbf{A}})$. The map $f^{{\mathbf{D}}}$ is well-defined, in $\operatorname{Pol}_{1}({\mathbf{D}})$, and $f^{{\mathbf{D}}}(D^{o})\subseteq D^{o}$.
  2. (2)
    For all $J,K\in[I]$, $\overline{R}_{JK}=\{f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\,:\,\text{$f\in\operatorname{Pol}_{1}({\mathbf{A}})$, $f(C_{K})\subseteq C_{J}$}\}$.
  3. (3)
    For all $J,K\in[I]$, $j\in J$, $k\in K$, and $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(0_{k})\in B_{j}$, there exists $g\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $g(0_{k})=0_{j}$ and $g^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}$.
Proof.
(1) Suppose $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},b^{\prime})$. Applying compatibility of $f$ to $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},b^{\prime})$ gives $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(f(a^{\prime}),f(b^{\prime}))$, so $f^{{\mathbf{D}}}$ is well-defined. Write $f(x)=t(x,c_{1},\ldots,c_{m})$ where $t(x,\mathbf{y})$ is a term. Replacing each $c_{i}$ with $(c_{i},c_{i})/\Delta$ produces a polynomial of ${\mathbf{D}}$ which is $f^{{\mathbf{D}}}$. If $x\in D^{o}$ then $f^{{\mathbf{D}}}(x)=t(x,(c_{1},c_{1})/\Delta,\ldots,(c_{m},c_{m})/\Delta)\in D^{o}$ as each $(c_{i},c_{i})/\Delta$ is in $D^{o}$ and $D^{o}$ is a subuniverse of ${\mathbf{D}}$ by Corollary 4.8(1).
(2) Note first that if $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $f(C_{K})\subseteq C_{J}$, then obviously $f^{{\mathbf{D}}}(\overline{0}_{K})\in E_{J}$. Since $f^{{\mathbf{D}}}(\overline{0}_{K})\in D^{o}$ by (1), we get $f^{{\mathbf{D}}}(\overline{0}_{K})\in E_{J}\cap D^{o}=\{\overline{0}_{J}\}$, proving $f^{{\mathbf{D}}}(\overline{0}_{K})=\overline{0}_{J}$. This explains one inclusion.
To prove the other, suppose $g\in\operatorname{Pol}_{1}({\mathbf{D}})$ with $g(\overline{0}_{K})=\overline{0}_{J}$. Pick a term $t(x,\mathbf{y})$ and parameters $(u_{1},v_{1})/\Delta,\ldots,(u_{m},v_{m})/\Delta$ defining $g$. Then for every $a\in C_{K}$ and $e\in C_{J}$,
$(e,e)/\Delta=\overline{0}_{J}$ $=g(\overline{0}_{K})=g((a,a)/\Delta)$
$=t((a,a)/\Delta,(u_{1},v_{1})/\Delta,\ldots,(u_{m},v_{m})/\Delta)$
$=(t(a,\mathbf{u}),t(a,\mathbf{v}))/\Delta.$
By a property of $\Delta$ (i.e., Theorem 4.6(1)), this implies $t(a,\mathbf{u})=t(a,\mathbf{v})$ for all $a\in C_{K}$. Define $f(x):=t(x,\mathbf{u})\in\operatorname{Pol}_{1}({\mathbf{A}})$. Then for all $a\in C_{K}$ and $e\in C_{J}$ we’ve shown $(f(a),f(a))/\Delta=(e,e)/\Delta$, which, by definition of $\Delta$, forces $f(a)\stackrel{{\scriptstyle\alpha}}{{\equiv}}e$ and hence $f(C_{K})\subseteq C_{J}$. Finally, for all $(a,b)/\Delta\in E_{K}$ we have $a,b\in C_{K}$, so
$g((a,b)/\Delta)$ $=t((a,b)/\Delta,(u_{1},v_{1})/\Delta,\ldots,(u_{m},v_{m})/\Delta)$
$=(t(a,\mathbf{u}),t(b,\mathbf{v}))/\Delta$
$=(t(a,\mathbf{u}),t(b,\mathbf{u}))/\Delta$ by an earlier calculation
$=(f(a),f(b)/\Delta$
$=f^{{\mathbf{D}}}((a,b)/\Delta).$
Thus $g{\upharpoonright}_{E_{K}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}$ as required.
(3) Let $e=f(0_{k})\in B_{j}$ and $g(x)=d(f(x),e,0_{j})$. Clearly $g(0_{k})=0_{j}$. Given $(a,b)/\Delta\in E_{K}$, we have
$g^{{\mathbf{D}}}((a,b)/\Delta)=(g(a),g(b))/\Delta$ $=(d(f(a),e,0_{j}),d(f(b),e,0_{j}))/\Delta=f^{{\mathbf{D}}}((a,b)/\Delta),$
where the last equality is by Lemma 5.4(2). ∎
Just as we defined ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ for each $i\in I$, for each $J\in[I]$ we let ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$ denote the left $\overline{\mathbb{F}}$-vector space with underlying group $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$ and canonical left action of $\overline{\mathbb{F}}$ given by $\overline{\lambda}\cdot x=\overline{\lambda}_{J}(x)$. An important fact about this vector space is that every range of a $\theta$-class contained in $C_{J}$ is a subspace.
Lemma 6.9.
For every $J\in[I]$ and $e\in C_{J}$, $\operatorname{ran}(e/\theta)$ is a subspace of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$.
Proof.
Let $x\in\operatorname{ran}(e/\theta)=\operatorname{ran}(\chi_{e})$, so $x=(a,e)/\Delta$ for some $a\in e/\theta$. Let $\overline{\lambda}\in\overline{\mathbb{F}}$. If $a=e$ then $x=\overline{0}_{J}$ and $\overline{\lambda}_{J}(x)=x\in\operatorname{ran}(e/\theta)$. Assume $a\neq e$. Let $\overline{\lambda}_{J}(x)=(c,u)/\Delta$. Since $\overline{\lambda}_{J}:E_{J}\rightarrow E_{J}$ we get $c,u\in C_{J}$. Pick $j\in J$ with $c,u\in B_{j}$. Using Lemma 5.7(3), obtain $k\in J$ so that $B_{i}\rightarrow B_{k}$ and $B_{j}\rightarrow B_{k}$ witnessed by $f,g\in\operatorname{Pol}_{1}({\mathbf{A}})$ respectively. This means that $f(B_{i})\subseteq B_{k}$, $g(B_{j})\subseteq B_{k}$, $(f(v),f(z))/\Delta=(v,z)/\Delta$ for all $v,z\in B_{i}$, and $(g(v),g(z))/\Delta=(v,z)/\Delta$ for all $v,z\in B_{j}$. Let $a^{\prime}=f(a)$ and $e^{\prime}=f(e)$. By Lemma 5.7(5), we may assume that $g$ was chosen satisfying $g(u)=e^{\prime}$. Under this assumption, let $c^{\prime}=g(c)$.
Thus we have $a^{\prime},c^{\prime},e^{\prime}\in B_{k}$, $(a^{\prime},e^{\prime})/\Delta=x$, and $(c^{\prime},e^{\prime})/\Delta=\overline{\lambda}_{J}(x)$. Since $a\neq e$, we get $a^{\prime}\neq e^{\prime}$ by a property of $\Delta$ (Theorem 4.6(1)). As $(a,e),(a^{\prime},e^{\prime})\in\theta$ and $\theta$ is abelian and minimal, there exists $h\in\operatorname{Pol}_{1}({\mathbf{A}})$ satisfying $h(e^{\prime})=e$ and $h(a^{\prime})=a$ by Corollary 3.2. Then $h(B_{k})\subseteq B_{i}$, so $h(C_{J})\subseteq C_{J}$. Thus $h^{{\mathbf{D}}}{\upharpoonright}_{E_{J}}\in\overline{R}_{JJ}$ by Lemma 6.8. Hence $h^{{\mathbf{D}}}\circ\overline{\lambda}_{J}=\overline{\lambda}_{J}\circ h^{{\mathbf{D}}}{\upharpoonright}_{E_{J}}$ by condition (7b) in the definition of $\overline{\mathbb{F}}$. Also observe that $h^{{\mathbf{D}}}(x)=h^{{\mathbf{D}}}((a^{\prime},e^{\prime})/\Delta)=(h(a^{\prime}),h(e^{\prime}))/\Delta=(a,e)/\Delta=x$. Thus
$\overline{\lambda}_{J}(x)$ $=\overline{\lambda}_{J}(h^{{\mathbf{D}}}(x))$
$=h^{{\mathbf{D}}}(\overline{\lambda}_{J}(x))\qquad\qquad\mbox{as $x\in E_{J}$}$
$=h^{{\mathbf{D}}}((c^{\prime},e^{\prime})/\Delta)$
$=(h(c^{\prime}),h(e^{\prime}))/\Delta=(h(c^{\prime}),e)/\Delta\in\operatorname{ran}(\chi_{e}).$
Corollary 6.10.
Fix $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$. Fix $J\in[I]$ and $i\in J$. Then for all $a,e\in B_{i}$ there exists a unique $b\in B_{i}$ satisfying $\overline{\lambda}_{J}((a,e)/\Delta)=(b,e)/\Delta$.
Proof.
Existence follows from the proof of Lemma 6.9. Uniqueness follows from injectivity of $\chi_{e}$ (Lemma 5.2). ∎
This leads to a natural action of $\overline{\mathbb{F}}$ on each group $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ in ${\mathbf{A}}$.
Definition 6.11.
Given $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$ and $e\in A$, choose $J\in[I]$ so that $e\in C_{J}$ and define $\overline{\lambda}_{e}:e/\theta\rightarrow e/\theta$ so that, for each $a\in e/\theta$, $\overline{\lambda}_{e}(a)$ is the unique $b\in e/\theta$ satisfying $\overline{\lambda}_{J}((a,e)/\Delta)=(b,e)/\Delta$.
In other words, $\overline{\lambda}_{e}$ is determined by the equation $\chi_{e}\circ\overline{\lambda}_{e}=\overline{\lambda}_{J}\circ\chi_{e}$. It follows that $\overline{\lambda}_{e}=\chi_{e}^{-1}\circ\overline{\lambda}_{J}{\upharpoonright}_{\operatorname{ran}(B_{i})}\circ\chi_{e}$, which proves the next lemma.
Lemma 6.12.
Let $\overline{\lambda}\in\overline{\mathbb{F}}$ and $e\in A$. Then $\overline{\lambda}_{e}\in\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,e))$.
Now we define an action of $\overline{\mathbb{F}}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ in the obvious way: if $\overline{\lambda}\in\overline{\mathbb{F}}$ and $a\in e/\theta$, define $\overline{\lambda}\cdot a:=\overline{\lambda}_{e}(a)$. It can be routinely checked that this gives $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ the structure of a left $\overline{\mathbb{F}}$-vector space, which we denote by ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. We get the following improvement to Lemma 5.2 and Corollary 5.8 from Section 5.
Proposition 6.13.
Let ${\mathbf{A}}$ be an algebra in a variety with a weak difference term, let $\theta$ be an abelian minimal congruence, let $\alpha,\Delta,\varphi,D^{o},{\mathbf{D}}$ be as in Notation 6.6, and let $\overline{\mathbb{F}}$ be the division ring associated to ${\mathbf{D}},\varphi,D^{o}.$ Let $C$ be an $\alpha$-class.
  1. (1)
    For each $e\in C$, the map $\chi_{e}:a\mapsto(a,e)/\Delta$ is a left $\overline{\mathbb{F}}$-vector space embedding of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ into ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{C})$.
  2. (2)
    The set $S_{C}:=\{\operatorname{ran}(B)\,:\,\text{$B\in A/\theta$, $B\subseteq C$}\}$ is a directed set of subspaces of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{C})$ whose union is ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{C})$.
Recall that we also have the left $\mathbb{F}$-vector space ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ where $\mathbb{F}$ is the division ring associated to ${\mathbf{A}}$, $\theta$ and a transversal $T$ for $\theta$ which we fixed at the beginning of the section. Let $(0_{i})_{i\in I}$ be as in Notation 6.1. In the remainder of this section will show that ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ are essentially the same.
Definition 6.14.
Define $\Psi:\overline{\mathbb{F}}\rightarrow\prod_{i\in I}\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$ by $\Psi(\overline{\lambda})=(\overline{\lambda}_{0_{i}})_{i\in I}$.
Lemma 6.15.
Let ${\mathbf{A}},\theta,T,\alpha,\Delta,{\mathbf{D}},D^{o}$ be as in Notation 6.6. Let $\mathbb{F}$ be the division ring associated to ${\mathbf{A}},\theta,T$ and let $\overline{\mathbb{F}}$ be the division ring associated to ${\mathbf{D}},\varphi,D^{o}$.
  1. (1)
    $\operatorname{ran}(\Psi)=\mathbb{F}$.
  2. (2)
    $\Psi$ is a ring isomorphism from $\overline{\mathbb{F}}$ to $\mathbb{F}$.
  3. (3)
    Fix $e\in A$. If $\cdot$ denotes the action of $\overline{\mathbb{F}}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and $\cdot^{e}$ denotes the action of $\mathbb{F}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$, then for all $\overline{\lambda}\in\overline{\mathbb{F}}$ and $a\in e/\theta$, $\overline{\lambda}\cdot a=\Psi(\overline{\lambda})\cdot^{e}a$.
Proof.
(1) Let $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$, so $\Psi(\overline{\lambda})=(\overline{\lambda}_{0_{i}})_{i\in I}$. To show $\Psi(\overline{\lambda})\in\mathbb{F}$, we must verify condition (2b) in Notation 6.1. Let $j,k\in I$ and $r\in R_{jk}$; thus $r=f{\upharpoonright}_{B_{k}}$ for some $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(0_{k})=0_{j}$. Let $J=[j]$ and $K=[k]$. Then $f(C_{K})\subseteq C_{J}$, so $f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\in\overline{R}_{JK}$. Condition (7b) in Notation 6.6 gives $\overline{\lambda}_{J}\circ f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\overline{\lambda}_{K}$. Also note that for all $a\in B_{k}$,
$\chi_{0_{j}}(r(a))=(f(a),0_{j})/\Delta=(f(a),f(0_{k}))/\Delta=f^{{\mathbf{D}}}((a,0_{k})/\Delta)=f^{{\mathbf{D}}}(\chi_{0_{k}}(a))$
which proves $\chi_{0_{j}}\circ r=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\chi_{0_{k}}$. Combining this with $\chi_{0_{j}}\circ\overline{\lambda}_{0_{j}}=\overline{\lambda}_{J}\circ\chi_{0_{j}}$ gives
$\chi_{0_{j}}\circ\overline{\lambda}_{0_{j}}\circ r=\overline{\lambda}_{J}\circ\chi_{0_{j}}\circ r=\overline{\lambda}_{J}\circ f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\chi_{0_{k}}$
and similarly
$\chi_{0_{j}}\circ r\circ\overline{\lambda}_{0_{k}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\chi_{0_{k}}\circ\overline{\lambda}_{0_{k}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\overline{\lambda}_{K}\circ\chi_{0_{k}}.$
Since $\overline{\lambda}_{J}\circ f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\overline{\lambda}_{K}$ and $\chi_{0_{j}}$ is injective (Lemma 5.2), we get $\overline{\lambda}_{0_{j}}\circ r=r\circ\overline{\lambda}_{0_{k}}$ as required by condition (2b) in Notation 6.1. This proves $\Psi(\overline{\lambda})\in\mathbb{F}$, so $\operatorname{ran}(\Psi)\subseteq\mathbb{F}$.
Towards proving the opposite inclusion, we first show the following.
Claim 6.15.1.
Suppose $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$. Then for all $J\in[I]$ and $i,j\in J$, and all $a\in B_{i}$ and $a^{\prime}\in B_{j}$, if $(a,0_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},0_{j})$, then $(\lambda_{i}(a),0_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(\lambda_{j}(a^{\prime}),0_{j})$.
Proof of Claim 6.15.1.
Using Lemma 5.7(3), choose $k\in J$ with $B_{i},B_{j}\rightarrow B_{k}$ witnessed by $f,g\in\operatorname{Pol}_{1}({\mathbf{A}})$ respectively. By Lemma 5.7(5), we can assume that $f(0_{i})=g(0_{j})=0_{k}$. Thus $(f(a),0_{k})/\Delta=(a,0_{i})/\Delta=(a^{\prime},0_{j})/\Delta=(g(a^{\prime}),0_{k})/\Delta$ by the definition of $\rightarrow$, and so $f(a)=g(a^{\prime})$ by injectivity of $\chi_{0_{k}}$. Let $r=f{\upharpoonright}_{B_{i}}$ and $s=g{\upharpoonright}_{B_{j}}$. Then $r\in R_{ki}$ and $s\in R_{kj}$. Then
$(\lambda_{i}(a),0_{i})/\Delta$ $=(r(\lambda_{i}(a)),0_{k})/\Delta$ as $r=f{\upharpoonright}_{B_{i}}$ and $f$ witnesses $B_{i}\rightarrow B_{k}$
$=(\lambda_{k}(r(a)),0_{k})/\Delta$ as $\lambda_{k}\circ r=r\circ\lambda_{i}$
and similarly $(\lambda_{j}(a^{\prime}),0_{j})/\Delta=(\lambda_{k}(s(a^{\prime})),0_{k})$. As $r(a)=s(a^{\prime})$, we get $(\lambda_{i}(a),0_{i})/\Delta=(\lambda_{j}(a^{\prime}),0_{j})/\Delta$ as required. ∎
Observe that for each $J\in[I]$, every element of $E_{J}$ can be written in the form $(a,0_{j})/\Delta$ for some $j\in J$ and $a\in B_{j}$. This observation, combined with Claim 6.15.1, justifies the following definition.
Definition 6.15.2.
For $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$ and $J\in[I]$, define $\lambda_{J}:E_{J}\rightarrow E_{J}$ as follows: if $i\in J$ and $a\in B_{i}$, then $\lambda_{J}((a,0_{i})/\Delta):=(\lambda_{i}(a),0_{i})/\Delta$.
In other words, $\lambda_{J}$ is determined by the equations $\lambda_{J}\circ\chi_{0_{i}}=\chi_{0_{i}}\circ\lambda_{i}$ for all $i\in J$. It follows that for each $i\in J$, $\lambda_{J}{\upharpoonright}_{\operatorname{ran}(B_{i})}=\chi_{0_{i}}\circ\lambda_{i}\circ\chi_{0_{i}}^{-1}$. Since $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$ is the union of the directed family $\{\operatorname{ran}(B_{i})\,:\,\text{$i\in J$}\}$ of subgroups by Corollary 5.8, it follows that $\lambda_{J}\in\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J}))$.
Definition 6.15.3.
Define $\Phi:\mathbb{F}\rightarrow\prod_{J\in[I]}\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J}))$ as follows: given $\lambda\in\mathbb{F}$, set $\Phi(\lambda):=(\lambda_{J})_{J\in[I]}$.
Claim 6.15.4.
$\operatorname{ran}(\Phi)\subseteq\overline{\mathbb{F}}$ and $\Psi\circ\Phi=\mathrm{id}_{\mathbb{F}}$. Thus $\operatorname{ran}(\Psi)=\mathbb{F}$.
Proof of Claim 6.15.4.
Let $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$. To prove $\Phi(\lambda)\in\overline{\mathbb{F}}$, we must show $\lambda_{J}\circ\overline{r}=\overline{r}\circ\lambda_{K}$ for all $J,K\in[I]$ and all $\overline{r}\in\overline{R}_{JK}$. Fix $J,K\in[I]$ and $\overline{r}\in\overline{R}_{JK}$. By Corollary 5.8, it is enough to prove $\lambda_{J}\circ\overline{r}{\upharpoonright}_{\operatorname{ran}(B_{k})}=\overline{r}\circ\lambda_{K}{\upharpoonright}_{\operatorname{ran}(B_{k})}$ for all $k\in K$. Fix $k\in K$. Choose $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(C_{K})\subseteq C_{J}$ and such that $\overline{r}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}$. By Lemma 6.8(3), we can assume $f(0_{k})=0_{j}$ for some $j\in J$. Define $r:=f{\upharpoonright}_{B_{k}}$. Then $r\in R_{jk}$. Now for any $a\in B_{k}$,
$\lambda_{J}(\overline{r}((a,0_{k})/\Delta))$ $=\lambda_{J}((r(a),0_{j})/\Delta)$
$=(\lambda_{j}(r(a)),0_{j})/\Delta$ by definition of $\lambda_{J}$
$=(r(\lambda_{k}(a)),0_{j})/\Delta$ as $\lambda_{j}\circ r=r\circ\lambda_{k}$
$=\overline{r}(\lambda_{K}((a,0_{k})/\Delta))$ by definition of $\mbox{by definition of $\lambda_{K}$}.$
This proves $\Phi(\lambda)\in\overline{\mathbb{F}}$ and hence $\operatorname{ran}(\Phi)\subseteq\overline{\mathbb{F}}$. Again let $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$ and let $\overline{\lambda}=(\lambda_{J})_{J\in[I]}=\Phi(\lambda)\in\overline{\mathbb{F}}$. To show $\Psi(\overline{\lambda})=\lambda$, we must show $\overline{\lambda}_{0_{i}}=\lambda_{i}$ for each $i\in I$. Fix $J\in[I]$ and $i\in J$. We have $\lambda_{J}\circ\chi_{0_{i}}=\chi_{0_{i}}\circ\lambda_{i}$ by the definition of $\lambda_{J}$. But this means that $\lambda_{i}$ satisfies the equation which determines $\overline{\lambda}_{0_{i}}$ (see the discussion following Definition 6.11). This proves $\Psi(\overline{\lambda})=\lambda$ and hence $\Psi\circ\Phi=\mathrm{id}_{\mathbb{F}}$, which completes the proof of Claim 6.15.4 and of (1). ∎
(2) It is a tedious but mostly routine exercise to show that $\Psi$ is a unital ring homomorphism. For example, to prove that $\Psi$ preserves addition, it suffices to prove that if $\overline{\lambda},\overline{\mu},\overline{\tau}\in\overline{\mathbb{F}}$ with $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}$ etc. and such that $\overline{\tau}=\overline{\lambda}+\overline{\mu}$ calculated in $\overline{\mathbb{F}}$, then for all $e\in A$ and $a\in e/\theta$ we have $\overline{\tau}_{e}(a)=\overline{\lambda}_{e}(a)+\overline{\mu}_{e}(a)$ calculated in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. The following calculation proves this. Given $J\in[I]$, $e\in C_{J}$ and $a\in e/\theta$,
$\chi_{e}(\overline{\tau}_{e}(a))$ $=\overline{\tau}_{J}(\chi_{e}(a))$ $\chi_{e}\circ\overline{\tau}_{e}=\overline{\tau}_{J}\circ\chi_{e}$, by definition of $\overline{\tau}_{e}$
$=\overline{\lambda}_{J}(\chi_{e}(a))+\overline{\mu}_{J}(\chi_{e}(a))$ definition of $\overline{\lambda}+\overline{\mu}$ in $\overline{\mathbb{F}}$
$=\chi_{e}(\overline{\lambda}_{e}(a))+\chi_{e}(\overline{\mu}_{e}(a))$ definitions of $\overline{\lambda}_{e}$ and $\overline{\mu}_{e}$
$=\chi_{e}(\overline{\lambda}_{e}(a)+\overline{\mu}_{e}(a))$ as $\chi_{e}$ is a group homomorphism.
Since $\chi_{e}$ is injective, we get $\overline{\tau}_{e}(a)=\overline{\lambda}_{e}(a)+\overline{\tau}_{e}(a)$ as required. The proofs that $\Psi$ preserves multiplication and the unit are no harder. Thus $\Psi$ is a surjective unital ring homomorphism, and since $\overline{\mathbb{F}}$ is a division ring and $\mathbb{F}$ is nontrivial, $\Psi$ must be an isomorphism.
(3) Let $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$, $e\in A$, and $a\in e/\theta$. Let $i\in I$ so that $a,e\in B_{i}$ and let $J=[i]$. Let $\cdot$ denote the action of $\overline{\mathbb{F}}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and let $\cdot^{e}$ denote the action of $\mathbb{F}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. Let $b=d(a,e,0_{i})$, $x=\overline{\lambda}_{0_{i}}(b)$, and $c=d(x,0_{i},e)$. Thus by definition,
$(x,0_{i})/\Delta=\overline{\lambda}_{J}((b,0_{i})/\Delta)\quad\mbox{and}\quad\Psi(\overline{\lambda})\cdot^{e}a=c.$
On the other hand, Lemma 5.4(1) gives
$(a,e)/\Delta=(b,0_{i})/\Delta\quad\mbox{and}\quad(x,0_{i})/\Delta=(c,e)/\Delta.$
Thus $(c,e)/\Delta=\overline{\lambda}_{J}((a,e)/\Delta)$, which implies $\overline{\lambda}\cdot a=\overline{\lambda}_{e}(a)=c$ as required. ∎
In summary, we have proved the following.
Proposition 6.16.
Let ${\mathbf{A}}$ be an algebra in a variety with a weak difference term, let $\theta$ be an abelian minimal congruence of ${\mathbf{A}}$, let $T$ be a transversal for $\theta$, and let $\mathbb{F}$ be the division ring associated to ${\mathbf{A}},\theta,T$. Let $\alpha=(0:\theta)$, $\Delta=\Delta_{\theta,\alpha}$, $\varphi=\overline{\alpha}/\Delta$, $D^{o}=\{0_{C}\,:\,\text{$C\in A/\alpha$}\}$, and ${\mathbf{D}}=D({\mathbf{A}},\theta)$, and let $\overline{\mathbb{F}}$ be the division ring associated to ${\mathbf{D}},\varphi,D^{o}$. Then there exists an isomorphism $\Psi:\overline{\mathbb{F}}\cong\mathbb{F}$ such that for all $e\in A$, if ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ are the two left vector space structures defined in the discussions following Definition 6.3 and Lemma 6.12, then ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ are the same vector space modulo the identification of $\mathbb{F}$ with $\overline{\mathbb{F}}$ via $\Psi$.
7. Similarity
In a 1983 paper [5], Freese introduced a relation of similarity between subdirectly irreducible algebras in a congruence modular variety. If the monoliths of the algebras are nonabelian, then the algebras are similar if and only if they are isomorphic. If the monoliths are abelian, then the relation is more interesting. Freese’s original definition in the latter case, as well as the equivalent formulation found in [6], depend on a certain “uniformity” of the blocks of an abelian monolith $\mu$ contained in a common $(0:\mu)$-block, represented in part by Corollary 5.3. As we have claimed (and will prove in Section 8), this uniformity is lost in varieties with a weak difference term.
Happily, Freese gave two other equivalent formulations of the similarity relation in the abelian monolith case, one of which is based on the construction in Section 4. Both formulations generalize nicely to varieties with a weak difference term. We develop this generalization in this section.
The following definition, in the context of congruence modular varieties, is found in [6, Definition 10.9].
Definition 7.1.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}}\in\mathcal{V}$ is subdirectly irreducible with monolith $\mu$. The algebra $D({\mathbf{A}})$ is defined as follows:
  1. (1)
    If $\mu$ is nonabelian, then $D({\mathbf{A}})={\mathbf{A}}$.
  2. (2)
    If $\mu$ is abelian, then $D({\mathbf{A}})=D({\mathbf{A}},\mu)$ as defined in Definition 4.7.
Observe that, in the context of the previous definition, $D({\mathbf{A}})$ is a subdirectly irreducible algebra in $\mathsf{HS}({\mathbf{A}}^{2})$; furthermore, the monoliths of ${\mathbf{A}}$ and $D({\mathbf{A}})$ are either both abelian or both nonabelian.
Definition 7.2.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible. We say that ${\mathbf{A}}$ and ${\mathbf{B}}$ are similar, and write ${\mathbf{A}}\sim{\mathbf{B}}$, if $D({\mathbf{A}})\cong D({\mathbf{B}})$.
By [5, Theorem 5.1] and [6, Theorem 10.11], our definition of similarity, when restricted to the congruence modular setting, is equivalent to the original definition given in [5] and [6]. It is in this sense that our definition extends the original one.
If we let $\mu$ and $\kappa$ denote the respective monoliths of ${\mathbf{A}}$ and ${\mathbf{B}}$, then Definition 7.2 can be rephrased as follows: ${\mathbf{A}}\sim{\mathbf{B}}$ if and only if one of the following holds:
  1. (1)
    $\mu$ and $\kappa$ are both nonabelian and ${\mathbf{A}}\cong{\mathbf{B}}$; or
  2. (2)
    $\mu$ and $\kappa$ are both abelian and $D({\mathbf{A}},\mu)\cong D({\mathbf{B}},\kappa)$.
Before stating the main theorem of this section, we introduce a variant of a crucial construct due to Zhuk [37, Section 6.6].
Definition 7.3.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are subdirectly irreducible algebras in a common signature with monoliths $\mu,\kappa$ respectively. A proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$ is a subuniverse $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ satisfying
  1. (B1)
    $\operatorname{pr}_{1,2}(T)=\mu$ and $\operatorname{pr}_{3,4}(T)=\kappa$.
  2. (B2)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $a_{1}=a_{2}$ if and only if $b_{1}=b_{2}$.
  3. (B3)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $(a_{i},a_{i},b_{i},b_{i})\in T$ for $i=1,2$.
If $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$, then we define its trace to be the set $\operatorname{tr}(T)=\{(a,b)\in A\times B\,:\,\text{$(a,a,b,b)\in T$}\}$, and its kernel to be the set
$\ker(T)=\{((a_{1},b_{1}),(a_{2},b_{2}))\,:\,\text{$(a_{1},a_{2},b_{1},b_{2})\in T$}\}.$
Observe that for any proper bridge $T$ we have $\operatorname{tr}(T)=\operatorname{pr}_{1,3}(T)=\operatorname{pr}_{2,4}(T)$ by property (B3), and hence $\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ by property (B1).
In the special case where one of the two monoliths in the previous definition is nonabelian, the existence of a proper bridge is possible only in the most trivial way.
Proposition 7.4.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Assume that one of $\mu$ or $\kappa$ is nonabelian. Then so is the other, $\operatorname{tr}(T)$ is the graph of an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$, and $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$.
Proof.
Assume with no loss of generality that $\kappa$ is nonabelian. We will first show that $\operatorname{tr}(T)$ is the graph of a surjective homomorphism $h:{\mathbf{A}}\rightarrow{\mathbf{B}}$. Let
$S=\{(b,c)\in B^{2}\,:\,\text{$\exists a\in A$ with $(a,b),(a,c)\in\operatorname{tr}(T)$}\}.$
$S$ is a tolerance of ${\mathbf{B}}$. Let
$Y=\left\{\begin{pmatrix}b&c\\ b^{\prime}&c^{\prime}\end{pmatrix}\in B^{2\times 2}~{}:~{}\mbox{$\exists\,(a,a^{\prime})\in\mu$ with $(a,a^{\prime},b,b^{\prime}),(a,a^{\prime},c,c^{\prime})\in T$}\right\}.$
One can easily show, using (B1) and the definitions of $S$ and $Y$, that $X(\kappa,S)\subseteq Y$ (see Definition 2.1). Since $Y$ is a subuniverse of ${\mathbf{B}}^{2\times 2}$, we get $M(\kappa,S)\subseteq Y$. Then from (B2) and Definition 2.2(2) we can deduce $C(S,\kappa;0_{B})$. Since $S$ is a tolerance, we can improve this to $C(\operatorname{Cg}(S),\kappa;0_{B})$ by [17, Theorem 2.19(2)]. Now assume that $S\neq 0_{B}$; then $\operatorname{Cg}(S)\geq\kappa$ since $\kappa$ is the monolith of ${\mathbf{B}}$; but then $\kappa$ would be abelian, contrary to our assumption. This proves $S=0_{B}$ and hence $\operatorname{tr}(T)$ is the graph of a surjective homomorphism $h:{\mathbf{A}}\rightarrow{\mathbf{B}}$.
Assume that $h$ is not injective. Then $\mu\subseteq\ker(h)$ as $\mu$ is the monolith of ${\mathbf{A}}$. Choose $(a,b)\in\mu$ with $a\neq b$. By (B1), there exists $(c,d)\in\kappa$ with $(a,b,c,d)\in T$. Then $c\neq d$ by (B2). By (B3), we have $(a,c),(b,d)\in\operatorname{tr}(T)$, or equivalently, $c=h(a)$ and $d=h(b)$. But then $(a,b)\in\mu\subseteq\ker(h)$ implies $c=d$, contradiction. Thus $h$ is an isomorphism.
Finally, for every $(a,b)\in\mu$ there is at least one pair $(c,d)\in\kappa$ with $(a,b,c,d)\in T$ by (B1). On the other hand, the only pair $(c,d)$ that can work is $(c,d)=(h(a),h(b))$ by (B3) and the definition of $h$. Hence $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$. ∎
Lemma 7.5.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Let ${\mathbf{C}}=\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and $\tau=\ker(T)$. Then $\tau\in\operatorname{Con}({\mathbf{C}})$. Hence
  1. (B4)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $(a_{2},a_{1},b_{2},b_{1})\in T$.
  2. (B5)
    If $(a_{1},a_{2},b_{1},b_{2})\in T$ and $(a_{2},a_{3},b_{2},b_{3})\in T$, then $(a_{1},a_{3},b_{1},b_{3})\in T$.
Proof.
Note that $\tau$ is a reflexive subuniverse of ${\mathbf{C}}^{2}$ by (B3), and that (B4) and (B5) simply express the symmetry and transitivity of $\tau$. Hence all we need to prove is that $\tau$ is symmetric and transitive.
If $\mu$ or $\kappa$ is nonabelian, then $C$ is the graph of an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$ and $\tau=\{((a,h(a)),(b,h(b))\,:\,\text{$(a,b)\in\mu$}\}$ by Proposition 7.4. Clearly $\tau$ is symmetric and transitive in this case.
If $\mu$ and $\kappa$ are both abelian, let $d(x,y,z)$ be a weak difference term for $\mathcal{V}$. Let
$\theta=\{((a_{1},b_{1}),(a_{2},b_{2}))\in C^{2}\,:\,\text{$(a_{1},a_{2})\in\mu$ and $(b_{1},b_{2})\in\kappa$}\}.$
Clearly $\theta\in\operatorname{Con}({\mathbf{C}})$. Because $\mu$ and $\kappa$ are abelian, the restriction of $d$ to any $\theta$-class is a Maltsev operation. We have that $\tau$ is a reflexive subuniverse of ${\mathbf{C}}^{2}$ and $\tau\subseteq\theta$ by (B1). Thus $\tau$ is symmetric and transitive by Lemma 3.1. ∎
Lemma 7.6.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Let ${\mathbf{C}}=\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and $\tau=\ker(T)$. Also let $\delta_{1},\delta_{2}\in\operatorname{Con}({\mathbf{C}})$ be the kernels of the two projection homomorphisms $\operatorname{pr}_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $\operatorname{pr}_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and let $\delta^{+}_{i}$ be the unique upper cover of $\delta_{i}$ in $\operatorname{Con}({\mathbf{C}})$ for $i=1,2$. Then $(\delta_{1},\delta^{+}_{1})\searrow(0,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$.
Proof.
Clearly $\tau\leq\delta^{+}_{1}\wedge\delta^{+}_{2}$ by (B1). Choose $(a_{1},a_{2})\in\mu$ with $a_{1}\neq a_{2}$. Again by (B1), there exist $b_{1},b_{2}\in B$ with $(a_{1},a_{2},b_{1},b_{2})\in T$ and hence $(a_{1},b_{1})\stackrel{{\scriptstyle\tau}}{{\equiv}}(a_{2},b_{2})$, which proves $\tau\nleq\delta_{1}$. A similar argument shows $\tau\nleq\delta_{2}$. This proves $\tau\vee\delta_{i}=\delta^{+}_{i}$ for $i=1,2$.
Finally, suppose $((a_{1},b_{1}),(a_{2},b_{2}))\in\tau\cap\delta_{1}$. Then $a_{1}=a_{2}$, so $b_{1}=b_{2}$ by (B2), so $(a_{1},b_{1})=(a_{2},b_{2})$. This shows $\tau\wedge\delta_{1}=0$, and a similar proof shows $\tau\wedge\delta_{2}=0$. This proves $(\delta_{1},\delta^{+}_{1})\searrow(0,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$ as required. ∎
In the next lemma, we show that proper bridges between subdirectly irreducible algebras with abelian monoliths satisfy additional properties.
Lemma 7.7.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly algebras with abelian monoliths $\mu,\kappa$ respectively, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Let $\alpha=(0:\mu)$ and $\beta=(0:\kappa)$; also let $\Delta_{{\mathbf{A}}}=\Delta_{\mu,\alpha}$ and $\Delta_{{\mathbf{B}}}=\Delta_{\kappa,\beta}$.
  1. (1)
    For all $(a,r),(b,s)\in\operatorname{tr}(T)$, $(a,b)\in\alpha\iff(r,s)\in\beta$.
  2. (2)
    For every $(\mu,\alpha)$-matrix $M=\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}$ there exists a $(\kappa,\beta)$-matrix $\begin{pmatrix}r&r^{\prime}\\ s&s^{\prime}\end{pmatrix}$ with $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$.
  3. (3)
    For all $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$, $(a,b)\stackrel{{\scriptstyle\Delta_{{\mathbf{A}}}}}{{\equiv}}(a^{\prime},b^{\prime})\iff(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r^{\prime},s^{\prime})$.
Proof.
(1) Define ${\mathbf{C}}=\operatorname{tr}(T)\leq{\mathbf{A}}\times{\mathbf{B}}$ and $\delta_{1},\delta^{+}_{1},\delta_{2},\delta^{+}_{2},\tau\in\operatorname{Con}({\mathbf{C}})$ as in Lemma 7.5. Observe that $\delta^{+}_{i}/\delta_{i}$ is abelian for $i=1,2$, because $\mu$ and $\kappa$ are abelian. Since $(\delta_{1},\delta^{+}_{1})\searrow(0,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$ by Lemma 7.5, we get $(\delta_{1}:\delta^{+}_{1})=(0:\tau)=(\delta_{2}:\delta^{+}_{2})$ by Proposition 3.13(1). On the other hand, letting $\operatorname{pr}_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ be the first projection homomorphism, we have $(\delta_{1}:\delta^{+}_{1})=(\operatorname{pr}_{1}^{-1}(0):\operatorname{pr}_{1}^{-1}(\mu))=\operatorname{pr}_{1}^{-1}((0:\mu))=\operatorname{pr}_{1}^{-1}(\alpha)$ by Lemma 2.3(2), and similarly $(\delta_{2}:\delta^{+}_{2})=\operatorname{pr}_{2}^{-1}(\beta)$. Together these facts prove $\operatorname{pr}_{1}^{-1}(\alpha)=\operatorname{pr}_{2}^{-1}(\beta)$, which is equivalent to (1).
(2) Recall that the set $M(\mu,\alpha)$ of all $(\mu,\alpha)$-matrices is the subuniverse of ${\mathbf{A}}^{2\times 2}$ generated by the set
$X(\mu,\alpha):=\left\{\begin{pmatrix}a&a\\ b&b\end{pmatrix}:(a,b)\in\mu\right\}\cup\left\{\begin{pmatrix}a&a^{\prime}\\ a&a^{\prime}\end{pmatrix}:(a,a^{\prime})\in\alpha\right\}.$
Let $S$ be the set of pairs $(M,N)=\left(\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix},\begin{pmatrix}r&r^{\prime}\\ s&s^{\prime}\end{pmatrix}\right)$ where $M$ is a $(\mu,\alpha)$-matrix, $N$ is a $(\kappa,\beta)$-matrix, and $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$. Then $S\leq{\mathbf{A}}^{2\times 2}\times{\mathbf{B}}^{2\times 2}$ and $\operatorname{pr}_{1}(S)$ is a subuniverse of $M(\mu,\alpha)$. As we are desiring to prove $\operatorname{pr}_{1}(S)=M(\mu,\alpha)$, it is therefore enough to prove $X(\mu,\alpha)\subseteq\operatorname{pr}_{1}(S)$. That is, to prove (2), it is enough to prove it for matrices $M\in X(\mu,\alpha)$.
Consider first the case where $M=\begin{pmatrix}a&a\\ b&b\end{pmatrix}$ with $(a,b)\in\mu$. By (B1), there exists $(r,s)\in\kappa$ with $(a,b,r,s)\in T$. Then $N:=\begin{pmatrix}r&r\\ s&s\end{pmatrix}\in M(\kappa,\beta)$ and $(M,N)\in S$ witnessed by $(a,b,r,s)$ and $(a^{\prime},b^{\prime},r^{\prime},s^{\prime}):=(a,b,r,s)$.
Consider next the case where $M=\begin{pmatrix}a&a^{\prime}\\ a&a^{\prime}\end{pmatrix}$ with $(a,a^{\prime})\in\alpha$. Pick $r,r^{\prime}\in B$ with $(a,r),(a^{\prime},r^{\prime})\in\operatorname{tr}(T)$. Then $(r,r^{\prime})\in\beta$ by (1), so $N:=\begin{pmatrix}r&r^{\prime}\\ r&r^{\prime}\end{pmatrix}\in M(\kappa,\beta)$, and we have $(M,N)\in S$ witnessed by $(a,a,r,r)$ and $(a^{\prime},a^{\prime},r^{\prime},r^{\prime})$. This completes the proof of (2).
We start the proof of (3) by proving the following special case.
Claim 7.7.1.
For all $(a,b)\in\mu$ and $(r,s),(u,v)\in\kappa$,
$(a,b,r,s),(a,b,u,v)\in T\implies(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(u,v).$
Proof of Claim 7.7.1.
Assume $(a,b,r,s),(a,b,u,v)\in T$. Then $(a,r),(a,u)\in\operatorname{tr}(T)$, so $(r,u)\in\beta$ by (1). Since $(r,s),(u,v)\in\kappa\leq\beta$, it follows that $r,s,u,v$ belong to a common $\beta$-class.
Let $d(x,y,z)$ be a weak difference term for $\mathcal{V}$. We claim that
($\ast$) $d(s,u,u)=d(s,v,v)\quad\mbox{and}\quad d(r,u,u)=d(s,v,u).$
To prove the first equality, observe that $d(\underline{u},u,u)=u=d(\underline{u},v,v)$ since $u,v$ belong to a common class of an abelian congruence ($\kappa$). Since $(u,s)\in\beta$ and $C(\beta,\kappa;0)$, we can replace the underlined occurrences of $u$ with $s$ to get $d(\underline{s},u,u)=d(\underline{s},v,v)$.
To prove the second equality, observe that the following three 4-tuples belong to $T$ (the third because of (B3)):
$(a,b,r,s),\quad(a,b,u,v),\quad(a,a,u,u).$
Applying the weak difference term coordinate-wise to these tuples gives
$(a,d(b,b,a),d(r,u,u),d(s,v,u))\in T.$
As $a,b$ belong to a common class of an abelian congruence ($\mu$), we get $d(b,b,a)=a$. Thus by bridge property (B2) we get $d(r,u,u)=d(s,v,u)$.
Now we build some $(\kappa,\beta)$-matrices. Let
$M_{1}=\begin{pmatrix}r&r\\ s&s\end{pmatrix},~{}~{}M_{2}=\begin{pmatrix}r&u\\ r&u\end{pmatrix},~{}~{}M_{3}=\begin{pmatrix}s&v\\ s&v\end{pmatrix},~{}~{}M_{4}=\begin{pmatrix}v&v\\ v&v\end{pmatrix},~{}~{}M_{5}=\begin{pmatrix}u&u\\ v&v\end{pmatrix}.$
Clearly $M_{1},\ldots,M_{5}\in M(\kappa,\beta)$. Applying the weak difference term gives
$N:=d(M_{1},M_{2},M_{2})$ $=\begin{pmatrix}r&d(r,u,u)\\ s&d(s,u,u)\end{pmatrix}\in M(\kappa,\beta)$
$N^{\prime}:=d(M_{3},M_{4},M_{5})$ $=\begin{pmatrix}d(s,v,u)&u\\ d(s,v,v)&v\end{pmatrix}\in M(\kappa,\beta).$
The second column of $N$ and the first column of $N^{\prime}$ are identical by ($\ast$7). Since $\Delta_{{\mathbf{B}}}$ is the “horizontal transitive closure” of $M(\kappa,\beta)$, it follows that $(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(u,v)$ as required. ∎
Now we prove (3). Fix $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$. Assume $(a,b)\stackrel{{\scriptstyle\Delta_{{\mathbf{A}}}}}{{\equiv}}(a^{\prime},b^{\prime})$. Then there exist matrices
$\begin{pmatrix}a_{0}&a_{1}\\ b_{0}&b_{1}\end{pmatrix},\begin{pmatrix}a_{1}&a_{2}\\ b_{1}&b_{2}\end{pmatrix},\ldots,\begin{pmatrix}a_{n-1}&a_{n}\\ b_{n-1}&b_{n}\end{pmatrix}\in M(\mu,\alpha)$
with $(a_{0},b_{0})=(a,b)$ and $(a_{n},b_{n})=(a^{\prime},b^{\prime})$. For each $i<n$, apply item (2) to the matrix $\begin{pmatrix}a_{i}&a_{i+1}\\ b_{i}&b_{i+1}\end{pmatrix}$ to get a matrix
$\begin{pmatrix}r_{i}&r_{i}^{\prime}\\ s_{i}&s_{i}^{\prime}\end{pmatrix}\in M(\kappa,\beta)$
with
$(a_{i},b_{i},r_{i},s_{i})\in T\quad\mbox{and}\quad(a_{i+1},b_{i+1},r_{i}^{\prime},s_{i}^{\prime})\in T.$
Then for each $i<n$ we have $(r_{i},s_{i})\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r_{i}^{\prime},s_{i}^{\prime})$ since these pairs are the two columns of a $(\kappa,\beta)$-matrix and $\Delta_{{\mathbf{B}}}$ is the transitive closure of $M_{\kappa,\beta}$. By Claim 7.7.1 we also get
$(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r_{0},s_{0}),\quad\mbox{$(r_{i}^{\prime},s_{i}^{\prime})\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r_{i+1},s_{i+1})$ for all $i<n$},\quad\mbox{and}\quad(r_{n}^{\prime},s_{n}^{\prime})\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r^{\prime},s^{\prime}).$
Then $(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r^{\prime},s^{\prime})$ by transitivity. The converse implication is proved similarly. ∎
Here is the main theorem of this section. The equivalence of items (1)–(3) was proved for congruence modular varieties by Freese [5, Theorems 2.1 and 5.1]; see also [6, Theorems 10.8 and 10.11].
Theorem 7.8.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible. The following are equivalent:
  1. (1)
    ${\mathbf{A}}\sim{\mathbf{B}}$.
  2. (2)
    There exist an algebra ${\mathbf{C}}\in\mathcal{V}$, surjective homomorphisms $f_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and congruences $\psi,\tau\in\operatorname{Con}({\mathbf{C}})$ with $\psi<\tau$, such that, letting $\delta_{i}=\ker(f_{i})$ and letting $\delta^{+}_{i}$ denote the unique upper cover of $\delta_{i}$ in $\operatorname{Con}({\mathbf{C}})$ for $i=1,2$, we have $(\delta_{1},\delta^{+}_{1})\searrow(\psi,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$.
  3. (3)
    There exist an algebra ${\mathbf{C}}\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and a congruence $\tau\in\operatorname{Con}({\mathbf{C}})$ which, with the projection homomorphisms $f_{1}=\operatorname{pr}_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}=\operatorname{pr}_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$ and $\psi:=0_{C}$, satisfy the conclusion of item (2).
  4. (4)
    There exists a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$.
Proof.
Throughout this proof, let $\mu,\kappa$ denote the monoliths of ${\mathbf{A}},{\mathbf{B}}$ respectively. The implication (3) $\Rightarrow$ (2) is clear, and (4) $\Rightarrow$ (3) holds by Lemma 7.6. We will first prove (2) $\Rightarrow$ (1) $\Rightarrow$ (4) in the case when one of $\mu$ or $\kappa$ is nonabelian. Then we will prove (2) $\Rightarrow$ (4) $\Leftrightarrow$ (1) in the case when both $\mu$ and $\kappa$ are abelian.
Case 1: $\mu$ or $\kappa$ is nonabelian.
(1) $\Rightarrow$ (4). Since ${\mathbf{A}}\sim{\mathbf{B}}$, it follows that both $\mu$ and $\kappa$ are nonabelian and there exists an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$. Then the set $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$.
(2) $\Rightarrow$ (1). Assume there exist ${\mathbf{C}}\in\mathcal{V}$, surjective homomorphisms $f_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and congruences $\psi,\tau\in\operatorname{Con}({\mathbf{C}})$ with $\psi<\tau$, such that, letting $\delta_{i}=\ker(f_{i})$ and letting $\delta^{+}_{i}$ denote the unique upper cover of $\delta_{i}$ for $i=1,2$, we have $(\delta_{1},\delta^{+}_{1})\searrow(\psi,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$. Because we are in Case 1, at least one of $\delta^{+}_{1}/\delta_{1}$ or $\delta^{+}_{2}/\delta_{2}$ is nonabelian. By Proposition 3.12(1), both are nonabelian, as is $\tau/\psi$. It follows that $\mu$ and $\kappa$ are both nonabelian.
For each $i=1,2$ we have $\delta_{i}\wedge\tau=\psi$ and hence $C(\delta_{i},\tau;\psi)$; see e.g. [17, Theorem 2.19(8)]; thus $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ by semidistributivity of the centralizer relation in the first variable; see e.g. [17, Theorem 2.19(5)]. Suppose $\delta_{1}\neq\delta_{2}$; then without loss of generality we can assume $\delta_{1}\vee\delta_{2}>\delta_{2}$. Hence $\delta_{1}\vee\delta_{2}\geq\delta^{+}_{2}$ (as $\delta_{2}$ is completely meet-irreducible) and thus $\delta_{1}\vee\delta_{2}\geq\tau$. Then $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ implies $C(\tau,\tau;\psi)$ by monotonicity, which contradicts the fact that $\tau/\psi$ is nonabelian. This proves $\delta_{1}=\delta_{2}$ and hence ${\mathbf{A}}\cong{\mathbf{C}}/\delta_{1}={\mathbf{C}}/\delta_{2}\cong{\mathbf{B}}$. Hence ${\mathbf{A}}\sim{\mathbf{B}}$.
Case 2: $\mu$ and $\kappa$ are abelian.
Thus $D({\mathbf{A}})=D({\mathbf{A}},\mu)$ and $D({\mathbf{B}})=D({\mathbf{B}},\kappa)$. For the remainder of this proof, let $\alpha=(0:\mu)$, $\Delta_{{\mathbf{A}}}=\Delta_{\mu,\alpha}$, and $\varphi_{{\mathbf{A}}}=\overline{\alpha}_{{\mathbf{A}}}/\Delta_{{\mathbf{A}}}$. Also let $D^{o}_{{\mathbf{A}}}\leq D({\mathbf{A}})$ as given in the proof of Corollary 4.8 for $({\mathbf{A}},\mu)$. Similarly define $\beta:=(0:\kappa)$, $\Delta_{{\mathbf{B}}}$, $\varphi_{{\mathbf{B}}}$, and $D^{o}_{{\mathbf{B}}}$ for $({\mathbf{B}},\kappa)$.
(4) $\Rightarrow$ (1). Assume that $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. By Lemma 7.7(3), the set
$\{((a,b)/\Delta_{{\mathbf{A}}},(r,s)/\Delta_{{\mathbf{B}}})\,:\,\text{$(a,b,r,s)\in T$}\}$
is the graph of a bijection $\lambda$. By the nature of its definition it is also a subalgebra of $D({\mathbf{A}})\times D({\mathbf{B}})$, and by (B1) the domain and range of $\lambda$ are $D({\mathbf{A}})$ and $D({\mathbf{B}})$ respectively. Thus $\lambda$ is an isomorphism from $D({\mathbf{A}})$ to $D({\mathbf{B}})$.
(1) $\Rightarrow$ (4). Assume $\lambda_{0}:D({\mathbf{A}})\cong D({\mathbf{B}})$. Recall that $D({\mathbf{A}})$ and $D({\mathbf{B}})$ are subdirectly irreducible with monoliths $\varphi_{{\mathbf{A}}},\varphi_{{\mathbf{B}}}$ respectively, by Corollary 4.8. Thus we must have $\lambda_{0}(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$. As $D^{o}_{{\mathbf{A}}}$ is a subuniverse of $D({\mathbf{A}})$ and a transversal for $\varphi_{{\mathbf{A}}}$, it follows that $\lambda_{0}(D^{o}_{{\mathbf{A}}})$, like $D^{o}_{{\mathbf{B}}}$, is a subuniverse of $D({\mathbf{B}})$ and a transversal for $\varphi_{{\mathbf{B}}}$. Thus by Lemma 3.9, there exists an isomorphism $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$ satisfying $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$. Now define
[…]
Because $\lambda$ is an isomorphism, we get $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ with $\operatorname{pr}_{1,2}(T)=\mu$ and $\operatorname{pr}_{3,4}(T)=\kappa$, proving (B1). Let $(a_{1},a_{2},b_{1},b_{2})\in T$ and assume $a_{1}=a_{2}$. Then $(a_{1},a_{2})/\Delta_{{\mathbf{A}}}\in D^{o}_{{\mathbf{A}}}$, As $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$, we get $(b_{1},b_{2})/\Delta_{{\mathbf{B}}}=\lambda((a_{1},a_{2})/\Delta_{{\mathbf{A}}})\in D^{o}_{{\mathbf{B}}}$, so $b_{1}=b_{2}$ (see Theorem 4.6(1)). A similar argument shows that if $b_{1}=b_{2}$, then $a_{1}=a_{2}$. This proves (B2).
Again suppose $(a_{1},a_{2},b_{1},b_{2})\in T$ and let $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}$. Observe that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{A}}}}}{{\equiv}}(a_{1},a_{2})/\Delta_{{\mathbf{A}}}$ and $(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}(b_{1},b_{2})/\Delta_{{\mathbf{B}}}$. By an earlier observation we have $\lambda(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$. Thus
$r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ $\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}\lambda((a_{1},a_{2})/\Delta_{{\mathbf{A}}})$
$=(b_{1},b_{2})/\Delta_{{\mathbf{B}}}$ as $(a_{1},a_{2},b_{1},b_{2})\in T$
$\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}(b_{1},b_{1})/\Delta_{{\mathbf{B}}}=s,$
i.e., $(r,s)\in\varphi_{{\mathbf{B}}}$. Also recall that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\in D^{o}_{{\mathbf{A}}}$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\in D^{o}_{{\mathbf{B}}}$, and since $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$ we also get $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})\in D^{o}_{{\mathbf{B}}}$. So $r,s\in D^{o}_{{\mathbf{B}}}$ and $(r,s)\in\varphi_{{\mathbf{B}}}$. But $D^{o}_{{\mathbf{B}}}$ is a transversal for $\varphi_{{\mathbf{B}}}$. These facts imply $r=s$, which implies $(a_{1},a_{1},b_{1},b_{1})\in T$. A similar argument gives $(a_{2},a_{2},b_{2},b_{2})\in T$. This proves (B3). Thus $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$.
(2) $\Rightarrow$ (4). Let ${\mathbf{C}}\in\mathcal{V}$, homomorphisms $f_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and congruences $\delta_{1},\delta_{2},\delta^{+}_{1},\delta^{+}_{2},\psi,\tau$ satisfy the conditions in (2). Because $(\delta_{i},\delta^{+}_{i})\searrow(\psi,\tau)$, we get $\delta^{+}_{i}=\delta_{i}\circ\tau\circ\delta_{i}$ for $i=1,2$ and $(\delta_{1}:\delta^{+}_{1})=(\delta_{2}:\delta^{+}_{2})$ by Proposition 3.13.
Claim 7.8.1.
$f_{1}(\tau)=\mu$, $f_{2}(\tau)=\kappa$, and $f_{1}^{-1}(\alpha)=f_{2}^{-1}(\beta)$.
Proof of Claim 7.8.1.
We have $f_{1}(\tau)\subseteq\mu$ because $\tau\leq\delta^{+}_{1}$. Given $(a,b)\in\mu$, pick $c\in f_{1}^{-1}(a)$ and $d\in f_{1}^{-1}(b)$. Then $(c,d)\in\delta^{+}_{1}$, so there exist $c^{\prime},d^{\prime}\in C$ with
$c\stackrel{{\scriptstyle\delta_{1}}}{{\equiv}}c^{\prime}\stackrel{{\scriptstyle\tau}}{{\equiv}}d^{\prime}\stackrel{{\scriptstyle\delta_{1}}}{{\equiv}}d.$
Then $(a,b)=(f_{1}(c^{\prime}),f_{1}(d^{\prime}))\in f_{1}(\tau)$, proving $\mu\subseteq f_{1}(\tau)$. Thus $f_{1}(\tau)=\mu$, and a similar proof shows $f_{2}(\tau)=\kappa$. Finally, observe that
$f_{1}^{-1}(\alpha)=f_{1}^{-1}((0:\mu_{{\mathbf{A}}}))=(f_{1}^{-1}(0):f_{1}^{-1}(\mu_{{\mathbf{A}}}))=(\delta_{1}:\delta^{+}_{1}).$
A similar calculation shows $f_{2}^{-1}(\beta)=(\delta_{2}:\delta^{+}_{2})$. Since $(\delta_{1}:\delta^{+}_{1})=(\delta_{2}:\delta^{+}_{2}$), we get $f_{1}^{-1}(\alpha)=f_{2}^{-1}(\beta)$. ∎
Now define
$T=\{(f_{1}(c),f_{1}(d),f_{2}(c),f_{2}(d))\,:\,\text{$(c,d)\in\tau$}\}.$
We will show that $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Clearly $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$, and $\operatorname{pr}_{1,2}(T)=\mu$ and $\operatorname{pr}_{3,4}(\tau)=\kappa$ by Claim 7.8.1, proving (B1). $T$ clearly satisfies (B3), as $\tau$ is reflexive. To prove (B2), let $(a_{1},a_{2},b_{1},b_{2})\in T$ and pick $(c,d)\in\tau$ witnessing this. If $a_{1}=a_{2}$, then $(c,d)\in\ker(f_{1})=\delta_{1}$, so $(c,d)\in\tau\wedge\delta_{1}=\psi\leq\delta_{2}=\ker(f_{2})$, proving $b_{1}=b_{2}$. The opposite implication is proved similarly. ∎
Item (1) in the next result extends the first claim of [6, Theorem 10.10] from congruence modular varieties to varieties with a weak difference term. (The remainder of [6, Theorem 10.10] was extended by Corollary 4.8.) Item (3) extends a fact established, in the congruence modular setting, within the proof of [6, Theorem 10.11].
Corollary 7.9.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}}\in\mathcal{V}$ is subdirectly irreducible with abelian monolith $\mu$.
  1. (1)
    ${\mathbf{A}}\sim D({\mathbf{A}})$.
  2. (2)
    Setting $\alpha=(0:\mu)$ and $\Delta=\Delta_{\mu,\alpha}$, the set
    $T_{{\mathbf{A}}}^{D}:=\{(a,b,(a,e)/\Delta,(b,e)/\Delta)\,:\,\text{$a\stackrel{{\scriptstyle\mu}}{{\equiv}}b\stackrel{{\scriptstyle\mu}}{{\equiv}}e$}\}$
    is a proper bridge from ${\mathbf{A}}$ to $D({\mathbf{A}})$.
  3. (3)
    If $(0:\mu)=\mu$ and there exists a subuniverse $A_{0}\leq{\mathbf{A}}$ which is a transversal for $\mu$, then ${\mathbf{A}}\cong D({\mathbf{A}})$.
Hence each similarity class in $\mathcal{V}$ consisting of subdirectly irreducible algebras with abelian monoliths contains a unique (up to isomorphism) subdirectly irreducible algebra ${\mathbf{S}}$ with abelian monolith $\varphi$ satisfying (i) $(0:\varphi)=\varphi$, and (ii) ${\mathbf{S}}$ has a subuniverse which is a transversal for $\varphi$.
Proof.
(1) Let $\eta_{1},\eta_{2}$ be the kernels of the two projection maps $\operatorname{pr}_{1},\operatorname{pr}_{2}:{\mathbf{A}}(\mu)\rightarrow{\mathbf{A}}$. By Theorem 4.6, if we set ${\mathbf{B}}=D({\mathbf{A}})$, ${\mathbf{C}}={\mathbf{A}}(\mu)$, $\delta_{1}=\eta_{1}$, $\delta_{2}=\Delta$, $\delta^{+}_{1}=\overline{\mu}$, $\delta^{+}_{2}=\overline{\alpha}$, $\psi=0$, and $\tau=\eta_{2}$, then the conditions of Theorem 7.8(2) are met, showing ${\mathbf{A}}\sim D({\mathbf{A}})$.
(2) This can be deduced from the above proof of item (1), following the proof of Theorem 7.8 (2) $\Rightarrow$ (4) in the abelian monolith case. Instead, we give a direct proof. Let $\varphi=\overline{\alpha}/\Delta$. Clearly $\operatorname{pr}_{1,2}(T_{{\mathbf{A}}}^{D})=\mu$, $\operatorname{pr}_{3,4}(T_{{\mathbf{A}}}^{D})\subseteq\varphi$, (B3) holds, and the ($\Rightarrow$) direction of (B2) holds. To prove the ($\Leftarrow$) direction of (B2), assume that $a,b,e$ belong to a common $\mu$ class and $(a,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e)$. Using the notation of Definition 5.1, we can express this as $\chi_{e}(a)=\chi_{e}(b)$. But $\chi_{e}$ is injective by Lemma 5.2, so $a=b$, completing the proof of (B2).
It remains to prove $\varphi\subseteq\operatorname{pr}_{3,4}(T_{{\mathbf{A}}}^{D})$. Let $((a_{1},a_{2})/\Delta,(b_{1},b_{2})/\Delta)\in\varphi$ be given; thus $(a_{1},a_{2}),(b_{1},b_{2})\in\mu$ and $a_{1},a_{2},b_{1},b_{2}$ belong to a common $\alpha$-class. By Lemma 5.7, there exists $e\in a_{1}/\alpha$ with $\operatorname{ran}(\chi_{a_{2}})\cup\operatorname{ran}(\chi_{b_{2}})\subseteq\operatorname{ran}(\chi_{e})$. As $(a_{1},a_{2})/\Delta\in\operatorname{ran}(\chi_{a_{2}})$ and $(b_{1},b_{2})/\Delta\in\operatorname{ran}(\chi_{b_{2}})$, this implies the existence of $a,b\in e/\mu$ with $\chi_{e}(a)=(a_{1},a_{2})/\Delta$ and $\chi_{e}(b)=(b_{1},b_{2})/\Delta$, which means
$((a_{1},a_{2})/\Delta,(b_{1},b_{2})/\Delta)=((a,e)/\Delta,(b,e)/\Delta)\in T_{{\mathbf{A}}}^{D}.$
(3) Since $(0:\mu)=\mu$, we have $D({\mathbf{A}})={\mathbf{A}}(\mu)/\Delta$ where $\Delta=\Delta_{\mu,\mu}$. Since $\mu$ is abelian, we get $\Delta=\left\{((a,b),(a^{\prime},b^{\prime})):\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in M(\mu,\mu)\right\}$ by Lemma 4.4(3). Let $\pi:{\mathbf{A}}\rightarrow{\mathbf{A}}_{0}$ be the homomorphism with $\pi{\upharpoonright}_{A_{0}}=\mathrm{id}_{A_{0}}$ and $\ker(\pi)=\mu$. Then define $\lambda:{\mathbf{A}}\rightarrow D({\mathbf{A}})$ by $\lambda(x)=(x,\pi(x))/\Delta$. Suppose $(a,b)\in\ker(\lambda)$. Then $((a,\pi(a),(b,\pi(b)))\in\Delta$, so $M:=\begin{pmatrix}a&b\\ \pi(a)&\pi(b)\end{pmatrix}\in M(\mu,\mu)$. Also $(a,b)\in\mu$, so $\pi(a)=\pi(b)$. Thus the bottom entries of $M$ are equal. Since $\mu$ is abelian, it follows that the top entries of $M$ are also equal, i.e., $a=b$. Thus $\lambda$ is injective.
To prove that $\lambda$ is surjective, let $(a,b)\in\mu$ be given and define $e=\pi(a)=\pi(b)$ and $c=d(a,b,e)$ where $d(x,y,z)$ is a weak difference term for $\mathcal{V}$. Then $(a,b)/\Delta=(c,e)/\Delta=\lambda(c)$. Thus $\lambda$ is an isomorphism from ${\mathbf{A}}$ to $D({\mathbf{A}})$. ∎
Recognizing when two subdirectly irreducible algebras with abelian monoliths are similar is not always easy. Our final result of this section gives some “obvious” necessary conditions.
Proposition 7.10.
Suppose $\mathcal{V}$ is a variety with a weak difference term and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with abelian monoliths $\mu,\kappa$ respectively. Define $\alpha=(0:\mu)$ and $\beta=(0:\kappa)$, let $\mathbb{F}_{{\mathbf{A}}}$ be the division ring associated to ${\mathbf{A}},\mu$ and some transversal for $\mu$, and let $\mathbb{F}_{{\mathbf{B}}}$ be the division ring associated to ${\mathbf{B}},\kappa$ and some transversal for $\kappa$. If ${\mathbf{A}}\sim{\mathbf{B}}$, then:
  1. (1)
    ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
  2. (2)
    $\mathbb{F}_{{\mathbf{A}}}\cong\mathbb{F}_{{\mathbf{B}}}$.
  3. (3)
    If ${\mathbf{A}}$ and ${\mathbf{B}}$ are finite, then there exists an isomorphism $f:{\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$ such that for all $C\in A/\alpha$, $\max\{|e/\mu|\,:\,\text{$e\in C$}\}=\max\{|e^{\prime}/\kappa|\,:\,\text{$e^{\prime}\in f(C)$}\}$.
Proof.
Let $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$. Let $\varphi_{{\mathbf{A}}},\varphi_{{\mathbf{B}}}$ be the monoliths of $D({\mathbf{A}}),D({\mathbf{B}})$ respectively. Then $\lambda$ induces an isomorphism $\lambda^{\ast}:D({\mathbf{A}})/\varphi_{{\mathbf{A}}}\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ which moreover satisfies $|E|=|\lambda^{\ast}(E)|$ for every $\varphi_{{\mathbf{A}}}$-class $E$.
(1) ${\mathbf{A}}/\alpha\cong D({\mathbf{A}})/\varphi_{{\mathbf{A}}}$ and ${\mathbf{B}}/\beta\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ by Corollary 4.8, so ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
(2) As in the proof of Theorem 7.8(1) $\Rightarrow$ (4) in the abelian monolith case, we can assume $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$. This implies that the division rings associated to $D({\mathbf{A}}),\varphi_{{\mathbf{A}}},D^{o}_{{\mathbf{A}}}$ and to $D({\mathbf{B}}),\varphi_{{\mathbf{B}}},D^{o}_{{\mathbf{B}}}$ are isomorphic. By Proposition 6.16, these division rings in turn are isomorphic to $\mathbb{F}_{{\mathbf{A}}}$ and $\mathbb{F}_{{\mathbf{B}}}$ respectively.
(3) Let $h_{{\mathbf{A}}}^{\ast}:{\mathbf{A}}/\alpha\cong D({\mathbf{A}})/\varphi_{{\mathbf{A}}}$ and $h_{{\mathbf{B}}}^{\ast}:{\mathbf{B}}/\beta\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ be given by Corollary 4.8 Then for each $C\in A/\alpha$, $h_{{\mathbf{A}}}^{\ast}(C)=0_{C}/\varphi_{{\mathbf{A}}}$, which is the universe of $\operatorname{Grp}_{D({\mathbf{A}})}(\varphi_{{\mathbf{A}}},0_{C})$. Thus by Lemma 5.2(1) and Corollary 5.8(2),
$\max\{|e/\mu|\,:\,\text{$e\in C$}\}=|0_{C}/\varphi_{{\mathbf{A}}}|=|h^{\ast}_{{\mathbf{A}}}(C)|.$
Similarly, for any $\beta$-class $C^{\prime}$ we have
$\max\{|e^{\prime}/\kappa|\,:\,\text{$e^{\prime}\in C^{\prime}$}\}=|h^{\ast}_{{\mathbf{B}}}(C^{\prime})|.$
If $C^{\prime}\in B/\beta$ is chosen so that $h_{{\mathbf{B}}}^{\ast}(C^{\prime})=\lambda^{\ast}(h_{{\mathbf{A}}}^{\ast}(C))$, then $|h^{\ast}_{{\mathbf{B}}}(C^{\prime})|=|h^{\ast}_{{\mathbf{A}}}(C)|$ by a comment in the preamble to this proof. Thus we can set $f=(h^{\ast}_{{\mathbf{B}}})^{-1}\circ\lambda^{\ast}\circ h^{\ast}_{{\mathbf{A}}}$. ∎
8. Examples
In this section we illustrate some of our earlier results by presenting a recipe for constructing finite subdirectly irreducible algebras with abelian monoliths in a varieties having weak difference terms. The algebras ${\mathbf{A}}$ can be constructed so that if $\mu$ is its monolith and $\alpha=(0:\mu)$, then
  1. (1)
    the division ring associated to ${\mathbf{A}},\mu$ and a transversal for $\mu$ can be any desired finite field $\mathbb{F}$;
  2. (2)
    $|A/\alpha|$ can be any positive integer.
  3. (3)
    The vector spaces ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{D({\mathbf{A}},\mu)}(\varphi,0_{C})$ as $C$ ranges over the $\alpha$-classes can have any finite dimensions.
  4. (4)
    The ranges of the $\mu$-classes within a given $\alpha$-class $C$ can be essentially any collection of subspaces of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}({\mathbf{A}},\mu)}(\varphi,0_{C})$ including $0_{C}/\varphi$ itself.
Our recipe is an adaptation of a known construction [22] (see also [3]). We first give a useful method for constructing varieties with a prescribed weak difference term. For this, we need the following result characterizing weak difference terms at the level of varieties. Following [19], given a variety $\mathcal{V}$, let ${\mathbf{F}}={\mathbf{F}}_{\mathcal{V}}(3)$ be the $\mathcal{V}$-free algebra freely generated by $\{x,y,z\}$, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ calculated in ${\mathbf{F}}$, and define $\beta_{0}=\beta$, $\gamma_{0}=\gamma$, $\beta_{n+1}=\beta\vee(\alpha\wedge\gamma_{n})$, and $\gamma_{n+1}=\gamma\vee(\alpha\wedge\beta_{n})$ for $n\geq 0$.
Theorem 8.1.
There exists an infinite sequence $\Sigma_{0},\Sigma_{1},\Sigma_{2},\ldots$ of finite sets of identities involving variables $x,y$ and ternary operation symbols $\mathsf{d},\mathsf{f}_{1},\mathsf{f}_{2},\ldots$, such that for every variety $\mathcal{V}$ and every ternary term $d(x,y,z)$ in the signature of $\mathcal{V}$, the following are equivalent:
  1. (1)
    $d$ is a weak difference term for $\mathcal{V}$.
  2. (2)
    For some $n\geq 0$ there exist idempotent ternary terms $f_{1},f_{2},\ldots$ in the signature of $\mathcal{V}$ so that, when the symbols $\mathsf{d},\mathsf{f}_{1},\mathsf{f}_{2},\ldots$ are interpreted by $d,f_{1},f_{2},\ldots$ respectively, the identities in $\Sigma_{n}$ are true in $\mathcal{V}$.
  3. (3)
    Letting ${\mathbf{F}}={\mathbf{F}}_{\mathcal{V}}(3)$ and interpreting $d$ as an element of ${\mathbf{F}}$ in the usual way, i.e. as $d^{{\mathbf{F}}}(x,y,z)$, we have $(x,d)\in\gamma_{m}$ and $(d,z)\in\beta_{m}$ for some $m\geq 0$.
Proof sketch.
(1) $\Rightarrow$ (3) can be extracted from the proof of Theorem 4.8 (1) $\Rightarrow$ (2) $\Rightarrow$ (3) in [19]. Standard arguments with roots in [29, 34] can produce sets $\Sigma_{0},\Sigma_{1},\Sigma_{2},\ldots$ of identities witnessing (2) $\Leftrightarrow$ (3). To prove (3) $\Rightarrow$ (1), assume that $(x,d)\in\gamma_{m}$ and $(d,z)\in\beta_{m}$. Then by [35, Theorem 2.1 (3) $\Rightarrow$ (4)], there exist idempotent ternary terms $f_{1},\ldots,f_{n},g_{1},\ldots,g_{n}$ (for some $n$) such that $\mathcal{V}\models f_{i}(x,y,x)\approx g_{i}(x,y,x)$ for each $i\in[n]$, and for all ${\mathbf{A}}\in\mathcal{V}$ and all $a,b\in A$, if
[…]
then $d(a,a,b)=b=d(b,a,a)$ and $d(b,b,a)=a=d(a,b,b)$. Now one can easily adapt the proof of [14, Theorem 1.2 (3) $\Rightarrow$ (1)] to get that $d$ is a weak difference term for $\mathcal{V}$. ∎
Definition 8.2.
A semilattice-over-Maltsev operation is any operation $d:A^{3}\rightarrow A$ constructed in the following way. Let $(S,\wedge)$ be a (meet) semilattice. For each $s\in S$ let $V_{s}$ be a non-empty set, and assume that $V_{s}\cap V_{t}=\varnothing$ for all $s,t\in S$ with $s\neq t$. For each pair $s,t\in S$ with $s\geq t$, assign a function $f_{(s,t)}:V_{s}\rightarrow V_{t}$, so that $f_{(s,s)}=\mathrm{id}_{V_{s}}$ and $f_{(s,u)}=f_{(t,u)}\circ f_{(s,t)}$ for all $s,t,u\in S$ with $s\geq t\geq u$. Finally, for each $s\in S$ let $m_{s}:(V_{s})^{3}\rightarrow V_{s}$ be a Maltsev operation.
From this data, define $A=\bigcup_{s\in S}V_{s}$ and $\nu:A\rightarrow S$ so that $a\in V_{\nu(a)}$ for all $a\in A$. Then define $d:A^{3}\rightarrow A$ by
$d(a,b,c)=m_{t}(f_{(\nu(a),t)}(a),f_{(\nu(b),t)}(b),f_{(\nu(c),t)}(c))\quad\mbox{where $t=\nu(a)\wedge\nu(b)\wedge\nu(c)$}.$
Lemma 8.3.
Suppose ${\mathbf{A}}$ is an algebra and $d$ is a ternary term operation of ${\mathbf{A}}$ which is semilattice-over-Maltsev. Then $d$ is a weak difference term for $\mathsf{HSP}({\mathbf{A}})$.
Proof.
Clearly $d$ is idempotent, and it can be easily checked that ${\mathbf{A}}$ satisfies the following identities:
(8.1) $d(d(x,y,y),x,x)$ $\approx d(x,y,y)\approx d(d(x,y,y),y,y)$
$d(x,x,d(x,x,y))$ $\approx d(x,x,y)\approx d(y,y,d(x,x,y)).$
Define ternary terms $f_{i},g_{i}$ for $i=0,1,2,3$ by
$f_{0}(x,y,z)$ $=x$ $g_{0}(x,y,z)$ $=d(x,z,z)$
$f_{1}(x,y,z)$ $=d(x,y,y)$ $g_{1}(x,y,z)$ $=d(d(x,y,y),z,z)$
$f_{2}(x,y,z)$ $=z$ $g_{2}(x,y,z)$ $=d(x,x,z)$
$f_{3}(x,y,z)$ $=d(y,y,z)$ $g_{3}(x,y,z)$ $=d(x,x,d(y,y,z)).$
From idempotency of $d$ and the identities (8.1), one can deduce that ${\mathbf{A}}$ satisfies the identities
(8.2) $f_{i}(x,y,x)\approx g_{i}(x,y,x)\quad\mbox{for all $i$}$
and
(8.3) $x$ $\approx f_{0}(x,y,y)$ $f_{2}(x,x,y)$ $\approx y$
$f_{1}(x,x,y)$ $\approx f_{0}(x,x,y)$ $f_{2}(x,y,y)$ $\approx f_{3}(x,y,y)$
$f_{1}(x,y,y)$ $\approx g_{1}(x,y,y)$ $g_{3}(x,x,y)$ $\approx f_{3}(x,x,y)$
$g_{0}(x,x,y)$ $\approx g_{1}(x,x,y)$ $g_{3}(x,y,y)$ $\approx g_{2}(x,y,y)$
$g_{0}(x,y,y)$ $\approx d(x,y,y)$ $d(x,x,y)$ $\approx g_{2}(x,x,y).$
The identities (8.2) and (8.3) therefore hold in $\mathsf{HSP}({\mathbf{A}})$. Let ${\mathbf{F}}={\mathbf{F}}_{\mathsf{HSP}({\mathbf{A}})}(3)$ be the free algebra in $\mathsf{HSP}({\mathbf{A}})$ freely generated by $\{x,y,z\}$. As in the discussion preceding Theorem 8.1, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ be calculated in ${\mathbf{F}}$. Interpreting $d$ and $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ as elements of ${\mathbf{F}}$ in the usual way, we find that the identities (8.2) and (8.3) insure that $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ witness
$(x,d)$ $\in\gamma\circ(\alpha\cap(\beta\circ(\alpha\cap\gamma)\circ\beta))\circ\gamma\subseteq\gamma\vee(\alpha\wedge(\beta\vee(\alpha\wedge\gamma)))=\gamma_{2}$
$(d,z)$ $\in\beta\circ(\alpha\cap(\gamma\circ(\alpha\cap\beta)\circ\gamma))\circ\beta\subseteq\beta\vee(\alpha\wedge(\gamma\vee(\alpha\wedge\beta)))=\beta_{2}.$
Thus $d$ is a weak difference term for $\mathsf{HSP}({\mathbf{A}})$ by Theorem 8.1. ∎
Here is our recipe for constructing interesting finite subdirectly irreducible algebras with abelian monoliths.
Example 8.4.
Fix a finite field $\mathbb{F}$. Let $V_{(0,0)},V_{(1,0)},\ldots,V_{(m,0)}$ be finite-dimensional vector spaces over $\mathbb{F}$ with $m\geq 0$ and $\dim(V_{(0,0)})>0$, chosen so that $V_{(0,0)},\ldots,V_{(m,0)}$ are pairwise disjoint. For each $\ell\leq m$ let $W_{(\ell,0)},W_{(\ell,1)},\ldots,W_{(\ell,n_{\ell})}$ be a list of (not necessarily distinct) subspaces of $V_{(\ell,0)}$ with $W_{(\ell,0)}=V_{(\ell,0)}$.
Choose vector spaces $V_{(\ell,i)}$ over $\mathbb{F}$ and isomorphisms $\sigma_{(\ell,i)}:V_{(\ell,i)}\cong W_{(\ell,i)}$ for $\ell\leq m$ and $1\leq i\leq n_{\ell}$, so that the spaces $V_{(\ell,i)}$ ($0\leq\ell\leq m$, $0\leq i\leq n_{\ell}$) are pairwise disjoint. Also set $\sigma_{(\ell,0)}=\mathrm{id}_{V_{(\ell,0)}}$ for each $\ell\leq m$. Let $C_{\ell}$ be the (disjoint) union of $V_{(\ell,0)},V_{(\ell,1)},\ldots,V_{(\ell,n_{\ell})}$, and let $A$ be the (disjoint) union of $C_{0},\ldots,C_{m}$. Then let $\mu$ and $\alpha$ be the equivalence relations on $A$ whose equivalence classes are the spaces $V_{(\ell,i)}$ ($\ell\leq m$, $i\leq n_{\ell}$) and the sets $C_{0},\ldots,C_{m}$ respectively. Also let $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$ and let $\psi$ be the equivalence relation on $B$ whose equivalence classes are $V_{(0,0)},\ldots,V_{(m,0)}$. We will build an algebra ${\mathbf{A}}$ with universe $A$ such that:
  • ${\mathbf{A}}$ belongs to a variety with a weak difference term.
  • ${\mathbf{A}}$ is subdirectly irreducible with abelian monolith $\mu$. In fact, $\mu$ will be the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
  • $\alpha=(0:\mu)$.
  • Letting $\Delta=\Delta_{\mu,\alpha}$ and $\varphi=\overline{\alpha}/\Delta$, there is a bijection from $D({\mathbf{A}})$ to $B$ which sends $\varphi$ to $\psi$ and sends each $\varphi$-class $(\mu\cap(C_{\ell})^{2})/\Delta$ to $V_{(\ell,0)}$.
  • Modulo this bijection, for each $\alpha$-class $C_{\ell}$, the ranges (see Definition 5.5) of the $\mu$-classes $V_{(\ell,i)}$ contained in $C_{\ell}$ correspond precisely to the subspaces $W_{(\ell,i)}$ in $V_{(\ell,0)}$.
  • The division ring associated to ${\mathbf{A}},\mu$, and any transversal for $\mu$ is isomorphic to $\mathbb{F}$.
Begin by letting $S=\{(\ell,i)\,:\,\text{$\ell\leq m$ and $i\leq n_{\ell}$}\}$. Thus $S$ is the set of indices for the vector spaces $V_{(\ell,i)}$ which constitute $A$. Define a binary operation $\wedge$ on $S$ by
$(\ell,i)\wedge(k,j)=\left\{\begin{array}[]{cl}(\ell,i)&\mbox{if $(\ell,i)=(k,j)$}\\ (\ell,0)&\mbox{if $k=\ell$ but $i\neq j$}\\ (0,0)&\mbox{if $k\neq\ell$.}\end{array}\right.$
Then $(S,\wedge)$ is the meet semilattice with least element $(0,0)$ pictured in Figure 2.
[…]
Figure 2. The semilattice $(S,\wedge)$
Recall that for each $s=(\ell,i)\in S$ we have a corresponding vector space $V_{s}$ over $\mathbb{F}$. To each pair of elements $s,t\in S$ with $s\geq t$ in the semilattice ordering, we associate an $\mathbb{F}$-linear map $f_{(s,t)}:V_{s}\rightarrow V_{t}$ as follows.
  1. (1)
    $f_{(s,s)}=\mathrm{id}_{V_{s}}$ for all $s\in S$.
  2. (2)
    $f_{((\ell,i),(\ell,0))}=\sigma_{(\ell,i)}$ for $(\ell,i)\in S$ with $i\neq 0$.
  3. (3)
    $f_{((\ell,i),(0,0))}=$ the constant zero map, for $(\ell,i)\in S$ with $\ell\neq 0$.
Now let $d:A^{3}\rightarrow A$ be the semilattice-over-Maltsev operation (see Definition 8.2) defined by the data $(S,\wedge)$, $(V_{s}:s\in S)$, $(f_{(s,t)}:s,t\in S,\,s\geq t)$, and the family of Maltsev operations $m_{s}(x,y,z)=x-y+z$ evaluated in each vector space $V_{s}$. That is, for all $a\in V_{s_{1}}$, $b\in V_{s_{2}}$ and $c\in V_{s_{3}}$, letting $t=s_{1}\wedge s_{2}\wedge s_{3}$ we have
[…]
For each $\ell\leq m$ define $\sigma_{\ell}:C_{\ell}\rightarrow V_{(\ell,0)}$ so that $\sigma_{\ell}{\upharpoonright}_{V_{(\ell,i)}}=\sigma_{(\ell,i)}$ for all $i\leq n_{\ell}$. For each $s\in S$ let $0_{s}$ denote the zero element of $V_{s}$. We define some unary operations on $A$.
  1. (1)
    For each $(\ell,i)\in S$ and each nonconstant $\mathbb{F}$-linear map $g:V_{(\ell,0)}\rightarrow V_{(\ell,i)}$, define $F_{(\ell,i)}^{g}:A\rightarrow A$ by
    $F_{(\ell,i)}^{g}(a)=\left\{\begin{array}[]{cl}(g\circ\sigma_{\ell})(a)&\mbox{if $x\in C_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in C_{k}$ with $k\neq\ell$.}\end{array}\right.$
  2. (2)
    For each $\ell\in\{1,\ldots,m\}$ with $\dim(V_{(\ell,0)})>0$, choose nonconstant $\mathbb{F}$-linear maps $g_{\ell}:V_{(0,0)}\rightarrow V_{(\ell,0)}$ and $h_{\ell}:V_{(\ell,0)}\rightarrow V_{(0,0)}$ and define $G_{\ell},G^{\prime}_{\ell}:A\rightarrow A$ by
    $G_{\ell}(a)=\left\{\begin{array}[]{cl}(g_{\ell}\circ\sigma_{0})(a)&\mbox{if $a\in C_{0}$}\\ 0_{(k,0)}&\mbox{if $a\in C_{k}$ with $k\neq 0$}\end{array}\right.$
    and
    $G_{\ell}^{\prime}(a)=\left\{\begin{array}[]{cl}(h_{\ell}\circ\sigma_{\ell})(a)&\mbox{if $a\in C_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in C_{k}$ with $k\neq\ell$}\end{array}\right.$
We also need the following binary operations.
  1. (3)
    Fix $1\in V_{(0,0)}$ with $1\neq 0_{(0,0)}$. For every $s\in S$, define $H_{s}:A^{2}\rightarrow A$ by
    $H_{s}(a,b)=\left\{\begin{array}[]{cl}1&\mbox{if $a,b\in V_{s}$}\\ 0_{(0,0)}&\mbox{otherwise.}\end{array}\right.$
  2. (4)
    For each $\ell\in\{1,\ldots,m\}$, define $K_{\ell}:A^{2}\rightarrow A$ by
    $K_{\ell}(a,b)=\left\{\begin{array}[]{cl}b&\mbox{if $a\in C_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\not\in C_{\ell}$ and $b\in C_{k}$}.\end{array}\right.$
Let $\mathscr{F}$ be the set of all the operations $F_{(\ell,i)}^{g}$, let $\mathscr{G}$ be the set of all the operations $G_{\ell}$ and $G_{\ell}^{\prime}$, let =$\mathscr{H}=\{H_{s}\,:\,\text{$s\in S$}\}$, and let $\mathscr{K}=\{K_{\ell}\,:\,\text{$1\leq\ell\leq m$}\}$. Finally, let
${\mathbf{A}}=(A,\{d\}\cup\mathscr{F}\cup\mathscr{G}\cup\mathscr{H}\cup\mathscr{K}).$
Claim 8.4.1.
  1. (1)
    $d$ is a weak difference term $\mathsf{HSP}({\mathbf{A}})$.
  2. (2)
    ${\mathbf{A}}$ is subdirectly irreducible with abelian monolith $\mu$. Moreover, $\mu$ is the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
  3. (3)
    $(0:\mu)=\alpha$.
  4. (4)
    $\Delta_{\mu,\alpha}$ is the set of all pairs $((a,b),(a^{\prime},b^{\prime}))$ such that for some $\ell\leq m$ and some $i,j\leq n_{\ell}$, $\{a,b\}\subseteq V_{(\ell,i)}$ and $\{a^{\prime},b^{\prime}\}\subseteq V_{(\ell,j)}$ and
    $\sigma_{(\ell,i)}(a)-\sigma_{(\ell,i)}(b)=\sigma_{(\ell,j)}(a^{\prime})-\sigma_{(\ell,j)}(b^{\prime})$ calculated in $V_{(\ell,0)}$.
Proof.
(1) follows from Lemma 8.3. It is easy to check that $\mu$ and $\alpha$ are congruences of ${\mathbf{A}}$. Let $\theta\in\operatorname{Con}({\mathbf{A}})$ satisfy $0\neq\theta\subseteq\mu$. (We want to show $\theta=\mu$.) We can choose $(\ell,i)\in S$ so that $\theta\cap(V_{(\ell,i)})^{2}$ contains a nontrivial pair $(c,c^{\prime})$, i.e., with $c\neq c^{\prime}$. Let $a=c-c^{\prime}$ calculated in $V_{(\ell,i)}$. Then $a\neq 0_{(\ell,i)}$ and $(a,0_{(\ell,i)})\in\theta$ because of $d(x,y,z)$. Now let $j\leq n_{\ell}$ and $b\in V_{(\ell,j)}$ with $b\neq 0_{(\ell,j)}$. Let $a^{\prime}=\sigma_{(\ell,i)}(a)$. Then $a^{\prime}\neq 0_{(\ell,0)}$. Hence there exists an $\mathbb{F}$-linear map $g:V_{(\ell,0)}\rightarrow V_{(n,j)}$ with $g(a^{\prime})=b$. Then $F_{(\ell,j)}^{g}(a)=b$ while $F_{(\ell,j)}^{g}(0_{(\ell,i)})=0_{(\ell,j)}$, proving $(b,0_{(\ell,j)})\in\theta$. As $j$ and $b$ were arbitrary, this proves $\mu\cap(C_{\ell})^{2}\subseteq\theta$.
If $\ell>0$, then we can use the operation $G_{\ell}^{\prime}$ to show that $\theta\cap(V_{(0,0)})^{2}$ contains a nontrivial pair. Then the argument in the previous paragraph gives $\mu\cap(C_{0})^{2}\subseteq\theta$. Now let $k\in\{1,\ldots,m\}$ be such that some $\mu$-class $V_{(k,j)}$ contained in $C_{k}$ has more than one element. Then we can use the operation $G_{k}$ to show that $\theta\cap(V_{(k,0)})^{2}$ contains a nontrivial pair. Again by the argument in the previous paragraph, we get $\mu\cap(C_{k})^{2}\subseteq\theta$. This finishes the proof that $\theta=\mu$; hence $0_{A}\prec\mu$.
Next let $(c,c^{\prime})$ be any nontrivial pair in $A^{2}$ with $(c,c^{\prime})\not\in\mu$. Say $c\in V_{s}$ and $c^{\prime}\in V_{t}$ with $s\neq t$. Then the polynomial $f(x)=H_{s}(x,c)$ sends $(c,c^{\prime})$ to $(1,0_{(0,0)})$, so $\operatorname{Cg}(c,c^{\prime})\cap\mu\neq 0$. As we have shown $0_{A}\prec\mu$, this proves $\mu$ is the monolith of ${\mathbf{A}}$. Combining the above argument with Corollary 3.2, we also get that for all $(a,b)\in\mu$ and all $(c,c^{\prime})\in A^{2}$ with $c\neq c^{\prime}$, there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ wth $f(c)=a$ and $f(c^{\prime})=b$. This proves that $\mu$ is the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
Next we will show that $C(\alpha,\mu;0)$ holds. Define $R$ to be the set of all matrices $\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in A^{2\times 2}$ such that for some $\ell\leq m$ and some $i,j\leq n_{\ell}$,
  • $0\in\{i,j\}$ or $i=j$.
  • $a,b\in V_{(\ell,i)}$ and $a^{\prime},b^{\prime}\in V_{(\ell,j)}$.
  • $\sigma_{\ell}(a)-\sigma_{\ell}(b)=\sigma_{\ell}(a^{\prime})-\sigma_{\ell}(b^{\prime})$ calculated in $V_{(\ell,0)}$.
One can show that $R$ is a subuniverse of ${\mathbf{A}}^{2\times 2}$ (exercise). Clearly $X(\mu,\alpha)\subseteq R$ (see Definition 2.1). Conversely, suppose
$\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in R\quad\mbox{with $a,b\in V_{(\ell,i)}$ and $a^{\prime},b^{\prime}\in V_{(\ell,j)}$}.$
We can assume with no loss of generality that $j=0$ or $i=j$. Then
$\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}=d\left(\begin{pmatrix}a&a^{\prime}\\ a&a^{\prime}\end{pmatrix},\begin{pmatrix}a&a\\ a&a\end{pmatrix},\begin{pmatrix}a&a\\ b&b\end{pmatrix}\right),$
proving $R\subseteq\operatorname{Sg}^{{\mathbf{A}}^{2\times 2}}(X(\mu,\alpha))$. These remarks prove $R=M(\mu,\alpha)$. Since every matrix in $R$ has the property that the first column is in $0_{A}$ if and only if the second column is in $0_{A}$, we have proved that $C(\alpha,\mu;0)$ holds and therefore $\mu$ is abelian and $(0:\mu)\geq\alpha$. To prove $(0:\mu)=\alpha$, suppose $\theta\in\operatorname{Con}({\mathbf{A}})$ and $\theta>\alpha$. Pick $(a,b)\in\theta\setminus\alpha$, say with $a\in C_{\ell}$ and $b\in C_{k}$ with $k\neq\ell$. Using the operation $K_{\ell}$ and the pair $(0_{(0,0)},1)\in\mu\setminus 0_{A}$, we have
$K_{\ell}(b,0_{(0,0)})=0_{(0,0)}=K_{\ell}(b,1)$
but
$K_{\ell}(a,0_{(0,0)})=0_{(0,0)}\neq 1=K_{\ell}(a,1),$
proving that $C(\theta,\mu;0)$ fails for every $\theta>\alpha$. Hence $(0:\mu)=\alpha$.
Recall from Lemma 4.4(1) that $\Delta_{\mathsf{h}}(\mu,\alpha)$ is the horizontal transitive closure of $M(\mu,\alpha)$. Thus $\Delta_{\mu,\alpha}$ is the transitive closure of
$\left\{((a,b),(a^{\prime},b^{\prime})):\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in R\right\}.$
The description of $\Delta_{\mu,\alpha}$ in (4) then follows. ∎
Recall that $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$. Let $\Delta=\Delta_{\mu,\alpha}$ and $D({\mathbf{A}})={\mathbf{A}}(\mu)/\Delta$. Observe that every element of $D({\mathbf{A}})$ can be uniquely written in the form $(a,0_{(\ell,0)})/\Delta$ for some $\ell\leq m$ and $a\in V_{(\ell,0)}$. Let $h:D({\mathbf{A}})\rightarrow B$ be the bijection sending $(a,0_{(\ell,0)})/\Delta\mapsto a$. Then by construction, for every $(\ell,i)\in S$ we have
$\operatorname{ran}(V_{(\ell,i)})$ $=\operatorname{ran}(\chi_{0_{(\ell,i)}})$ Definition 5.5
=$=\{(a,0_{(\ell,i)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$
=$=\{(\sigma_{(\ell,i)}(a),0_{(\ell,0)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$ Claim 8.4.1(4)
=$=\{(a^{\prime},0_{(\ell,0)})/\Delta\,:\,\text{$a^{\prime}\in W_{(\ell,i)}$}\}$ as $\sigma_{(\ell,i)}(V_{(\ell,i)})=W_{(\ell,i)}$.
Hence $h(\operatorname{ran}(V_{(\ell,i)}))=W_{(\ell,i)}$ as desired.
Finally, let $T$ be the transversal for $\mu$ given by $T=\{0_{s}\,:\,\text{$s\in S$}\}$, and let $\mathbb{F}^{\prime}$ be the division ring associated to ${\mathbf{A}},\mu,T$. We will verify that $\mathbb{F}^{\prime}\cong\mathbb{F}$. For each $\zeta\in\mathbb{F}$ and $s\in S$ let $\zeta_{s}:V_{s}\rightarrow V_{s}$ be the “scalar multiplication by $\zeta$ in $V_{s}$” map and let $\Theta(\zeta)=(\zeta_{s})_{s\in S}$. Every $n$-ary basic operation of ${\mathbf{A}}$, when restricted to an $n$-tuple of $\mu$-classes, is $\mathbb{F}$-affine, and this property is inherited by polynomial operations. It follows that $\Theta(\zeta)\in\mathbb{F}^{\prime}$. To prove $\operatorname{ran}(\Theta)=\mathbb{F}^{\prime}$, let $\lambda=(\lambda_{s})_{s\in S}\in\mathbb{F}^{\prime}$. Then
($\ast$) $\lambda_{s}\circ f{\upharpoonright}_{V_{t}}=f\circ\lambda_{t}$ for all $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $s,t\in S$ with $f(0_{t})=0_{s}$.
Fix $\ell\leq m$ with $\dim(V_{(\ell,0)})>0$. Since every $\mathbb{F}$-linear map $V_{(\ell,0)}\rightarrow V_{(\ell,0)}$ is a restriction to $V_{(\ell,0)}$ of some unary polynomial, ($\ast$8.4) implies that $\lambda_{(\ell,0)}$ commutes with every $r\in\operatorname{End}_{\mathbb{F}}(V_{(\ell,0)})$. Hence there exists $\zeta=\zeta^{(\ell)}\in\mathbb{F}$ such that $\lambda_{(\ell,0)}=\zeta_{(\ell,0)}$. If $\ell\neq 0$, then recall the nonconstant $\mathbb{F}$-linear map $g_{\ell}:V_{(0,0)}\rightarrow V_{(\ell,0)}$ used to define the basic operation $G_{\ell}$. Then $g_{\ell}=G_{\ell}{\upharpoonright}_{V_{(0,0)}}$, so ($\ast$8.4) gives $\lambda_{(\ell,0)}\circ g_{\ell}=g_{\ell}\circ\lambda_{(0,0)}$. Let $\zeta=\zeta^{(0,0)}$ and $\zeta^{\prime}=\zeta^{(\ell,0)}$. Choose $a\in V_{(0,0)}$ with $b:=g_{\ell}(a)\neq 0_{(\ell,0)}$. Then
$\zeta\cdot b=\zeta\cdot g_{\ell}(a)$ $=g_{\ell}(\zeta\cdot a)$ $\mathbb{F}$-linearity of $g_{\ell}$
$=g_{\ell}(\zeta_{(0,0)}(a))$ definition of $\zeta_{(0,0)}$
$=g_{\ell}(\lambda_{(0,0)}(a))$
$=\lambda_{(\ell,0)}(g_{\ell}(a))$ by ($\ast$8.4)
$=\zeta^{\prime}\cdot b.$
Since $b\neq 0_{(\ell,0)}$, we get $\zeta=\zeta^{\prime}$, i.e., $\zeta^{(0,0)}=\zeta^{(\ell,0)}$. A similar argument using operations in $\mathscr{F}$ gives $\zeta^{(\ell,i)}=\zeta^{(\ell,0)}$ whenever $\ell\leq m$ and $\dim(V_{(\ell,i)})>0$. Hence there exists $\zeta\in\mathbb{F}$ such that $\zeta^{s}=\zeta$ for all $s\in S$ with $\dim(V_{s})>0$. It follows that $\lambda=\Theta(\zeta)$, so $\Theta(\mathbb{F})=\mathbb{F}^{\prime}$. The reader can verify that $\Theta:\mathbb{F}\cong\mathbb{F}^{\prime}$.
9. Conclusion and questions
In this paper we have extended Freese’s analysis of abelian minimal congruences, and his similarity relation between subdirectly irreducible algebras, from the congruence modular setting to the setting of varieties with a weak difference term. We also established a connection between similarity and certain constructs which we have called “proper bridges.” In the companion paper [33], we will compare proper bridges to Zhuk’s original bridges (which are superficially weaker) and explain some related elements of Zhuk’s work in terms of centrality and similarity. In a third paper [32], we will unite the analyses of “rectangular critical relations” by Zhuk [37, Lemma 8.21] (in the idempotent locally finite Taylor case) and by Kearnes and Szendrei [20] (in the congruence modular case), showing in particular the role of similarity, associated finite fields, and linear equations.
In the congruence modular setting, Freese and McKenzie [6] established some nontrivial results concerning the similarity relation, such as bounding the number of similarity classes of subdirectly irreducible algebras in finitely generated varieties. One can ask whether any of these properties persist in varieties with a weak difference term, or with a difference term. For example:
Problem 1.
If ${\mathbf{A}}$ is a finite algebra and $\mathsf{HSP}({\mathbf{A}})$ has a weak difference term, does it follow that there are only finitely many similarity classes of subdirectly irreducible algebras in $\mathsf{HSP}({\mathbf{A}})$ with abelian monolith?
Because of their connection to the Constraint Satisfaction Dichotomy Theorem, finite subdirectly irreducible algebras with abelian monoliths in an idempotent locally finite Taylor variety are of particular interest. We have seen one result (Lemma 5.9) showing how the assumption of idempotency can simplify our analysis. In what other ways does idempotency affect the analysis? For example:
Problem 2.
Suppose ${\mathbf{A}}$ is a finite idempotent subdirectly irreducible algebra in a Taylor variety with abelian monolith $\mu$, $\alpha=(0:\mu)$, and $C$ is an $\alpha$-class. If $q$ is the maximal size of a $\mu$-class in $C$, does it follow that every $\mu$-class in $C$ has size $q$ or 1?