### v2

# Abelian congruences and similarity in varieties with a weak difference term

Abstract.
This is the first of three papers motivated by the author’s desire to understand and explain “algebraically” one aspect of Dmitriy Zhuk’s proof of the CSP Dichotomy Theorem. In this paper we study abelian congruences in varieties having a weak difference term. Each class of the congruence supports an abelian group structure; if the congruence is minimal, each class supports the structure of a vector space over a division ring determined by the congruence. A construction due to J. Hagemann, C. Herrmann and R. Freese in the congruence modular setting extends to varieties with a weak difference term, and provides a “universal domain” for the abelian groups or vector spaces that arise from the classes of the congruence within a single class of the annihilator of the congruence. The construction also supports an extension of Freese’s similarity relation (between subdirectly irreducible algebras) from the congruence modular setting to varieties with a weak difference term.
Key words and phrases:
universal algebra, abelian congruence, weak difference term, similarity
2010 Mathematics Subject Classification:
08A05, 08B99 (Primary).
The support of the Natural Sciences and Engineering Research Council (NSERC) of Canada is gratefully acknowledged.
1. Introduction
Arguably the most important result in universal algebra in the last ten years is the positive resolution of the Constraint Satisfaction Problem (CSP) Dichotomy Conjecture, announced independently in 2017 by Andrei Bulatov [2] and Dmitriy Zhuk [34, 35]. One particular feature of Zhuk’s proof is his analysis of multi-sorted “rectangular critical” relations $R\leq_{sd}{\mathbf{A}}_{1}\times\cdots\times{\mathbf{A}}_{n}$ of finite algebras ${\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{n}$ in certain locally finite idempotent Taylor varieties. Zhuk showed that, in the varieties he considers, such relations produce derived relations, which he called “bridges,” between certain meet-irreducible congruences $\rho_{i}\in\operatorname{Con}({\mathbf{A}}_{i})$ determined by $R$. Zhuk established a number of useful properties of his bridges and used them to tease out implicit linear equations in the CSP instances he considered.
Strikingly, Zhuk’s definition of “bridge” formally resembles some aspects of Hagemann and Herrmann’s formulation [10] from 1979 of the centrality relation on congruences of a single algebra. Moreover, the “there-exists-a-bridge” relation, under suitable restrictions, has some resemblance to the “similarity” relation of Freese [5] from 1983; this latter relation is between subdirectly irreducible algebras in a congruence modular variety, and can be viewed as an amplification of a certain correspondence involving the centrality relation. As well, Zhuk’s bridges arising in rectangular critical subdirect products in locally finite idempotent Taylor varieties resemble the “graphs of similarity” which Kearnes and Szendrei proved arise in rectangular critical relations of the form $R\leq{\mathbf{A}}^{n}$ in congruence modular varieties [19, Theorem 2.5(6)].
Our main goal in this paper and two companion papers [31, 30] is to establish precise connections between Zhuk’s bridges, centrality, and similarity, and then to apply these connections to the study of rectangular critical relations among finite algebras in arbitrary locally finite Taylor varieties. In order to connect Zhuk’s bridges to similarity, we need to first generalize the similarity relation from its original setting (congruence modular varieties) to a broader class also containing all locally finite idempotent Taylor varieties. For this purpose, we focus on the class of varieties which have a weak difference term. Roughly speaking, a “weak difference term” is an idempotent ternary term that specializes to a Maltsev operation whenever it is restricted to a block of an abelian congruence. Such terms are weak versions of terms, known as “difference terms,” that exist and play an important role in congruence modular varieties [9]. Weak difference terms arose organically in tame congruence theory in the 1980s, when Hobby and McKenzie [12] showed that every locally finite Taylor has such a term. The property of a (general) variety having a weak difference term was first explicitly identified in the 1990s in work of Lipparini [21], Kearnes [14], and Kearnes and Szendrei [18], who explored the extent to which centralizer properties in congruence modular varieties extend to varieties with a weak difference term. Recent work concerning varieties with a weak difference term is represented by [16, 33, 20, 15].
This paper is organized around a construction, due to Freese [5] (building on Hagemann and Herrmann [10]) in the congruence modular setting, which we call the difference algebra of an abelian congruence. Whenever $\theta$ is an abelian congruence of an algebra in a variety with a weak difference term, every $\theta$-class naturally supports the structure of an abelian group. Letting $(0:\theta)$ denote the largest congruence centralizing $\theta$, it turns out that the family $\mathscr{F}$ of abelian groups supported on the $\theta$-classes within a fixed $(0:\theta)$-class $E$ can be coherently superimposed on each other, forming a single abelian group $G_{E}$ for which $\mathscr{F}$ is a directed family of subgroups whose union is $G_{E}$. Roughly speaking, the difference algebra of $\theta$ replaces all the $\theta$-classes within each $(0:\theta)$-class $E$ with a single class (of a new abelian congruence $\varphi$) supporting the group $G_{E}$. We define the difference algebra and its congruence $\varphi$, and prove their basic properties, in Sections 4. We make precise the connection between the abelian groups on $\theta$-classes and $\varphi$-classes in Section 5.
In Section 7, we show that Freese’s similarity relation generalizes nicely to varieties with a weak difference term. In the abelian monolith case, Freese’s original definition of similarity, as well as the reformulation in [6], does not generalize in a straightforward way. Instead, we take as our starting point two equivalent (in the congruence modular setting) characterizations of similarity due to Freese – one based on the difference algebra construction, the other a congruence lattice-theoretic formulation – and show that both characterizations generalize to, and are provably equivalent in, varieties with a weak difference term (Theorems 4.6 and 7.7). We also give a third equivalent characterization using a variant of Zhuk’s bridges, which we call “similarity bridges” (Definition 7.3).
Along the way, we expend some energy explaining a folklore result: that each class of an abelian minimal congruence $\theta$ (in an algebra $\mathbf{A}$} with a weak difference term) naturally supports the structure of a left vector space over a division ring $\mathbb{F}_{\theta}$. Here $\mathbb{F}_{\theta}$ depends on $\mathbf{A}$} and $\theta$ but not on the particular $\theta$-class. We provide a new, simpler construction of $\mathbb{F}_{\theta}$ based on the difference algebra construction in Section 6, and show in Appendix 1 that it agrees with the folklore construction. We also show that if two subdirectly irreducible algebras in a variety with a weak difference term have abelian monoliths and are similar, then the division rings determined by their monoliths are isomorphic.
In the companion paper [31], we will develop Zhuk’s theory of bridges and describe the precise relationships between the objects in Zhuk’s theory, the centrality relation, similarity, and similarity bridges as defined here. In a third paper [30], we will study rectangular critical relations in locally finite Taylor varieties, extending the work of Zhuk [35] and connecting these relations to similarity (as was done by Kearnes and Szendrei [19] in the congruence modular setting). In particular, we will show explicitly how each such relations can be understood as encoding a family of “parallel” linear equations over a finite field.
Acknowledgement. I thank Keith Kearnes for several helpful conversations, and Dmitriy Zhuk for providing the inspiration.
2. Definitions
We assume that the reader is familiar with the fundamentals of universal algebra as given in [4], [24] or [1]. Our notation generally follows that in [24], [1] and [16]. We follow [16] and refer to a set of operation symbols with assigned arities as a signature.
If $\mathbf{A}$} is an algebra and $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\leq\beta$, then $\beta/\alpha$ denotes the congruence of ${\mathbf{A}}/\alpha$ corresponding to $\beta$ via the Correspondence Theorem ([24, Theorem 4.12] or [1, Theorem 3.6]). If in addition $\gamma,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\gamma\leq\delta$, then we write $(\alpha,\beta)\nearrow(\gamma,\delta)$ and say that $(\alpha,\beta)$ is perspective up to $(\gamma,\delta)$, or transposes up to $(\gamma,\delta)$, if $\beta\wedge\gamma=\alpha$ and $\beta\vee\gamma=\delta$. The notation $(\gamma,\delta)\searrow(\alpha,\beta)$ means the same thing. We write $\alpha\prec\beta$ and say that $\beta$ covers $\alpha$, or is an upper cover of $\alpha$, if $\alpha<\beta$ and there does not exist a congruence $\gamma$ satisfying $\alpha<\gamma<\beta$. A congruence $\alpha$ is minimal if it covers 0, and is completely meet-irreducible if $\alpha\neq 1$ and there exists $\alpha^{+}$ with $\alpha\prec\alpha^{+}$ and such that $\alpha<\beta\implies\alpha^{+}\leq\beta$ for all $\beta\in\operatorname{Con}({\mathbf{A}})$. A subset $T\subseteq A$ is a transversal for a congruence $\alpha$ if it contains exactly one element from each $\alpha$-class.
$[n]$ denotes $\{1,2,\ldots,n\}$. If $\mathbf{A}$} is an algebra, an $n$-ary term $t(x_{1},\ldots,x_{n})$ of $\mathbf{A}$} is idempotent if $\mathbf{A}$} satisfies the identity $t(x,\ldots,x)\approx x$, and is a Taylor term if it is idempotent, $n>1$, and for each $i\in[n]$, $\mathbf{A}$} satisfies an identity of the form
$t(u_{1},\ldots,u_{n})\approx t(v_{1},\ldots,v_{n})$
where each $u_{j}$ and $v_{k}$ is the variable $x$ or $y$, and $\{u_{i},v_{i}\}=\{x,y\}$. A particularly important example of a Taylor term is a weak near-unanimity term (WNU), which is an $n$-ary idempotent term $w(x_{1},\ldots,x_{n})$ with $n>1$ which satisfies
$w(y,x,x,\ldots,x)\approx w(x,y,x,\ldots,x)\approx w(x,x,y,\ldots,x)\approx\cdots\approx w(x,\ldots,x,y).$
Another important example of a Taylor term is a Maltsev term; this is a ternary term $p(x,y,z)$ satisfying the identities
(2.1) $p(x,x,y)\approx y\approx p(y,x,x).$
The identities (2.1) are called the Maltsev identities. Any ternary operation (whether a term or not) satisfying them is called a Maltsev operation.
Before defining “weak difference term,” we recall the ternary centralizer relation on congruences and the notion of abelian congruences. Given a non-empty set $A$, let $A^{2\times 2}$ denote the set of all $2\times 2$ matrics over $A$. If $\mathbf{A}$} is an algebra, let ${\mathbf{A}}^{2\times 2}$ denote the algebra with universe $A^{2\times 2}$ which is isomorphic to ${\mathbf{A}}^{4}$ via the bijection
$\begin{pmatrix}a_{1}&a_{3}\\ a_{2}&a_{4}\end{pmatrix}\mapsto(a_{1},a_{2},a_{3},a_{4}).$
Definition 2.1.
Suppose $\mathbf{A}$} is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    ${\mathbf{A}}(\theta,\varphi)$ is the subalgebra of ${\mathbf{A}}^{2\times 2}$ consisting of the matrices whose columns belong to $\theta$ and rows belong to $\varphi$.
  2. (2)
    $M(\theta,\varphi)$ is the subuniverse of ${\mathbf{A}}(\theta,\varphi)$ generated by the set
    $X(\theta,\varphi):=\left\{\begin{pmatrix}c&c\\ d&d\end{pmatrix}:(c,d)\in\theta\right\}\cup\left\{\begin{pmatrix}a&b\\ a&b\end{pmatrix}:(a,b)\in\varphi\right\}.$
    The matrices in $M(\theta,\varphi)$ are called $(\theta,\varphi)$-matrices.
Definition 2.2.
Suppose $\theta,\varphi,\delta\in\operatorname{Con}({\mathbf{A}})$. We say that $\varphi$ centralizes $\theta$ modulo $\delta$, and write $C(\varphi,\theta;\delta)$, if any of the following equivalent conditions holds:
  1. (1)
    For every matrix in $M(\varphi,\theta)$, if one row is in $\delta$, then so is the other row.
  2. (2)
    For every matrix in $M(\theta,\varphi)$, if one column is in $\delta$, then so is the other column.
  3. (3)
    For every term $t(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})$ and all pairs $(a_{i},b_{i})\in\varphi$ ($i\in[m]$) and $(c_{j},d_{j})\in\theta$ ($j\in[n]$),
    if $t(\mathbf{a},\mathbf{c})\stackrel{{\scriptstyle\delta}}{{\equiv}}t(\mathbf{a},\mathbf{d})$, then $t(\mathbf{b},\mathbf{c})\stackrel{{\scriptstyle\delta}}{{\equiv}}t(\mathbf{b},\mathbf{d})$.
More generally, if $S,T$ are tolerances of $\mathbf{A}$} (i.e., reflexive symmetric subuniverses of ${\mathbf{A}}^{2}$), then we can define ${\mathbf{A}}(S,T)$, $X(S,T)$, $M(S,T)$, and the relation $C(S,T;\delta)$ in the same way; see [16, §2.5].
Let $\theta,\delta$ be congruences of an algebra $\mathbf{A}$}. The annihilator of $\theta$ modulo $\delta$, denoted $(\delta:\theta)$, is the unique largest congruence $\varphi$ for which $C(\varphi,\theta;\delta)$ holds. We say that $\theta$ is abelian if $C(\theta,\theta;0)$ holds; equivalently, if $\theta\leq(0:\theta)$. More generally, if $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\theta$, then we say that $\theta$ is abelian modulo $\delta$ if any of the following equivalent conditions hold: (i) $\theta/\delta$ is an abelian congruence of ${\mathbf{A}}/\delta$; (ii) $C(\theta,\theta;\delta)$ holds; (iii) $\theta\leq(\delta:\theta)$. The following lemma expresses that annihilators behave nicely under certain quotients and surjective homomorphisms.
Lemma 2.3.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are algebras in a common signature.
  1. (1)
    If $\delta^{\prime},\delta,\theta\in\operatorname{Con}({\mathbf{A}})$ with $\delta^{\prime}\leq\delta\leq\theta$, then $(\delta/\delta^{\prime}:\theta/\delta^{\prime})=(\delta:\theta)/\delta^{\prime}$.
  2. (2)
    Suppose $f:{\mathbf{A}}\rightarrow{\mathbf{B}}$ is a surjective homomorphism and $\lambda,\mu\in\operatorname{Con}({\mathbf{B}})$ with $\lambda\leq\mu$. Then $f^{-1}((\lambda:\mu))=(f^{-1}(\lambda):f^{-1}(\mu))$.
Proof.
[…]
$\overline{f}((\delta:\theta)/\delta^{\prime})$ $=\overline{f}((\delta/\delta^{\prime}:\theta/\delta^{\prime}))$ by (1)
$=(\overline{f}(\delta/\delta^{\prime}):\overline{f}(\theta/\delta^{\prime}))$ as $\overline{f}$ is an isomorphism
$=(\lambda:\mu),$
which proves $(\delta:\theta)=f^{-1}((\lambda:\mu))$. ∎
Definition 2.4.
Let $\mathcal{V}$ be a variety, ${\mathbf{A}}\in\mathcal{V}$, and $d(x,y,z)$ a ternary term in the signature of $\mathcal{V}$.
  1. (1)
    $d$ is a weak difference term for $\mathbf{A}$} if $d$ is idempotent and for every pair $\delta,\theta$ of congruences with $\delta\leq\theta$ and $\theta/\delta$ abelian, we have
    (2.2) $d(a,a,b)\stackrel{{\scriptstyle\delta}}{{\equiv}}b\stackrel{{\scriptstyle\delta}}{{\equiv}}d(b,a,a)$ for all $(a,b)\in\theta$.
  2. (2)
    $d$ is a weak difference term for $\mathcal{V}$ if it is a weak difference term for every algebra in $\mathcal{V}$.
Note in particular that if $d$ is a weak difference term for $\mathbf{A}$} and $\theta$ is an abelian congruence, then setting $\delta=0$ in (2.2) gives that the restriction of $d$ to any $\theta$-class is a Maltsev operation on that class.
Because the definition of “weak difference term” is not given in terms of identities, it can happen that an algebra has a weak difference term, but does not belong to any variety having a weak difference term. Therefore, we will take care to distinguish “having a weak difference term” from the stronger “belonging to a variety with a weak difference term.”
At the level of varieties, the property of having a weak difference term can be characterized by the existence of idempotent terms satisfying a certain pattern of identities, i.e., by an idempotent Maltsev condition, as shown by Kearnes and Szendrei [18, Theorem 4.8]. (See Theorem A2.1 for a slight improvement.) In the context of locally finite varieties, the situation is simpler.
Theorem 2.5.
For a locally finite variety $\mathcal{V}$, the following are equivalent:
  1. (1)
    $\mathcal{V}$ has a Taylor term.
  2. (2)
    $\mathcal{V}$ has a WNU term.
  3. (3)
    $\mathcal{V}$ has a weak difference term.
Proof.
This follows by combining [12, Theorem 9.6], [28, Corollary 5.3], [22, Theorem 2.2], and [18, Theorem 4.8]. ∎
Lemmas 5.2 and 5.7 and Corollary 5.8 mention difference terms; these are weak difference terms which moreover satisfy the identity $d(x,x,y)\approx y$. By an old theorem of Gumm [9], every congruence modular variety has a difference term.
The proof of Proposition 3.4 will make use of the two-term condition. This is the requirement on a congruence $\theta\in\operatorname{Con}({\mathbf{A}})$ that for all $(\theta,\theta)$-matrices
$\begin{pmatrix}a_{1}&a_{3}\\ a_{2}&a_{4}\end{pmatrix},\begin{pmatrix}b_{1}&b_{3}\\ b_{2}&b_{4}\end{pmatrix}\in M(\theta,\theta),$
if $a_{i}=b_{i}$ for $i=1,2,3$, then $a_{4}=b_{4}$. It is easy an easy exercise to show that if $\theta$ satisfies the two-term condition, then $\theta$ is abelian. The converse implication is also true if $\mathbf{A}$} belongs to a variety having a weak difference term, by [21, Theorem 3.1] or [18, Corollary 4.5]. (See Lemma 3.3 for a slight improvement.)
3. Tools
In this section we gather some technical results which will be used in later sections. We start with results about algebras having a weak difference term.
The first result is an easily proved generalization of the well-known fact (see e.g. [1, Theorem 4.65(1)]) that if an algebra $\mathbf{A}$} has a Maltsev term, then every reflexive subuniverse of ${\mathbf{A}}^{2}$ is a congruence of $\mathbf{A}$}. (A proof can be easily extracted from the proof of [16, Corollary 6.22].)
Lemma 3.1.
Suppose $\mathbf{A}$} is an algebra and $\rho$ is a reflexive subuniverse of ${\mathbf{A}}^{2}$. Suppose $\mathbf{A}$} has a ternary term $d(x,y,z)$ such that for all $(a,b)\in\rho$ we have $d(a,a,b)=b$ and $d(a,b,b)=a$. Then $\rho\in\operatorname{Con}({\mathbf{A}})$.
Corollary 3.2.
Suppose $\mathbf{A}$} is an algebra with a weak difference term, $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian, and $\theta$ is minimal. Then for all $(a,b),(a^{\prime},b^{\prime})\in\theta$, if $a\neq b$ then there exists a unary polynomial $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ satisfying $f(a)=a^{\prime}$ and $f(b)=b^{\prime}$.
Proof.
Let $\rho=\{(f(a),f(b))\,:\,\text{$f\in\operatorname{Pol}_{1}({\mathbf{A}})$}\}$. Then $\rho$ is a reflexive subuniverse of ${\mathbf{A}}^{2}$ satisfying $0_{A}\neq\rho\subseteq\theta$. Hence $\rho\in\operatorname{Con}({\mathbf{A}})$ by Lemma 3.1, and thus $\rho=\theta$ as $\theta$ is minimal. ∎
Lemma 3.3.
Suppose the algebra $\mathbf{A}$} has a weak difference term and $\theta\in\operatorname{Con}({\mathbf{A}})$. If $\theta$ is abelian, then $\theta$ satisfies the two-term condition.
Proof.
See Lipparini’s proof of [21, Theorem 3.1(a)]. His Theorem 3.1(a) assumes that $\mathbf{A}$} belongs to a variety with a difference term, but in fact all that the argument requires is that $\mathbf{A}$} itself have a weak difference term. ∎
As a consequence, a well-known property of difference terms in congruence modular varieties due to Gumm [9] partially extends to algebras with weak difference terms.
Proposition 3.4.
Suppose $\mathbf{A}$} is an algebra with a weak difference term $d(x,y,z)$, and $\theta\in\operatorname{Con}({\mathbf{A}})$ is an abelian congruence of $\mathbf{A}$}. Then for every polynomial operation $t(x_{1},\ldots,x_{n})$ and all $a_{i},b_{i},c_{i}\in A$ $(1\leq i\leq n)$ with $a_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}b_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}c_{i}$,
$d(t(\mathbf{a}),t(\mathbf{b}),t(\mathbf{c}))=t(d(a_{1},b_{1},c_{1}),\ldots,d(a_{n},b_{n},c_{n})).$
Proof.
Obviously $t(\mathbf{a}),t(\mathbf{b}),t(\mathbf{c})$ belong to a common $\theta$-class. Define polynomial operations $r(\mathbf{x},\mathbf{y})=d(t(\mathbf{a}),t(\mathbf{x}),t(\mathbf{y}))$ and $s(\mathbf{x},\mathbf{y})=t(d(a_{1},x_{1},y_{1}),\ldots,d(a_{n},x_{n},y_{n}))$. Then
$\begin{pmatrix}r(\mathbf{a},\mathbf{b})&r(\mathbf{a},\mathbf{c})\\ r(\mathbf{b},\mathbf{b})&r(\mathbf{b},\mathbf{c})\end{pmatrix}=\begin{pmatrix}t(\mathbf{b})&t(\mathbf{c})\\ t(\mathbf{a})&d(t(\mathbf{a}),t(\mathbf{b}),t(\mathbf{c}))\end{pmatrix}=:M_{1}$
while
$\begin{pmatrix}s(\mathbf{a},\mathbf{b})&s(\mathbf{a},\mathbf{c})\\ s(\mathbf{b},\mathbf{b})&s(\mathbf{b},\mathbf{c})\end{pmatrix}=\begin{pmatrix}t(\mathbf{b})&t(\mathbf{c})\\ t(\mathbf{a})&t(d(a_{1},b_{1},c_{1}),\ldots,d(a_{n},b_{n},c_{n}))\end{pmatrix}=:M_{2}.$
Obviously $M_{1},M_{2}$ are $(\theta,\theta)$-matrices, and since they agree at all positions except one, they form an instance of the two-term condition for $\theta$. As $\theta$ is abelian and $\mathbf{A}$} has a weak difference term, $\theta$ satisfies the two-term condition by Lemma 3.3. Thus the lower-right entries of $M_{1}$ and $M_{2}$ are equal, as desired. ∎
The next result is an immediate consequence of the definition of weak difference term and a classical theorem of Gumm [8] and Herrmann [11].
Lemma 3.5.
Suppose $\mathbf{A}$} is an algebra having a weak difference term $d(x,y,z)$, and $\theta$ is an abelian congruence of $\mathbf{A}$}. Then for all $e\in A$, the structure $(e/\theta,+,e)$ is an abelian group where $x+y:=d(x,e,y)$ for $x,y\in e/\theta$. Moreover, in this group we have x=d(e,x,e)-x=d(e,x,e) and d(x,y,z)=xy+zd(x,y,z)=x-y+z for all $x,y,z\in e/\theta$.
Proof.
Let $C=e/\theta$. Then $(C,d{\upharpoonright}_{C^{3}})$ is an abelian algebra with Maltsev operation $d{\upharpoonright}_{C^{3}}$. Now apply Gumm [8, Theorem 4.7] or Herrmann’s main theorem from [11] (or see [29, Corollary 2] for a short proof). ∎
While an algebra may have many different weak difference terms, they all agree whenever restricted to a class of an abelian congruence.
Lemma 3.6.
Suppose $\mathbf{A}$} is an algebra and $\theta$ is an abelian congruence. If $d(x,y,z)$ and $t(x,y,z)$ are two weak difference terms for $\mathbf{A}$}, then for every $\theta$-class $C$ we have that $d{\upharpoonright}_{C^{3}}=t{\upharpoonright}_{C^{3}}$.
Proof.
Again this is a well-known consequence of $(C,d{\upharpoonright}_{C^{3}})$ being an abelian Maltsev operation. For $a,b,c\in C$, Proposition 3.4 yields
d(t(a,b,b),t(b,b,b),t(b,b,c))=t(d(a,b,b),d(b,b,b),d(b,b,c))d(t(a,b,b),t(b,b,b),t(b,b,c))=t(d(a,b,b),d(b,b,b),d(b,b,c))
which simplifies to $d(a,b,c)=t(a,b,c)$. ∎
Definition 3.7.
Suppose $\mathbf{A}$} has a weak difference term, $\theta$ is an abelian congruence, and $e\in A$. We let $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ denote the abelian group with universe $e/\theta$, zero element $e$, and group operation $x+y:=d(x,e,y)$, where $d$ is a weak difference term.
The operation ++ in the previous definition does not depend on the choice of the weak difference term, by Lemma 3.6. The notation $M(\theta,e)$ has been used in [6] and elsewhere for $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$.
The next result is known in the congruence modular setting (e.g., [6, Chapter 5]), and is folklore for varieties with a weak difference term. We sketch a proof showing that only the algebra $\mathbf{A}$} itself needs to have a weak difference term.
Proposition 3.8.
Suppose $\mathbf{A}$} is an algebra having a weak difference term, and $\theta$ is an abelian congruence.
  1. (1)
    Suppose $C_{1},C_{2}$ are $\theta$-classes and $e_{i}\in C_{i}$ for $i=1,2$. If $r\in\operatorname{Pol}_{1}({\mathbf{A}})$ satisfies $r(e_{1})=e_{2}$, then $r{\upharpoonright}_{C_{1}}$ is a group homomorphism from $\operatorname{Grp}_{{\mathbf{A}}}(\theta_{1},e_{1})$ to $\operatorname{Grp}_{{\mathbf{A}}}(\theta_{2},e_{2})$.
  2. (2)
    Suppose $f\in\operatorname{Pol}_{n}({\mathbf{A}})$ and $C_{1},\ldots,C_{n},C$ are $\theta$-classes with $f(C_{1},\ldots,C_{n})\subseteq C$. Fix $\mathbf{e}=(e_{1},\ldots,e_{n})\in C_{1}\times\cdots\times C_{n}$ and $e\in C$. There exist $r_{1},\ldots,r_{n}\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $r_{i}(e_{i})=e$ for $i\in[n]$, such that for all $\mathbf{a}=(a_{1},\ldots,a_{n})\in C_{1}\times\cdots\times C_{n}$,
[…]
Proof sketch.
(1) follows from Proposition 3.4. (2) is proved by induction on $n$. When $n=1$, define $r_{1}(x)=d(f(x),f(e_{1}),e)$, whose restriction to $C_{1}$ is just $f{\upharpoonright}_{C_{1}}(x)-f(e_{1})$ computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. When $n>1$, define $g\in\operatorname{Pol}_{n-1}({\mathbf{A}})$ by
$g(x_{1},\ldots,x_{n-1})=f(x_{1},\ldots,x_{n-1},e_{n})$
and inductively get $r_{1},\ldots,r_{n-1}\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $r_{t}(e_{t})=e$ for $t\in[n-1]$, such that $g(a_{1},\ldots,a_{n-1})=\sum_{t=1}^{n-1}r_{t}(a_{t})+f(\mathbf{e})$ for all $\mathbf{a}\in C_{1}\times\cdots\times C_{n-1}$. Define $r_{n}(x)=d(f(e_{1},\ldots,e_{{n-1}},x),f(\mathbf{e}),e)$ and note that $r_{n}(e_{n})=e$. Then for all $\mathbf{a}\in C_{1}\times\cdots\times C_{n}$, computing in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$,
$\sum_{t=1}^{n}r_{t}(a_{t})+f(\mathbf{e})$ $=\left(\sum_{t=1}^{n-1}r_{t}(a_{t})+f(\mathbf{e})\right)+r_{n}(a_{n})$
$=f(a_{1},\ldots,a_{n-1},e_{n})+(f(e_{1},\ldots,e_{n},a_{n})-f(\mathbf{e}))$
$=d(f(a_{1},\ldots,a_{n-1},e_{n}),f(e_{1},\ldots,e_{n}),f(e_{1},\ldots,e_{{n-1}},a_{n})).$
Using Proposition 3.4, the last expression displayed above is equal to
$f(d(a_{1},e_{1},e_{1}),\ldots,d(a_{n-1},e_{{n-1}},e_{{n-1}}),d(e_{n},e_{n},a_{n}))$
which simplifies to $f(\mathbf{a})$. ∎
The following result is folklore; we include it here for completeness.
Theorem 3.9.
Suppose $\mathbf{A}$} is a finite algebra with a weak difference term, $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian, and $0\prec\theta$. There exists a prime $p$ such that for all $e\in A$, $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ is an elementary abelian $p$-group. Hence every $\theta$-class has cardinality $p^{k}$ for some $k\geq 0$ (where $k$ can depend on the $\theta$-class).
Proof sketch.
This was proved by Freese [5, §4] in the congruence modular case, and essentially the same proof works here. We recast his argument using tools from a recent paper of Mayr and Szendrei [23]. (See also the arguments in Section 6, especially Corollary 6.7, where a generalization is proved.)
First note that we can assume that $\mathbf{A}$} has no nullary operations (by replacing each nullary operation with a constant unary operation). Let ${\mathbf{I}}={\mathbf{A}}/\theta$, and let $\nu_{\theta}:{\mathbf{A}}\rightarrow{\mathbf{I}}$ be the natural map. The pair $({\mathbf{A}},\nu_{\theta})$ is called an algebra over $\mathbf{I}$} in [23, Definition 2.1]. By an old construction of Novotný [26], one can obtain a multisorted algebra ${\mathbf{M}}=\mathfrak{M}({\mathbf{A}},\nu_{\theta})$ whose sort domains are the $\theta$-classes of $\mathbf{A}$}; then applying another old construction of Gardner [7], one can convert $\mathbf{M}$} to a one-sorted algebra whose universe is the Cartesian product of the $\theta$-classes of $\mathbf{A}$}. Call this final algebra $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$. For precise definitions, see [23, Propositions 2.2 and 2.5]. Because $\theta$ is minimal and abelian, it follows from [23, Corollary 3.5 and Theorem 3.6] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple and abelian. And because $\mathbf{A}$} has a term $d(x,y,z)$ whose restriction to each $\theta$-class is a Maltsev operation, it follows from the definitions and [23, Remark 2.8] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ has a Maltsev term. In fact, writing the universe of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ as $C_{1}\times\cdots\times C_{k}$ where $C_{1},\ldots,C_{k}$ is an enumeration of the $\theta$-classes, then the Maltsev operation $m(x,y,z)$ of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is given coordinatewise by $d$.
Since $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple, abelian, and has a Maltsev term, it follows that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is polynomially equivalent to a (finite) simple module by a theorem of Gumm [8] and Herrmann [11]. If we fix a $k$-tuple $\mathbf{e}=(e_{1},\ldots,e_{k})\in C_{1}\times\cdots\times C_{k}$, then we can assume that $\mathbf{e}$ is the zero element of this module and addition is given by $x+y=m(x,\mathbf{e},y)$. By Schur’s Lemma and finiteness, the endomorphism ring of this module is a finite field $\mathbb{F}$, say of cardinality $p^{m}$. It follows that the abelian group $(C_{1}\times\cdots\times C_{k},+,\mathbf{e})$ inherits the structure of a vector space over $\mathbb{F}$, which forces $px=\mathbf{e}$ for all $x\in C_{1}\times\cdots\times C_{k}$. Interpreting this equation coordinatewise gives that each group $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ satisfies $px=e_{i}$ as required. ∎
In the context of the proof of Theorem 3.9, it can be further shown that the finite field $\mathbb{F}$ acts on $C_{1}\times\cdots\times C_{k}$ in a coordinate-wise fashion, meaning that each $\theta$-class $C_{i}$ can be viewed as a vector space over $\mathbb{F}$. In congruence modular varieties, Freese proved an analogous result when $\mathbf{A}$} is infinite: there is a division ring $\mathbb{F}$ determined by $\mathbf{A}$} and the abelian minimal congruence $\theta$ such that every $\theta$-class inherits the structure of a left $\mathbb{F}$-vector space. In section 6 we will extend Freese’s result to arbitrary algebras in varieties with a weak difference term, via a new method for constructing the division ring and its action.
The next lemma will be needed in Section 7. Its proof, in the congruence modular case, essentially can be extracted from the proof of [6, Theorem 10.11].
Lemma 3.10.
Suppose $\mathbf{A}$} is an algebra with a weak difference term and $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian. Let $D_{1},D_{2}$ be subuniverses of $\mathbf{A}$} which are transversals for $\theta$. Then there exists an automorphism $\sigma$ of $\mathbf{A}$} satisfying $\sigma(D_{1})=D_{2}$ and $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}x$ for all $x\in A$.
Proof.
For $i=1,2$ let $\pi_{i}:A\rightarrow D_{i}$ be the unique map given by $\pi_{i}(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}x$ for all $x\in A$, and note that $\pi_{1}$ and $\pi_{2}$ are retractions of $\mathbf{A}$} with kernel $\theta$. Let $d(x,y,z)$ be a weak difference term for $\mathcal{V}$ and define $\sigma:A\rightarrow A$ and $\tau:A\rightarrow A$ by
$\sigma(x)$ $=d(x,\pi_{1}(x),\pi_{2}(x))$
$\tau(x)$ $=d(x,\pi_{2}(x),\pi_{1}(x)).$
Observe first that $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}\tau(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}d(x,x,x)=x$ for all $x\in A$. Fix a $\theta$-class $C$; by construction, there exist $b,c\in C$ such that $\pi_{1}(x)=b$ and $\pi_{2}(x)=c$ for all $x\in C$. Pick an abelian group operation ++ on $C$ such that d(x,y,z)=xy+zd(x,y,z)=x-y+z for all $x,y,z\in C$. Then $\sigma(x)=x-b+c$ and $\tau(x)=x-c+b$ for all $x\in C$, and since this is true for every $\theta$-class, we get $\sigma\circ\tau=\tau\circ\sigma=\mathrm{id}_{A}$, proving $\sigma$ is a bijection. Our aim is to prove that $\sigma$ is the desired automorphism.
Let $f$ be an $n$-ary basic operation. Let $a_{1},\ldots,a_{n}\in A$ and put $a=f(a_{1},\ldots,a_{n})$; we need to prove $\sigma(a)=f(\sigma(a_{1}),\ldots,\sigma(a_{n}))$. Let $b_{i}=\pi_{1}(a_{i})$ and $c_{i}=\pi_{2}(a_{i})$, and note that $a_{i},b_{i},c_{i}$ belong to a common $\theta$-class $C_{i}$ for each $i=1,\ldots,n$. Let $b=f(b_{1},\ldots,b_{n})$ and $c=f(c_{1},\ldots,c_{n})$, and note that $b\in D_{1}$ and $c\in D_{2}$ (since $D_{1},D_{2}\leq{\mathbf{A}}$), so $\pi_{1}(a)=b$ and $\pi_{2}(a)=c$. Thus what must be shown is
$d(f(a_{1},\ldots,a_{n}),f(b_{1},\ldots,b_{n}),f(c_{1}\ldots,c_{n}))=f(d(a_{1},b_{1},c_{1}),\ldots,d(a_{n},b_{n},c_{n}))$
given that $a_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}b_{i}\stackrel{{\scriptstyle\theta}}{{\equiv}}c_{i}$ for all $i=1,\ldots,n$. This follows from Proposition 3.4. Hence $\sigma$ is an automorphism of $\mathbf{A}$}.
Finally, suppose $a\in D_{1}$. Then $\pi_{1}(a)=a$, so $\sigma(a)=d(a,a,\pi_{2}(a))=\pi_{2}(a)\in D_{2}$. This proves $\sigma(D_{1})\subseteq D_{2}$. A similar argument shows $\tau(D_{2})\subseteq D_{1}$, so $\sigma(D_{1})=D_{2}$. ∎
Next we give some results concerning varieties with a weak difference term. The first three Propositions are extracted from [16, 17].
Proposition 3.11.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    If $C(\theta,\theta;\delta)$, then $\theta\vee\delta=\delta\circ\theta\circ\delta$ and $(\theta\vee\delta)/\delta$ is abelian.
  2. (2)
    If $\theta$ is abelian, then $C(\theta,\theta;\delta)$.
Proof.
These are special cases of Lemmas 6.7 and 6.8 respectively from [16]. ∎
Proposition 3.12.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\leq\beta$. If $\beta/\alpha$ is abelian, then the interval $I[\alpha,\beta]$ is a modular lattice of permuting equivalence relations.
Proof.
This is a special case of [16, Theorem 6.16]. ∎
Proposition 3.13.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\sigma,\tau,\delta,\varepsilon\in\operatorname{Con}({\mathbf{A}})$ with $\sigma\leq\tau$ and $\delta\leq\varepsilon$ and $(\sigma,\tau)\nearrow(\delta,\varepsilon)$.
  1. (1)
    $\tau/\sigma$ is abelian if and only if $\varepsilon/\delta$ is abelian.
  2. (2)
    If $\tau/\sigma$ is abelian, then the map $J:I[\sigma,\tau]\rightarrow I[\delta,\varepsilon]$ given by $J(x)=x\vee\delta$ is surjective.
Proof.
(1) follows from [16, Theorem 3.27] as corrected in [17]. (2) is a special case of [16, Theorem 6.24]. ∎
Our last result in this section generalizes [13, Lemma 2.6].
Proposition 3.14.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\delta,\varepsilon,\sigma,\tau\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\varepsilon$ and $\sigma\leq\tau$ where $(\sigma,\tau)\nearrow(\delta,\varepsilon)$.
  1. (1)
    If $\varepsilon/\delta$ is abelian, then $\varepsilon=\delta\circ\tau\circ\delta$ and $(\delta:\varepsilon)=(\sigma:\tau)$.
  2. (2)
    If $\delta\prec\varepsilon$, then $(\delta:\varepsilon)=(\sigma:\tau)$.
Proof.
(1) We have $C(\varepsilon,\varepsilon;\delta)$ by hypothesis, and hence $C(\tau,\tau;\delta)$ by monotonicity. Then $\varepsilon=\tau\vee\delta=\delta\circ\tau\circ\delta$ by Proposition 3.11(1).
$\delta\wedge\tau=\sigma$ implies $C(\delta,\tau;\sigma)$ by e.g. [16, Theorem 2.19(8)], proving $(\sigma:\tau)\geq\delta$. Thus to prove $(\delta:\varepsilon)=(\sigma:\tau)$, it is enough to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\sigma)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. We have $C(\alpha,\tau;\sigma)\iff C(\alpha,\tau;\delta)$ by [16, Theorem 2.19(4)], since $\tau\wedge\sigma=\tau\wedge\delta$. Thus we only need to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\delta)$. The forward implication follows immediately by monotonicity. Conversely, suppose $C(\alpha,\tau;\delta)$. Then $C(\alpha,\delta\circ\tau\circ\delta;\delta)$ by [16, Theorem 2.19(3)], i.e., $C(\alpha,\varepsilon;\delta)$.
(2) If $\varepsilon/\delta$ is abelian, then we can apply item (1). Assume $\varepsilon/\delta$ is nonabelian. Then $\tau/\sigma$ is also nonabelian by Proposition 3.13(1). As in the proof of (1), we only need to prove $C(\alpha,\tau;\delta)\implies C(\alpha,\varepsilon;\delta)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. Assume $C(\alpha,\tau;\delta)$. If $\alpha\geq\varepsilon$, then $C(\tau,\tau;\delta)$ by monotonicity, which contradicts nonabelianness of $\tau/\sigma$. Thus $\alpha\ngeq\varepsilon$. Since $\alpha\geq\delta$ and $\delta\prec\varepsilon$, we get $\alpha\wedge\varepsilon=\delta$ and hence $C(\alpha,\varepsilon;\delta)$. ∎
4. The difference algebra of an abelian congruence
In this section we revisit an old construction of Hagemann and Herrmann [10] and Freese [5] in congruence modular varieties. We will show that the construction adapts nicely to varieties with a weak difference term. In the following definition, we follow the notation of [6].
Definition 4.1.
Suppose $\mathbf{A}$} is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$ with $\theta\leq\alpha$.
  1. (1)
    ${\mathbf{A}}(\theta)$ denotes $\theta$ viewed as a subalgebra of ${\mathbf{A}}^{2}$.
  2. (2)
    $\Delta_{\theta,\varphi}$ denotes the congruence of ${\mathbf{A}}(\theta)$ generated by $\{((a,a),(b,b))\,:\,\text{$(a,b)\in\varphi$}\}$.
The next notations depart from those in [6].
Definition 4.2.
Let ${\mathbf{A}},\theta,\varphi$ be as in Definition 4.1.
  1. (1)
    $\operatorname{pr}_{1},\operatorname{pr}_{2}:{\mathbf{A}}(\theta)\rightarrow{\mathbf{A}}$ denote the two projection homomorphisms.
  2. (2)
    $\eta_{i}=\ker(\operatorname{pr}_{i})\in\operatorname{Con}({\mathbf{A}}(\theta))$ for $i=1,2$.
  3. (3)
    If $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\theta$, then $\overline{\alpha}$ denotes
    $\overline{\alpha}=\{((a_{1},a_{2}),(b_{1},b_{2}))\in A(\theta)\times A(\theta)\,:\,\text{$(a_{1},b_{1})\in\alpha$}\}.$
Note that in the context of Definition 4.2 we have $\overline{\alpha}\in\operatorname{Con}({\mathbf{A}}(\theta))$ and $\overline{\alpha}$ is also given by $\overline{\alpha}=\{((a_{1},a_{2}),(b_{1},b_{2}))\in A(\theta)\times A(\theta)\,:\,\text{$(a_{2},b_{2})\in\alpha$}\}$. Also observe that for each $i=1,2$, $\operatorname{Con}({\mathbf{A}})$ is naturally isomorphic to the interval $I[\eta_{i},1]$ in $\operatorname{Con}({\mathbf{A}}(\theta))$ via $\operatorname{pr}_{i}^{-1}$, and $\alpha$ maps to $\overline{\alpha}$ under either isomorphism; see Figure 1.
It will be useful to view $\Delta_{\theta,\varphi}$ as a set of $(\theta,\varphi)$-matrices. Conversely, it will be useful to view $M(\theta,\varphi)$ as a tolerance of ${\mathbf{A}}(\theta)$.
Definition 4.3.
Suppose $\mathbf{A}$} is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    Let $\Delta_{\mathsf{h}}(\theta,\varphi)$ denote the subuniverse of ${\mathbf{A}}(\theta,\varphi)$ “corresponding to” $\Delta_{\theta,\varphi}$:
    $\Delta_{\mathsf{h}}(\theta,\varphi)=\left\{\begin{pmatrix}a&c\\ b&d\end{pmatrix}:((a,b),(c,d))\in\Delta_{\theta,\varphi}\right\}.$
  2. (2)
    Let $M_{\theta,\varphi}$ denote the subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$ “corresponding to” $M(\theta,\varphi)$:
    $M_{\theta,\varphi}=\left\{((a,b),(c,d)):\begin{pmatrix}a&c\\ b&d\end{pmatrix}\in M(\theta,\varphi)\right\}.$
Lemma 4.4.
Suppose $\mathbf{A}$} is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    $\Delta_{\theta,\varphi}$ is the transitive closure of of $M_{\theta,\varphi}$. Thus $\Delta_{\mathsf{h}}(\theta,\varphi)$ is the “horizontal transitive closure of $M(\theta,\varphi)$.”
  2. (2)
    $\Delta_{\mathsf{h}}(\theta,\varphi)$ is “vertically symmetric.” I.e.,
    $\mbox{If~}\begin{pmatrix}a&c\\ b&d\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi),~\mbox{then}~\begin{pmatrix}b&d\\ a&c\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi).\\$
  3. (3)
    If $\mathbf{A}$} has a weak difference term and $\theta$ is abelian, then $\Delta_{\mathsf{h}}(\theta,\varphi)$ is “vertically transitive.” I.e.,
    If $\begin{pmatrix}a&c\\ b&d\end{pmatrix},\begin{pmatrix}b&d\\ r&s\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi),~\mbox{then}~\begin{pmatrix}a&c\\ r&s\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi).$
  4. (4)
    If $\mathbf{A}$} has a weak difference term and $\theta$ is abelian, then $\Delta_{\mathsf{h}}(\theta,\theta)=M(\theta,\theta)$.
Proof.
(1) $M_{\theta,\varphi}$ is the reflexive, symmetric subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$ generated by $\{((a,a),(b,b))\,:\,\text{$(a,b)\in\varphi$}\}$.
(2) This follows from (1) and the fact that $M(\theta,\varphi)$ is vertically symmetric.
(3) We can consider $\Delta_{\mathsf{h}}(\theta,\varphi)$ as a binary relation on the set of its rows. As such, it is a reflexive subuniverse of ${\mathbf{A}}(\varphi)\times{\mathbf{A}}(\varphi)$, and if $\theta$ is abelian, a weak difference term for $\mathbf{A}$}, interpreted coordinatewise in ${\mathbf{A}}(\varphi)$, will satisfy the hypothesis of Lemma 3.1 for this relation.
(4) Let $\rho=M_{\theta,\theta}$; then $\rho$ is a reflexive subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$. If $\theta$ is abelian and $\mathbf{A}$} has a weak difference term $d$, then $d$ interpreted coordinatewise in ${\mathbf{A}}(\theta)$ satisfies the hypothesis of Lemma 3.1 for $\rho$. Thus $\rho$ is transitive by Lemma 3.1, implying $\Delta_{\theta,\theta}=M_{\theta,\theta}$ by (1) and hence $\Delta_{\mathsf{h}}(\theta,\theta)=M(\theta,\theta)$. ∎
Remark 4.5.
Assuming the hypotheses of Lemma 4.4(3), $\Delta_{\mathsf{h}}(\theta,\varphi)$ is an example of what A. Moorhead calls a “2-dimensional congruence” [25]. In fact, $\Delta_{\mathsf{h}}(\theta,\varphi)$ is the smallest 2-dimensional congruence generated by $M(\theta,\varphi)$, which Moorhead denotes by $\Delta(\theta,\varphi)$.
The next theorem has its roots in [10, Proposition 1.1].
Theorem 4.6.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is a nonzero abelian congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. Let $\Delta=\Delta_{\theta,\alpha}$, let $\eta_{1},\eta_{2},\overline{\theta}$ be as in Definition 4.2, and let $\varepsilon=\overline{\theta}\wedge\Delta$. Then:
  1. (1)
    For any $a\in A$, $(a,a)/\Delta=\{(b,b)\,:\,\text{$(a,b)\in\alpha$}\}$.
  2. (2)
    $\Delta<\overline{\alpha}$ and $(\Delta:\overline{\alpha})=\overline{\alpha}$.
  3. (3)
    $\overline{\theta}$ is abelian, and the interval $I[0,\overline{\theta}]$ in $\operatorname{Con}({\mathbf{A}}(\theta))$ is a modular lattice consisting of permuting congruences.
  4. (4)
    $\{0,\eta_{1},\eta_{2},\varepsilon,\overline{\theta}\}$ is a sublattice of $I[0,\overline{\theta}]$ isomorphic to ${\mathbf{M}}_{3}$.
  5. (5)
    Each of $(\varepsilon,\overline{\theta}),(0,\eta_{1}),(0,\eta_{2})$ transposes up to $(\Delta,\overline{\alpha})$.
  6. (6)
    If $\theta$ is minimal, i.e., $0\prec\theta$, then the interval $I[0,\overline{\theta}]$ has height $2$, $\Delta\prec\overline{\alpha}$, and $\Delta$ is completely meet-irreducible.
Proof.
(1) follows from the fact that $\alpha$ centralizes $\theta$. In $\operatorname{Con}({\mathbf{A}}(\theta))$ we have $\eta_{1}\wedge\eta_{2}=0$ and $\eta_{i}<\overline{\theta}$ for $i=1,2$. Clearly $\Delta\leq\overline{\alpha}$. Item (1) and $\theta\neq 0$ imply $\overline{\theta}\nleq\Delta$ and hence $\Delta<\overline{\alpha}$ and $\varepsilon<\overline{\theta}$. If $\varphi$ is any congruence of $\mathbf{A}$} satisfying $\theta\leq\varphi\leq\alpha$, then we can show that $(\Delta\wedge\overline{\varphi})\vee\eta_{1}=(\Delta\wedge\overline{\varphi})\vee\eta_{2}=\overline{\varphi}$. Indeed, if $((a_{1},a_{2}),(b_{1},b_{2}))\in\overline{\varphi}$ then
$(a_{1},a_{2})\stackrel{{\scriptstyle\eta_{i}}}{{\equiv}}(a_{i},a_{i})\stackrel{{\scriptstyle\Delta\wedge\overline{\varphi}}}{{\equiv}}(b_{i},b_{i})\stackrel{{\scriptstyle\eta_{i}}}{{\equiv}}(b_{1},b_{2}).$
In particular, this proves $\Delta\vee\eta_{i}=\overline{\alpha}$ and $\varepsilon\vee\eta_{i}=\overline{\theta}$ for $i=1,2$; hence $\varepsilon\neq 0$. The first equality also proves $\Delta\vee\overline{\theta}=\overline{\alpha}$. We also get $\Delta\wedge\eta_{1}=\Delta\wedge\eta_{2}=0$ by (1); hence $\varepsilon\wedge\eta_{i}=0$ for $i=1,2$. And we can easily verify $\eta_{1}\vee\eta_{2}=\overline{\theta}$. This proves (4) and (5). See Figure 1 showing the placement of these congruences in $\operatorname{Con}({\mathbf{A}}(\theta))$.
$\scriptstyle{\eta_{1}}$$\scriptstyle{\eta_{2}}$$\scriptstyle{0}$$\scriptstyle{1=\overline{1}}$$\scriptstyle{\overline{\theta}}$$\operatorname{Con}({\mathbf{A}}(\theta))=$$\operatorname{Con}({\mathbf{A}})=$$\scriptstyle{0}$$\scriptstyle{\theta}$$\scriptstyle{1}$$\scriptstyle{\alpha}$$\scriptstyle{\overline{\alpha}}$$\scriptstyle{\Delta}$$\scriptstyle{\varepsilon}$
Figure 1. $\operatorname{Con}({\mathbf{A}})$ and $\operatorname{Con}({\mathbf{A}}(\theta))$
Since $\theta$ is abelian, we have $C(\overline{\theta},\overline{\theta};\eta_{i})$ for $i=1,2$. As $\eta_{1}\wedge\eta_{2}=0$, we get $C(\overline{\theta},\overline{\theta};0)$, i.e., $\overline{\theta}$ is abelian. Hence by Proposition 3.12, the interval $I[0,\overline{\theta}]$ is a modular lattice of permuting equivalence relations. This proves (3).
As $\overline{\theta}$ is abelian, we use Proposition 3.11(2) to get $C(\overline{\theta},\overline{\theta};\varepsilon)$, i.e., $\overline{\theta}/\varepsilon$ is abelian. As $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get that $\overline{\alpha}/\Delta$ is abelian by Proposition 3.13(1). Since also $\overline{\theta}/\eta_{1}$ is abelian and $(\eta_{1},\overline{\theta})\searrow(0,\eta_{2})\nearrow(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get $(\eta_{1}:\overline{\theta})=(0:\eta_{2})=(\varepsilon:\overline{\theta})=(\Delta:\overline{\alpha})$ by Proposition 3.14(1). Finally, $(\eta_{1}:\overline{\theta})=\overline{(0:\theta)}=\overline{\alpha}$ by Lemma 2.3(2) and the fact that $(0:\theta)=\alpha$, so $(\Delta:\overline{\alpha})=\overline{\alpha}$. This completes the proof of (2).
Now assume $0\prec\theta$. Then $\eta_{i}\prec\overline{\theta}$ for $i=1,2$. Since $\eta_{1}\wedge\eta_{2}=0$, we get $0\prec\eta_{1},\eta_{2}$ (see e.g. [24, Corollary 2.29(ii)]), and every maximal chain in the interval $I[0,\overline{\theta}]$ has length 2 [24, Theorem 2.37]. Thus $I[0,\overline{\theta}]$ has height 2. Note that this also implies $0\prec\varepsilon\prec\overline{\theta}$. Because $\overline{\theta}/\varepsilon$ is abelian and $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, the map $J:I[\varepsilon,\overline{\theta}]\rightarrow I[\Delta,\overline{\alpha}]$ given by $J(\gamma)=\gamma\vee\varepsilon$ is surjective by Proposition 3.13(2). Thus $\Delta\prec\overline{\alpha}$. Now let $\varphi\in\operatorname{Con}({\mathbf{A}}(\mu))$ with $\Delta\leq\varphi$ but $\overline{\alpha}\nleq\varphi$. Then $\overline{\alpha}\wedge\varphi=\Delta$, so $C(\varphi,\overline{\alpha};\Delta)$, so $\varphi\leq(\Delta:\overline{\alpha})=\overline{\alpha}$, so $\varphi=\overline{\alpha}\wedge\varphi=\Delta$. This proves (6). ∎
The next definition is essentially from [5, 6] in the congruence modular setting; we extend it to varieties with a weak difference term.
Definition 4.7.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\theta\in\operatorname{Con}({\mathbf{A}})$ where $\theta$ is abelian. Let $\alpha=(0:\theta)$.
  1. (1)
    The difference algebra for $\theta$, denoted $D({\mathbf{A}},\theta)$, is the algebra ${\mathbf{A}}(\theta)/\Delta_{\theta,\alpha}$.
  2. (2)
    The derived congruence, which we denote by $\varphi$, is the congruence $\overline{\alpha}/\Delta_{\theta,\alpha}$.
  3. (3)
    The canonical transversal is the set $D^{o}:=\{0_{E}\,:\,\text{$E\in A/\alpha$}\}$.
Corollary 4.8.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is an abelian congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. Let $D({\mathbf{A}},\theta)$ be the difference algebra for $\theta$, let $\varphi$ be the derived congruence, let $D^{o}$ be the canonical transversal, and let $\nu:{\mathbf{A}}(\theta)\rightarrow D({\mathbf{A}},\theta)$ be the natural projection map. Then:
  1. (1)
    $\varphi$ is an abelian congruence of $D({\mathbf{A}},\theta)$.
  2. (2)
    $(0:\varphi)=\varphi$.
  3. (3)
    $D^{o}$ is a transversal for $\varphi$ and $D^{o}\leq D({\mathbf{A}},\theta)$.
  4. (4)
    $\nu^{-1}(D^{o})=0_{A}$.
  5. (5)
    There exists an isomorphism $h:{\mathbf{A}}/\alpha\cong D({\mathbf{A}},\theta)/\varphi$ such that for all $(a,b)\in\theta$, $\nu(a,b)/\varphi=h(a/\alpha)$.
If in addition $\theta$ is a minimal congruence of $\mathbf{A}$}, then $D({\mathbf{A}},\theta)$ is subdirectly irreducible with monolith $\varphi$.
Proof.
Items (1) and (2) follow from Theorem 4.6(2). Let $\Delta=\Delta_{\theta,\alpha}$. For each $\alpha$-class $E$ observe that $0_{E}=\{(c,c)\,:\,\text{$c\in E$}\}$ is a $\Delta$-class by Theorem 4.6(1). Hence $D^{o}\subseteq D({\mathbf{A}},\theta)$. It is easy to check that $D^{o}\leq D({\mathbf{A}},\theta)$. Given $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(a,a)$ and so $(a,b)/\Delta\stackrel{{\scriptstyle\varphi}}{{\equiv}}(a,a)/\Delta\in D^{o}$. Thus $D^{o}$ meets every $\varphi$-class. Suppose $0_{E},0_{E^{\prime}}\in D^{o}$ and $0_{E}\stackrel{{\scriptstyle\varphi}}{{\equiv}}0_{E^{\prime}}$. Pick $c\in E$ and $c^{\prime}\in E^{\prime}$. Then $(c,c)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(c^{\prime},c^{\prime})$, so $(c,c^{\prime})\in\alpha$, so $E=E^{\prime}$. Thus $D^{o}$ is a transversal for $\varphi$.
To prove item (4), we must show that for all $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(c,c)$ for some $c\in A$ if and only if $a=b$; and this follows from Theorem 4.6(1). To prove item (5), define $h$ by $h(a/\alpha)=((a,a)/\Delta)/\varphi$.
Finally, if $\theta$ is a minimal congruence, then Theorem 4.6(6) implies that $D({\mathbf{A}},\theta)$ is subdirectly irreducible with monolith $\varphi$. ∎
5. The classes of an abelian congruence
Suppose $\mathbf{A}$} is an algebra in a congruence modular variety, $\theta$ is an abelian congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. A classical result of Gumm [9, see 3.3] states that if $C,C^{\prime}$ are two $\theta$-classes contained in a common $\alpha$-class $E$, then $|C|=|C^{\prime}|$ and in fact $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e^{\prime})$ are isomorphic for any $e\in C$ and $e^{\prime}\in C^{\prime}$. Moreover, it can be shown that these groups are all isomorphic to $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ where $\varphi$ is the derived congruence of the difference algebra $D({\mathbf{A}},\theta)$. These results extend to varieties with a difference term, as we shall see in Lemma 5.2 below. However they fail in general for varieties with a weak difference term, and in Appendix 2 we will show just how badly they can fail there.
Nonetheless, we shall show that in varieties with a weak difference term, $D({\mathbf{A}},\theta)$ and $\varphi$ have the following property: for every $\alpha$-class $E$, the group $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ forms a sort of “universal domain” for all the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ with $e\in E$.
Definition 5.1.
Suppose $\mathbf{A}$} is an algebra in a variety $\mathcal{V}$ with a weak difference term, $\theta$ is an abelian congruence, $\alpha=(0:\theta)$, $\Delta=\Delta_{\theta,\alpha}$, and $\varphi=\overline{\alpha}/\Delta$. Given $e\in A$, the map $\chi_{e}:e/\theta\rightarrow D({\mathbf{A}},\theta)$ is defined by $\chi_{e}(x)=(x,e)/\Delta$.
Lemma 5.2.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1.
  1. (1)
    For all $e\in A$, the map $\chi_{e}$ is an embedding of $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ into $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{e/\alpha})$.
  2. (2)
    If $\mathcal{V}$ has a difference term, then $\chi_{e}$ is an isomorphism.
Proof.
(1) Let $E=e/\alpha$. Clearly $\chi_{e}(e)=0_{E}$. If $(x,e)/\Delta=(y,e)/\Delta$, then $x,y,e$ all belong to a common $\theta$-class. So
$(x,e)$ $\stackrel{{\scriptstyle\Delta}}{{\equiv}}(y,e),\quad(y,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(y,e),\quad(y,y)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(y,y).$
Applying the weak difference term $d$ to these pairs and using $(x,y),(y,e)\in\theta$ gives
$(x,y)=(d(x,y,y),d(e,e,y))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(d(y,y,y),d(e,e,y))=(y,y).$
By Theorem 4.6(1), we get $x=y$. Thus $\chi_{e}$ is injective. It is easily checked that $\chi_{e}$ preserves ++.
(2) Suppose that $d(x,y,z)$ is a difference term, i.e., $\mathcal{V}\models d(x,x,y)\approx y$. Given $(r,s)/\Delta\in 0_{E}/\varphi$, so $(r,s)\in\theta\cap E^{2}$, define $a:=d(r,s,e)$. Then $a\stackrel{{\scriptstyle\theta}}{{\equiv}}d(r,r,e)=e$, so $a\in e/\theta$. Observe that
$\begin{pmatrix}r&r\\ s&s\end{pmatrix},\ \begin{pmatrix}s&s\\ s&s\end{pmatrix},\begin{pmatrix}s&e\\ s&e\end{pmatrix}\in M(\theta,\alpha).$
Also observe that $d(r,s,s)=r$ as $(r,s)\in\theta$. Thus applying $d(x,y,z)$ to the above matrices gives
$\begin{pmatrix}r&a\\ s&e\end{pmatrix}=\begin{pmatrix}d(r,s,s)&d(r,s,e)\\ d(s,s,s)&d(s,s,e)\end{pmatrix}\in M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha).$
Hence $(r,s)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,e)$, so $\chi_{e}(a)=(r,s)/\Delta$, proving $\chi_{e}$ is surjective. ∎
Lemma 5.3.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1.
  1. (1)
    If $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b\stackrel{{\scriptstyle\theta}}{{\equiv}}e$, then $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(d(a,b,e),e)$.
  2. (2)
    If $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b$ and $a\stackrel{{\scriptstyle\alpha}}{{\equiv}}c\stackrel{{\scriptstyle\alpha}}{{\equiv}}e$, then $(d(a,e,c),d(b,e,c))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$.
  3. (3)
    For all $e\in A$, $\operatorname{ran}(\chi_{e})=C^{2}/\Delta$ where $C=e/\theta$.
Proof.
(1) Let $x=d(a,b,e)$. Observe that
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\ \begin{pmatrix}b&b\\ b&b\end{pmatrix},\begin{pmatrix}e&b\\ e&b\end{pmatrix}\in M(\theta,\theta).$
Applying $d$ to these matrices gives
$\begin{pmatrix}x&a\\ e&b\end{pmatrix}=\begin{pmatrix}d(a,b,e)&d(a,b,b)\\ d(b,b,e)&d(b,b,b)\end{pmatrix}\in M(\theta,\theta)\subseteq M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha).$
Hence $(x,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ as required.
(2) This is proved similarly to (1), by applying $d$ to the matrices
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\quad\begin{pmatrix}e&a\\ e&a\end{pmatrix},\quad\begin{pmatrix}c&a\\ c&a\end{pmatrix}$.
(3) Clearly $\operatorname{ran}(\chi_{e})\subseteq C^{2}/\Delta$. Let $a,b\in C$. Then $d(a,b,e)\in C$ and $\chi_{e}(d(a,b,e))=(d(a,b,e),e)/\Delta=(a,b)/\Delta$ by (1), proving $C^{2}/\Delta\subseteq\operatorname{ran}(\chi_{e})$. ∎
The previous lemma justifies the next definition.
Definition 5.4.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. If $C$ is a $\theta$-class and $E$ is the $\alpha$-class containing $C$, then let $\operatorname{ran}(C)$ denote the subgroup of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ with universe $C^{2}/\Delta$ (equivalently, with universe $\operatorname{ran}(\chi_{e})$ for any $e\in C$). We call $\operatorname{ran}(C)$ the range of $C$.
The next definitions and lemma contain an important tool that will help us understand the arrangement of ranges of $\theta$-classes in their corresponding $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$.
Definition 5.5.
Suppose $\mathbf{A}$} is an algebra with a weak difference term $d$, and $E\subseteq A$.
  1. (1)
    A basic positive $(d,E)$-translation is a unary polymial of $\mathbf{A}$} of the form $d(x,e,e^{\prime})$ or $d(e,e^{\prime},x)$ with $e,e^{\prime}\in E$.
  2. (2)
    A positive $(d,E)$-translation is any composition of zero or more basic positive $(d,E)$-translations. (The composition of zero unary polynomials is $\mathrm{id}_{A}$.)
Definition 5.6.
Suppose $\mathbf{A}$} is an algebra with a weak difference term $d$, $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian, and $\alpha=(0:\theta)$. Let $C,C^{\prime}$ be $\theta$-classes contained in a common $\alpha$-class $E$. We write $C\rightarrow C^{\prime}$ to mean there exists a positive $(d,E)$-translation $f$ such that $f(C)\subseteq C^{\prime}$. We call such $f$ a witness to $C\rightarrow C^{\prime}$.
Thus $\rightarrow$ is a reflexive transitive directed edge relation on $A/\theta$ which does not connect $\theta$-classes from different $\alpha$-classes. Note that $\rightarrow$ depends on the choice of $d$.
Lemma 5.7.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. Fix a weak difference term $d$ for $\mathcal{V}$. Let $E$ be an $\alpha$-class and let $C,C^{\prime}$ be $\theta$-classes contained in $E$.
  1. (1)
    If $C\rightarrow C^{\prime}$ witnessed by $f$, then $\operatorname{ran}(C)\subseteq\operatorname{ran}(C^{\prime})$, and $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ for all $a,b\in C$.
  2. (2)
    There exists a $\theta$-class $C^{\prime\prime}\subseteq E$ with $C\rightarrow C^{\prime\prime}$ and $C^{\prime}\rightarrow C^{\prime\prime}$.
  3. (3)
    If $C\rightarrow C^{\prime}$, then for all $e\in C$ and $e^{\prime}\in C^{\prime}$ there exists a witness $f$ to $C\rightarrow C^{\prime}$ satisfying $f(e)=e^{\prime}$.
  4. (4)
    If $d$ is a difference term for $\mathbf{A}$}, then $C_{1}\rightarrow C_{2}$ for all $\theta$-classes $C_{1},C_{2}$ contained in a common $\alpha$-class.
Proof.
(1) It suffices to prove this for basic positive $(d,E)$-translations. Assume $f(x)=d(x,e,e^{\prime})$ with $e,e^{\prime}\in E$ and $f(C)\subseteq C^{\prime}$. For all $a,b\in C$, $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ follows from Lemma 5.3(2). This also proves $C^{2}/\Delta\subseteq(C^{\prime})^{2}/\Delta$, so $\operatorname{ran}(C)\subseteq\operatorname{ran}(C^{\prime})$. A symmetrical proof works if $f(x)=d(e,e^{\prime},x)$.
(2) Choose $e\in C$ and $e^{\prime}\in C^{\prime}$ and let $C^{\prime\prime}$ be the $\theta$-class containing $d(e,e,e^{\prime})$. $C^{\prime\prime}\subseteq E$ by idempotence of $d$. Define $f_{1}(x)=d(x,e,e^{\prime})$ and $f_{2}(x)=d(e,e,x)$. Then $f_{1},f_{2}$ witness $C\rightarrow C^{\prime\prime}$ and $C^{\prime}\rightarrow C^{\prime\prime}$ respectively.
(3) Suppose $f$ witnesses $C\rightarrow C^{\prime}$, $e\in C$, and $e^{\prime}\in C^{\prime}$. Let $g(x)=d(x,f(e),e^{\prime})$ and $h=g\circ f$. Then $h$ witnesses $C\rightarrow C^{\prime}$ and $h(e)=e^{\prime}$.
(4) Taking $e_{i}\in C_{i}$, the polynomial $f(x)=d(x,e_{1},e_{2})$ will witness $C_{1}\rightarrow C_{2}$, since $\mathcal{V}\models d(x,x,y)$. ∎
Corollary 5.8.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1. Let $E\in A/\alpha$.
  1. (1)
    The set $S_{E}:=\{\operatorname{ran}(C)\,:\,\text{$C\in A/\theta$ and $C\subseteq E$}\}$ is a directed set of subgroups of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ whose union is $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$.
  2. (2)
    Hence if $E$ is finite, there exists $C\in E/\theta$ with $\operatorname{ran}(C)=\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$.
  3. (3)
    If $\mathcal{V}$ has a difference term, then $S_{E}=\{\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})\}$.
Proof.
For all $C_{1},C_{2}\in A/\theta$ with $C_{1},C_{2}\subseteq E$, there exists $C\in A/\theta$ with $C\subseteq E$ and $\operatorname{ran}(C_{1})\cup\operatorname{ran}(C_{2})\subseteq\operatorname{ran}(C)$ by Lemma 5.7; this means $S_{E}$ is directed. To show that the union of $S_{E}$ is all of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$, let $(a,b)/\Delta\in\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$. Let $C=b/\theta$; then $a\in C$ so $(a,b)/\Delta\in\operatorname{ran}(C)\subseteq\bigcup S_{E}$. This proves (1) and (2). Item (3) follows from Lemma 5.2(2). ∎
As mentioned at the start of this section, if ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1 and $\mathcal{V}$ is congruence modular, or more generally has a difference term, then it follows from Lemma 5.2(2) that all $\theta$-classes in a common $\alpha$-class have the same cardinality. As we will show in Appendix 2, this uniformity is no longer to be expected in varieties with a weak difference term. However, we do know of one nontrivial restriction on the sizes of the classes of an abelian minimal congruence $\theta$, when the algebra is idempotent and $(0:\theta)=1$.
Lemma 5.9.
Suppose $\mathbf{A}$} is a finite idempotent algebra in a variety with a Taylor term, $\theta$ is an abelian minimal congruence, and $(0:\theta)=1$. Let $q$ be the maximal size of a $\theta$-class. Then every $\theta$-class has size $1$ or $q$.
Proof.
$\mathsf{HSP}({\mathbf{A}})$ has a weak difference term by finiteness of $\mathbf{A}$} and Theorem 2.5. Pick a $\theta$-class $C_{\mathsf{max}}$ of size $q$. Suppose there exists a $\theta$-class $C$ with $1<|C|<q$. Let $\Delta=\Delta_{\theta,1}$. Fix $e\in C_{\mathsf{max}}$ and $0\in C$. We have $\operatorname{ran}(C)\subset A(\theta)/\Delta=\operatorname{ran}(C_{\mathsf{max}})$ by Lemma 5.2 and Corollary 5.8. Fix $b\in C\setminus\{0\}$ and let $b^{\prime}\in C_{\mathsf{max}}\setminus\{e\}$ be such that $\chi_{0}(b)=\chi_{e}(b^{\prime})$. Also pick $u\in C_{\mathsf{max}}$ with $\chi_{e}(u)\not\in\operatorname{ran}(C)$.
By Corollary 3.2, there exists a unary polynomial $f\in\operatorname{Pol}_{1}({\mathbf{B}})$ with $f(b^{\prime})=u$ and $f(e)=e$. Choose a $(1+n)$-ary term $t(x,\mathbf{y})$ and $\mathbf{c}\in A^{n}$ so that $f(x)=t(x,\mathbf{c})$. Let $u^{\prime}=t(b,0,\ldots,0)$. We have the following:
$(b^{\prime},e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,0),\quad(c_{1},c_{1})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(0,0),\quad\ldots,\quad(c_{n},c_{n})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(0,0)$
where the first congruence holds because $\chi_{e}(b^{\prime})=\chi_{0}(b)$ while the remaining congruences are because $(0:\theta)=1$. Now applying $t$ coordinatewise and using idempotency gives $(u,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(u^{\prime},0)$ which contradicts our choice of $u$. ∎
6. The division ring of an abelian minimal congruence
Let $\mathbf{A}$} be an algebra with a weak difference term, let $\theta$ be an abelian congruence, and assume that $\theta$ is minimal, i.e., $0\prec\theta$. In this section we will explain in detail how to construct a division ring $\mathbb{F}_{\theta}$ and equip each $\theta$-class with the structure of a left vector space over $\mathbb{F}_{\theta}$. One construction of $\mathbb{F}_{\theta}$ was sketched by Freese in [5, Section 4] assuming $\mathbf{A}$} belongs to a congruence modular variety. Our construction, which uses the difference algebra of $\theta$, is simpler and better adapted to varieties which have a weak difference term but no difference term.
6.1. Easy case
We first consider an easy case. Throughout this subsection we make the following blanket assumptions:
Assumption 6.1.
$\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{D}}\in\mathcal{V}$, $\varphi\in\operatorname{Con}{\mathbf{D}}$, $0\prec\varphi$, $\varphi$ is abelian, and $\mathbf{D}$} has a subuniverse $T\leq{\mathbf{D}}$ which is a transversal for $\varphi$.
Definition 6.2.
Given Assumption 6.1,
$F=F_{{\mathbf{D}},\varphi,T}$ $:=\{\lambda\in\operatorname{End}({\mathbf{D}})\,:\,\text{$\lambda(\varphi)\subseteq\varphi$ and $\lambda(e)=e$ for all $e\in T$}\}.$
$0=0_{{\mathbf{D}},\varphi,T}$ $:=\mbox{the retraction ${\mathbf{D}}\rightarrow{\mathbf{T}}$ given by $x\stackrel{{\scriptstyle\varphi}}{{\equiv}}0(x)\in T$.}$
$1$ $:=\mathrm{id}_{D}$
If $\lambda,\mu\in F$, then $\lambda+\mu:D\rightarrow D$ and $\lambda:D\rightarrow D$ are defined by
[…]
$(-\lambda)(x)$ $:=d(0(x),\lambda(x),0(x))$
where $d$ is a weak difference term for $\mathcal{V}$. (Note that this definition does not depend on the choice of $d$, by Lemma 3.6.)
Lemma 6.3.
Given Assumption 6.1:
  1. (1)
    If $\lambda,\mu\in F$, then $\lambda+\mu\in F$, $\lambda\in F$, and $\lambda\circ\mu\in F$.
  2. (2)
    $(F,+,-,\circ,0,1)$ is a division ring.
Proof.
(1) Let $\lambda,\mu\in F$. Clearly $\lambda\circ\mu\in F$. Put $\sigma:=\lambda+\mu$. Fix a $\varphi$-class $C$, let $e$ be the unique element of $C\cap T$, and observe that $\lambda(C)\subseteq C$ and $\mu(C)\subseteq C$ while $0(C)=\{e\}$. Thus $\sigma(C)=d(\lambda(C),e,\sigma(C))\subseteq C$ and $\sigma(e)=d(\lambda(e),e,\sigma(e))=d(e,e,e)=e$, which proves $\sigma(\varphi)\subseteq\varphi$ and $\sigma|_{T}=\mathrm{id}_{T}$.
To prove $\sigma\in\operatorname{End}({\mathbf{D}})$, let $f$ be a $k$-ary basic operation and let $\mathbf{a}=(a_{1},\ldots,a_{k})\in D^{k}$. Let $a^{\prime}=f(\mathbf{a})$, for each $i\in[k]$ let $e_{i}=0(a_{i})\in T$ and let $e^{\prime}=0(a^{\prime})\in T$. Thus $(a^{\prime},e^{\prime})\in\varphi$ and $(a_{i},e_{i})\in\varphi$ for all $i\in[k]$. Observe that $f(\mathbf{e})\stackrel{{\scriptstyle\varphi}}{{\equiv}}f(\mathbf{a})=a^{\prime}\stackrel{{\scriptstyle\varphi}}{{\equiv}}e^{\prime}$ and $f(\mathbf{e})\in T$ (since $T\leq{\mathbf{D}}$), which implies $f(\mathbf{e})=e^{\prime}$. Let $\lambda(\mathbf{a}):=(\lambda(a_{1}),\ldots,\lambda(a_{k}))$ and $\mu(\mathbf{a}):=(\mu(a_{1}),\ldots,\mu(a_{k})$. Since $\varphi$ is abelian, we can apply Proposition 3.4 to $f$ and $\lambda(a_{i})\stackrel{{\scriptstyle\varphi}}{{\equiv}}e_{i}\stackrel{{\scriptstyle\varphi}}{{\equiv}}\mu(a_{i})$ for $i\in[k]$ to get
$d(f(\lambda(\mathbf{a})),f(\mathbf{e}),f(\mu(\mathbf{a})))=f(d(\lambda(a_{1}),e_{1},\mu(a_{1}),\ldots,d(\lambda(a_{k}),e_{k},\mu(a_{k})))).$
Using $f(\lambda(\mathbf{a}))=\lambda(f(\mathbf{a}))$ and $f(\mu(\mathbf{a}))=\mu(f(\mathbf{a}))$, the above equation simplifies to
$\sigma(f(\mathbf{a}))=f(\sigma(a_{1}),\ldots,\sigma(a_{k})),$
proving $\sigma\in\operatorname{End}({\mathbf{D}})$. Thus $\sigma\in F$, and a similar proof shows $\lambda\in F$.
(2) Let $\lambda,\mu\in F$. Observe that for each $\varphi$-class $C$, if $e$ is the unique element of $C\cap T$, then $\lambda{\upharpoonright}_{C}$ and $\mu{\upharpoonright}_{C}$ are endomorphisms of the abelian group $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$, $0{\upharpoonright}_{C}$ is the constant zero endomorphism of this group, and $(\lambda+\mu){\upharpoonright}_{C}$ is the pointwise sum of $\lambda{\upharpoonright}_{C}$ and $\mu{\upharpoonright}_{C}$ in this group. Thus the map $\Phi:F\rightarrow\prod_{e\in T}\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e))$ given by $\Phi(\lambda)_{e}=\lambda{\upharpoonright}_{C}$ where $C=e/\varphi$, is a unital embedding of $(F,+,-,\circ,0,1)$ into the direct product of the endomorphism rings of the groups $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$ for $e\in T$. Hence $(F,+,-,\circ,0,1)$ is a unital ring. It remains to show that each element of $F\setminus\{0\}$ is invertible.
Claim. $T$ is a maximal proper subuniverse of $\mathbf{D}$}.
Clearly $T\neq D$ since $\varphi\neq 0$ and $T$ is a transversal for $\varphi$. To prove the Claim, suppose $T<S\leq{\mathbf{D}}$ and $b\in D$; we will show $b\in S$. Fix $a\in S\setminus T$. Let $e=0(a)$ and $e^{\prime}=0(b)$, so $e,e^{\prime}\in T$ and $(a,e),(b,e^{\prime})\in\varphi$. Observe that $a\neq e$ (since $a\not\in T$), so $\operatorname{Cg}(a,e)=\varphi$, so $(b,e^{\prime})\in\operatorname{Cg}(a,e)$. By Corollary 3.2, there exists $f\in\operatorname{Pol}_{1}({\mathbf{D}})$ with $f(a)=b$ and $f(e)=e^{\prime}$. Pick a term $t(x,\mathbf{y})$ and parameters $c_{1},\ldots,c_{n}\in D$ with $f(x)=t(x,\mathbf{c})$. For each $i\in[n]$ let $e_{i}=0(c_{i})$, so $e_{i}\in T$ and $(c_{i},e_{i})\in\varphi$. Then
$t(e,e_{1},\ldots,e_{n})\stackrel{{\scriptstyle\varphi}}{{\equiv}}t(e,c_{1},\ldots,c_{n})=f(e)=e^{\prime}.$
Since $e^{\prime}$ and $t(e,e_{1},\ldots,e_{n})$ are in $T$ and $T$ is a transversal for $\varphi$, it follows that they are equal, so
(6.1) $t(\underline{e},e_{1},\ldots,e_{n})=t(\underline{e},c_{1},\ldots,c_{n}).$
As $\varphi$ is abelian, we can replace the underlined occurrences of $e$ in equation (6.1) with $a$ to get
(6.2) $t(a,e_{1},\ldots,e_{n})=t(a,c_{1},\ldots,c_{n}).$
Note that the left-hand side of equation 6.2 is in $S$, since $\{a,e_{1},\ldots,e_{n}\}\subseteq S$, while the right-hand side is $f(a)=b$, proving the Claim.
Now we finish the proof of (2). Let $\lambda\in F\setminus\{0\}$. One can easily check that $T\leq\operatorname{ran}(\lambda)\leq{\mathbf{D}}$ and $T\leq\lambda^{-1}(T)\leq{\mathbf{D}}$. Because $\lambda\neq 0$, we get $\operatorname{ran}(\lambda)\neq T$ and $\lambda^{-1}(T)\neq D$. Hence by the Claim, $\operatorname{ran}(\lambda)=D$ and $\lambda^{-1}(T)=T$. Thus $\lambda$ is surjective, and $\lambda{\upharpoonright}_{C}$ has trivial kernel for each $\varphi$-class $C$, so $\lambda$ is injective as well. Letting $\lambda^{-1}$ be the functional inverse of $\lambda$, we can easily check that $\lambda^{-1}\in F$. ∎
Definition 6.4.
Given Assumption 6.1, we denote the division ring $(F,+,-,\circ,0,1)$ just constructed by $\mathbb{F}_{{\mathbf{D}},\varphi,T}$.
Corollary 6.5.
Given Assumption 6.1, let $\mathbb{F}:=\mathbb{F}_{{\mathbf{D}},\varphi,T}$. For each $e\in T$, the group $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$ becomes a left $\mathbb{F}$-vector space $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$ via the action $\lambda\cdot a:=\lambda(a)$.
6.2. The general case
In this subsection we revoke Assumption 6.1. We will instead assume only that $\mathcal{V}$ is a variety with a weak difference term, $\mathbf{A}$} is an algebra in $\mathcal{V}$, and $\theta$ is an abelian minimal congruence of $\mathbf{A}$}. Let ${\mathbf{D}}=D({\mathbf{A}},\theta)$ be the difference algebra for $\theta$, let $\varphi=\overline{\alpha}/\Delta$ be the derived congruence, where $\alpha=(0:\theta)$ and $\Delta=\Delta_{\theta,\alpha}$, and let $D^{o}=\{0_{E}\,:\,\text{$E\in A/\alpha$}\}$ be the canonical transversal. By Corollary 4.8, $\mathbf{D}$} is subdirectly irreducible, $\varphi$ is its monolith, $\varphi$ is abelian, and $D^{o}$ is a transversal for $\varphi$ which is a subuniverse of $\mathbf{D}$}. Thus we can appeal to the previous subsection to obtain the division ring $\mathbb{F}_{{\mathbf{D}},\varphi,D^{o}}$. In this section we will explain how $\mathbb{F}_{{\mathbf{D}},\varphi,D^{o}}$ acts naturally on each group $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ for any $e\in A$. The next lemma explains the crucial fact needed.
Lemma 6.6.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is an abelian minimal congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. Let ${\mathbf{D}}:=D({\mathbf{A}},\theta)$ be the difference algebra for $\theta$, let $\varphi$ be the derived congruence, let $D^{o}$ be the canonical transversal, and let $\mathbb{F}:=\mathbb{F}_{{\mathbf{D}},\varphi,D^{o}}$ be the division ring defined in the previous subsection. Let $E$ be an $\alpha$-class. For every $\theta$-class $C$ contained in $E$, $\operatorname{ran}(C)$ is an $\mathbb{F}$-subspace of the vector space $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{E})$.
Proof.
Let $\Delta=\Delta_{\theta,\alpha}$. Fix an $\alpha$-class $E$ and a $\theta$-class $C\subseteq E$. By Definition 5.4, we only need to show that $\operatorname{ran}(C)$ is closed under the action of $\mathbb{F}$. Let $a,e\in C$, so $q:=(a,e)/\Delta$ is a typical element of $\operatorname{ran}(C)$. Let $\lambda\in\mathbb{F}$. If $a=e$, then $q=0_{E}\in D^{o}$ and so $\lambda(q)=q\in\operatorname{ran}(C)$ as required. For the remainder of the proof, assume $a\neq e$.
Choose $(b,u)\in\theta$ with $\lambda((a,e)/\Delta)=(b,u)/\Delta$ and let $C_{0}=u/\theta$. By Corollary 5.8, we can choose a $\theta$-class $C^{\prime}$ with $\operatorname{ran}(C)\cup\operatorname{ran}(C_{0})\subseteq\operatorname{ran}(C^{\prime})$. Fix $e^{\prime}\in C^{\prime}$ and recall that $\operatorname{ran}(C^{\prime})=\operatorname{ran}(\chi_{e^{\prime}})$. Hence there exist $a^{\prime},b^{\prime}\in C^{\prime}$ with $(a^{\prime},e^{\prime})/\Delta=(a,e)/\Delta=q$ and $(b^{\prime},e^{\prime})/\Delta=(b,u)/\Delta=\lambda(q)$.
Since $(a,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},e^{\prime})$ and $a\neq e$, we get $a^{\prime}\neq e^{\prime}$ by Theorem 4.6(1). Thus $(a,e)\in\operatorname{Cg}(a^{\prime},e^{\prime})$ by minimality of $\theta$, and so Corollary 3.2 gives a polynomial $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(a^{\prime})=a$ and $f(e^{\prime})=e$. Observe that $f(C^{\prime})\subseteq C$.
Let $b^{\ast}=f(b^{\prime})\in C$. We will show $\lambda(q)=(b^{\ast},e)/\Delta$, which will finish the proof. Choose a term $t(x,\mathbf{y})$ and parameters $c_{1},\ldots,c_{n}\in A$ so that $f(x)=t(x,\mathbf{c})$. For each $i\in[n]$ let $E_{i}=c_{i}/\alpha$. Then calculating in $\mathbf{D}$},
$t(q,0_{E_{1}},\ldots,0_{E_{n}})$ $=t((a^{\prime},e^{\prime})/\Delta,(c_{1},c_{1})/\Delta,\ldots,(c_{n},c_{n})/\Delta)$
$=(t(a^{\prime},\mathbf{c}),t(e^{\prime},\mathbf{c}))/\Delta$
$=(a,e)/\Delta=q.$
Recall that $\lambda\in\mathbb{F}$; in particular, $\lambda\in\operatorname{End}({\mathbf{D}})$ and $\lambda(0_{E^{\prime}})=0_{E^{\prime}}$ for every $\alpha$-class $E^{\prime}$. Thus
$\lambda(q)$ $=\lambda(t(q,0_{E_{1}},\ldots,0_{E_{n}}))$
$=t(\lambda(q),0_{E_{1}},\ldots,0_{E_{n}})$
$=t((b^{\prime},e^{\prime})/\Delta,(c_{1},c_{1})/\Delta,\ldots,(c_{n},c_{n})/\Delta)$
$=(t(b^{\prime},\mathbf{c}),t(e^{\prime},\mathbf{c}))/\Delta$
$=(b^{\ast},e)/\Delta$
as required. ∎
As a consequence, in the context of Lemma 6.6, we can “translate” the $\mathbb{F}$-vector space action of the vector space $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{E})$ to the abelian groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ via $\chi_{e}$ and $\chi_{e}^{-1}$, for each $e\in E$.
Corollary 6.7.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is an abelian minimal congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. Let ${\mathbf{D}}:=D({\mathbf{A}},\theta)$ be the difference algebra for $\theta$, let $\varphi$ be the derived congruence, let $D^{o}$ be the canonical transversal, and let $\mathbb{F}:=\mathbb{F}_{{\mathbf{D}},\varphi,D^{o}}$. For every $\theta$-class $C$ and every $e\in C$, the rule
$\lambda\cdot a:=(\chi_{e}^{-1}\circ\lambda\circ\chi_{e})(a),\quad\lambda\in\mathbb{F},~a\in C$
defines a left $\mathbb{F}$-vector space action on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. With respect to this action, $\chi_{e}$ is a left $\mathbb{F}$-vector space embedding of $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ into $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{C})$.
Definition 6.8.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\theta$ is an abelian minimal congruence of $\mathbf{A}$}.
  1. (1)
    The division ring associated to $\theta$ is the division ring $\mathbb{F}_{\theta}:=\mathbb{F}_{{\mathbf{D}},\varphi,D^{o}}$ specified in the previous corollary.
  2. (2)
    For each $e\in A$, the canonical action of $\mathbb{F}_{\theta}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ is the action defined in the previous corollary.
The following result will be useful.
Proposition 6.9.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$, and $\theta\in\operatorname{Con}({\mathbf{A}})$, $\kappa\in\operatorname{Con}({\mathbf{B}})$ are such that $0\prec\theta$, $0\prec\kappa$, and $\theta,\kappa$ are both abelian. If $D({\mathbf{A}},\theta)\cong D({\mathbf{B}},\kappa)$, then $\mathbb{F}_{\theta}\cong\mathbb{F}_{\kappa}$.
Proof.
Let $\varphi_{{\mathbf{A}}}$ be the monolith of $D({\mathbf{A}},\theta)$, let $\varphi_{{\mathbf{B}}}$ be the monolith of $D({\mathbf{B}},\kappa)$, and let $\psi:D({\mathbf{A}},\theta)\cong D({\mathbf{B}},\kappa)$ be an isomorphism. Let $D^{o}_{{\mathbf{A}}}$ and $D^{o}_{{\mathbf{B}}}$ be the canonical transversals for $\varphi_{{\mathbf{A}}}$ and $\varphi_{{\mathbf{B}}}$ respectively. Clearly $\psi(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$ since the monolith of any subdirectly irreducible algebra is a characteristic congruence (i.e., preserved by all automorphisms). Hence $\psi(D^{o}_{{\mathbf{A}}})$ is a transversal for $\varphi_{{\mathbf{B}}}$. By Lemma 3.10, there exists an automorphism $\sigma\in\operatorname{Aut}(D({\mathbf{B}},\kappa))$ satisfying $\sigma(\psi(D^{o}_{{\mathbf{A}}}))=D^{o}_{{\mathbf{B}}}$. Thus $\sigma\circ\psi:(D({\mathbf{A}},\theta),\varphi_{{\mathbf{A}}},D^{o}_{{\mathbf{A}}})\cong(D({\mathbf{B}},\kappa),\varphi_{{\mathbf{B}}},D^{o}_{{\mathbf{B}}})$. Since these triples are exactly the data used to define $\mathbb{F}_{\theta}$ and $\mathbb{F}_{\kappa}$, it follows that the map $\lambda\mapsto(\sigma\circ\psi)\circ\lambda\circ(\sigma\circ\psi)^{-1}$ is an isomorphism from $\mathbb{F}_{\theta}$ to $\mathbb{F}_{\kappa}$. ∎
Note that if $\gamma,\theta\in\operatorname{Con}({\mathbf{A}})$ with $\gamma\prec\theta$ and $\theta/\gamma$ abelian, then the construction above provides a division ring $\mathbb{F}_{\theta/\gamma}$ associated to the abelian minimal congruence $\theta/\gamma$ of ${\mathbf{A}}/\gamma$. Thus every abelian cover pair of congruences (of an algebra in a variety with a weak difference term) can be assigned a division ring. It turns out that the division rings assigned to perspective abelian cover pairs in this way are isomorphic.
Theorem 6.10.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\gamma,\theta,\delta,\varepsilon\in\operatorname{Con}({\mathbf{A}})$ with $\gamma\prec\theta$ and $\delta\prec\varepsilon$, $\theta/\gamma$ and $\varepsilon/\delta$ are abelian, and $(\gamma,\theta)\nearrow(\delta,\varepsilon)$. Let $\alpha=(\gamma:\theta)$.
  1. (1)
    The rule
    $(a/\gamma,b/\gamma)/\Delta_{\theta/\gamma,\alpha/\gamma}\mapsto(a/\delta,b/\delta)/\Delta_{\varepsilon/\delta,\alpha/\delta}\quad\mbox{for}~(a,b)\in\theta$
    is a well-defined isomorphism $D({\mathbf{A}}/\gamma,\theta/\gamma)\cong D({\mathbf{A}}/\delta,\varepsilon/\delta)$.
  2. (2)
    $\mathbb{F}_{\theta/\gamma}\cong\mathbb{F}_{\varepsilon/\delta}$.
Proof.
Observe that $(0:\theta/\gamma)=\alpha/\gamma$ by Lemma 2.3(1), so we may assume with no loss of generality that $\gamma=0$. Let $\overline{{\mathbf{A}}}$, $\overline{\varepsilon}$ and $\overline{\alpha}$ denote ${\mathbf{A}}/\delta$, $\varepsilon/\delta$ and $\alpha/\delta$ respectively. Note that $(0:\overline{\varepsilon})=\overline{\alpha}$ by Lemma 2.3(1) and Proposition 3.14, so $D(\overline{{\mathbf{A}}},\overline{\varepsilon})=\overline{{\mathbf{A}}}(\overline{\varepsilon})/\Delta_{\overline{\varepsilon},\overline{\alpha}}$.
(1) Let $\psi:{\mathbf{A}}(\theta)\rightarrow\overline{{\mathbf{A}}}(\overline{\varepsilon})$ be the homomorphism given by by $\psi((a,b))=(a/\delta,b/\delta)$, let $\nu:\overline{{\mathbf{A}}}(\overline{\varepsilon})\rightarrow D(\overline{{\mathbf{A}}},\overline{\varepsilon})$ be the natural projection map, and let $\Psi=\nu\circ\psi$. Since $\varepsilon=\delta\circ\theta\circ\delta$ by Proposition 3.14, $\psi$ is surjective, so $\Psi$ is also surjective. We must show that $\ker(\Psi)=\Delta_{\theta,\alpha}$. Let $\Delta=\Delta_{\theta,\alpha}$. Because $\varepsilon=\delta\circ\theta\circ\delta$, we have
(6.3) $M(\overline{\varepsilon},\overline{\alpha})=\left\{\begin{pmatrix}a/\delta&a^{\prime}/\delta\\ b/\delta&b^{\prime}/\delta\end{pmatrix}:\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in M(\theta,\alpha)\right\}.$
Hence $\Delta\subseteq\ker(\Psi)$. Before proving the opposite inclusion, we establish the following.
Claim. If $(a,b),(r,s)\in\theta$ and $(a,r),(b,s)\in\delta$, then $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r,s)$.
Proof of Claim.
Let $d$ be a weak difference term for $\mathcal{V}$. Note first that $d(a,r,r)\stackrel{{\scriptstyle\theta}}{{\equiv}}d(b,s,r)$ and $d(a,r,r)\stackrel{{\scriptstyle\delta}}{{\equiv}}d(a,a,a)=a=d(b,b,a)\stackrel{{\scriptstyle\delta}}{{\equiv}}d(b,s,r)$. Since $\theta\cap\delta=0$, this proves
(6.4) $d(a,r,r)=d(b,s,r).$
Next, note that
(6.5) $d(\underline{r},r,r)=r=d(\underline{r},s,s),$
so since $(r,s)\in\theta$ and $(r,b)\in\delta\vee\theta=\varepsilon\leq\alpha$ and $C(\alpha,\theta;0)$, we can replace the underlined occurrences of “rr” in equation 6.5 by “bb” to get
(6.6) $d(b,r,r)=d(b,s,s).$
Now we build some $(\theta,\alpha)$-matrices. Let
$M_{1}=\begin{pmatrix}a&a\\ b&b\end{pmatrix},~~M_{2}=\begin{pmatrix}a&r\\ a&r\end{pmatrix},~~M_{3}=\begin{pmatrix}b&s\\ b&s\end{pmatrix},~~M_{4}=\begin{pmatrix}s&s\\ s&s\end{pmatrix},~~M_{5}=\begin{pmatrix}r&r\\ s&s\end{pmatrix}$.
Clearly $M_{1},\ldots,M_{5}\in M(\theta,\alpha)$. Applying the weak difference term gives
$N:=d(M_{1},M_{2},M_{2})$ $=\begin{pmatrix}a&d(a,r,r)\\ b&d(b,r,r)\end{pmatrix}\in M(\theta,\alpha)$
$N^{\prime}:=d(M_{3},M_{4},M_{5})$ $=\begin{pmatrix}d(b,s,r)&r\\ d(b,s,s)&s\end{pmatrix}\in M(\theta,\alpha).$
The second column of $N$ and the first column of $N^{\prime}$ are identical by equations (6.4) and (6.6). Since $\Delta$ is the “horizontal transitive closure” of $M(\theta,\alpha)$, this proves $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r,s)$, proving the Claim. ∎
Now we prove $\ker(\Psi)\subseteq\Delta$. Assume $((a,b),(a^{\prime},b^{\prime}))\in\ker(\Psi)$; that is,
$(a,b),(a^{\prime},b^{\prime})\in\theta$ and $(a/\delta,b/\delta)\stackrel{{\scriptstyle\Delta_{\overline{\varepsilon},\overline{\alpha}}}}{{\equiv}}(a^{\prime}/\delta,b^{\prime}/\delta)$.
By definition of $\Delta_{\overline{\varepsilon},\overline{\alpha}}$ and equation 6.3, there exist $n\geq 0$ and
(6.7) $\begin{pmatrix}r_{i}&a_{i}\\ s_{i}&b_{i}\end{pmatrix}\in M(\theta,\alpha)\quad\mbox{for $i\leq n$}$
such that $(a/\delta,b/\delta)=(r_{0}/\delta,s_{0}/\delta)$, $(a_{i}/\delta,b_{i}/\delta)=(r_{i+1}/\delta,s_{i+1}/\delta)$ for all $i<n$, and $(a^{\prime}/\delta,b^{\prime}/\delta)=(a_{n}/\delta,b_{n}/\delta)$. The Claim then gives $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r_{0},s_{0})$, $(a_{i},b_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r_{i+1},s_{i+1})$ for $i<n$, and $(a_{n},b_{n})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},b^{\prime})$. As equation 6.7 gives $(r_{i},s_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a_{i},b_{i})$ for $i<n$, we get $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},b^{\prime})$ by transitivity as required.
(2) follows from (1) and Proposition 6.9. ∎
7. Similarity
In his 1983 paper [5], Freese introduced a relation of similarity between subdirectly irreducible algebras in a congruence modular variety. If the monoliths of the algebras are nonabelian, then the algebras are similar if and only if they are isomorphic. If the monoliths are abelian, then the relation is more interesting. Freese’s original definition in the latter case, as well as the equivalent formulation found in [6], relied heavily on having a difference term, and does not easily adapt to varieties with a weak difference term.
Happily, Freese gave two other equivalent formulations of the similarity relation in the abelian monolith case, one of which is based on the construction in Section 4. Both formulations generalize nicely to varieties with a weak difference term. We develop this generalization in this section.
The following definition, in the context of congruence modular varieties, is found in [6, Definition 10.9].
Definition 7.1.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}}\in\mathcal{V}$ is subdirectly irreducible with monolith $\mu$. The algebra $D({\mathbf{A}})$ is defined as follows:
  1. (1)
    If $\mu$ is nonabelian, then $D({\mathbf{A}})={\mathbf{A}}$.
  2. (2)
    If $\mu$ is abelian, then $D({\mathbf{A}})$ is $D({\mathbf{A}},\mu)$, the difference algebra for $\mu$.
Observe that, in the context of the previous definition, $D({\mathbf{A}})$ is a subdirectly irreducible algebra in $\mathsf{HS}({\mathbf{A}}^{2})$; furthermore, the monoliths of $\mathbf{A}$} and $D({\mathbf{A}})$ are either both abelian or both nonabelian.
Definition 7.2.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible. We say that $\mathbf{A}$} and $\mathbf{B}$} are similar, and write ${\mathbf{A}}\sim{\mathbf{B}}$, if $D({\mathbf{A}})\cong D({\mathbf{B}})$.
By [5, Theorem 5.1] and [6, Theorem 10.11], our definition of similarity, when restricted to the congruence modular setting, is equivalent to the original definition given in [5] and [6]. It is in this sense that our definition extends the original one.
If we let $\mu$ and $\kappa$ denote the respective monoliths of $\mathbf{A}$} and $\mathbf{B}$}, then Definition 7.2 can be rephrased as follows: ${\mathbf{A}}\sim{\mathbf{B}}$ if and only if one of the following holds:
  1. (1)
    $\mu$ and $\kappa$ are both nonabelian and ${\mathbf{A}}\cong{\mathbf{B}}$; or
  2. (2)
    $\mu$ and $\kappa$ are both abelian and $D({\mathbf{A}},\mu)\cong D({\mathbf{B}},\kappa)$.
Before stating the main theorem of this section, we introduce a variant of a crucial construct due to Zhuk [35, Section 6.6].
Definition 7.3.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are subdirectly irreducible algebras in a common signature with monoliths $\mu,\kappa$ respectively. A similarity bridge from $\mathbf{A}$} to $\mathbf{B}$} is a subuniverse $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ satisfying
  1. (B1)
    $\operatorname{pr}_{1,2}(T)=\mu$ and $\operatorname{pr}_{3,4}(T)=\kappa$.
  2. (B2)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $a_{1}=a_{2}$ if and only if $b_{1}=b_{2}$.
  3. (B3)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $(a_{i},a_{i},b_{i},b_{i})\in T$ for $i=1,2$.
If $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}, then we define its trace to be the set $\operatorname{tr}(T)=\{(a,b)\in A\times B\,:\,\text{$(a,a,b,b)\in T$}\}$, and its kernel to be the set
$\ker(T)=\{((a_{1},b_{1}),(a_{2},b_{2}))\,:\,\text{$(a_{1},a_{2},b_{1},b_{2})\in T$}\}.$
Observe that for any similarity bridge $T$ we have $\operatorname{tr}(T)=\operatorname{pr}_{1,3}(T)=\operatorname{pr}_{2,4}(T)$ by property (B3), and hence $\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ by property (B1).
In the special case where one of the two monoliths in the previous definition is nonabelian, the existence of a similarity bridge is possible only in the most trivial way.
Proposition 7.4.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}. Assume that one of $\mu$ or $\kappa$ is nonabelian. Then so is the other, $\operatorname{tr}(T)$ is the graph of an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$, and $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$.
Proof.
Assume with no loss of generality that $\kappa$ is nonabelian. We will first show that $\operatorname{tr}(T)$ is the graph of a surjective homomorphism $h:{\mathbf{A}}\rightarrow{\mathbf{B}}$. Let
$S=\{(b,c)\in B^{2}\,:\,\text{$\exists a\in A$ with $(a,b),(a,c)\in\operatorname{tr}(T)$}\}.$
$S$ is a tolerance of $\mathbf{B}$}. Let
$Y=\left\{\begin{pmatrix}b&c\\ b^{\prime}&c^{\prime}\end{pmatrix}\in B^{2\times 2}~:~\mbox{$\exists\,(a,a^{\prime})\in\mu$ with $(a,a^{\prime},b,b^{\prime}),(a,a^{\prime},c,c^{\prime})\in T$}\right\}.$
One can easily show, using (B1) and the definitions of $S$ and $Y$, that $X(\kappa,S)\subseteq Y$ (see Definition 2.1). Since $Y$ is a subuniverse of ${\mathbf{B}}^{2\times 2}$, we get $M(\kappa,S)\subseteq Y$. Then from (B2) and Definition 2.2(2) we can deduce $C(S,\kappa;0_{B})$. Since $S$ is a tolerance, we can improve this to $C(\operatorname{Cg}(S),\kappa;0_{B})$ by [16, Theorem 2.19(2)]. Now assume that $S\neq 0_{B}$; then $\operatorname{Cg}(S)\geq\kappa$ since $\kappa$ is the monolith of $\mathbf{B}$}; but then $\kappa$ would be abelian, contrary to our assumption. This proves $S=0_{B}$ and hence $\operatorname{tr}(T)$ is the graph of a surjective homomorphism $h:{\mathbf{A}}\rightarrow{\mathbf{B}}$.
Assume that $h$ is not injective. Then $\mu\subseteq\ker(h)$ as $\mu$ is the monolith of $\mathbf{A}$}. Choose $(a,b)\in\mu$ with $a\neq b$. By (B1), there exists $(c,d)\in\kappa$ with $(a,b,c,d)\in T$. Then $c\neq d$ by (B2). By (B3), we have $(a,c),(b,d)\in\operatorname{tr}(T)$, or equivalently, $c=h(a)$ and $d=h(b)$. But then $(a,b)\in\mu\subseteq\ker(h)$ implies $c=d$, contradiction. Thus $h$ is an isomorphism.
Finally, for every $(a,b)\in\mu$ there is at least one pair $(c,d)\in\kappa$ with $(a,b,c,d)\in T$ by (B1). On the other hand, the only pair $(c,d)$ that can work is $(c,d)=(h(a),h(b))$ by (B3) and the definition of $h$. Hence $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$. ∎
Lemma 7.5.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}. Let ${\mathbf{C}}=\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and $\tau=\ker(T)$. Then $\tau\in\operatorname{Con}({\mathbf{C}})$. Hence
  1. (B4)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $(a_{2},a_{1},b_{2},b_{1})\in T$.
  2. (B5)
    If $(a_{1},a_{2},b_{1},b_{2})\in T$ and $(a_{2},a_{3},b_{2},b_{3})\in T$, then $(a_{1},a_{3},b_{1},b_{3})\in T$.
Proof.
Note that $\tau$ is a reflexive subuniverse of ${\mathbf{C}}^{2}$ by (B3), and that (B4) and (B5) simply express the symmetry and transitivity of $\tau$. Hence all we need to prove is that $\tau$ is symmetric and transitive.
If $\mu$ or $\kappa$ is nonabelian, then $C$ is the graph of an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$ and $\tau=\{((a,h(a)),(b,h(b))\,:\,\text{$(a,b)\in\mu$}\}$ by Proposition 7.4. Clearly $\tau$ is symmetric and transitive in this case.
If $\mu$ and $\kappa$ are both abelian, let $d(x,y,z)$ be a weak difference term for $\mathcal{V}$. Let
$\theta=\{((a_{1},b_{1}),(a_{2},b_{2}))\in C^{2}\,:\,\text{$(a_{1},a_{2})\in\mu$ and $(b_{1},b_{2})\in\kappa$}\}.$
Clearly $\theta\in\operatorname{Con}({\mathbf{C}})$. Because $\mu$ and $\kappa$ are abelian, the restriction of $d$ to any $\theta$-class is a Maltsev operation. We have that $\tau$ is a reflexive subuniverse of ${\mathbf{C}}^{2}$ and $\tau\subseteq\theta$ by (B1). Thus $\tau$ is symmetric and transitive by Lemma 3.1. ∎
Lemma 7.6.
[…]
Proof.
Clearly $\tau\leq\delta^{+}_{1}\wedge\delta^{+}_{2}$ by (B1). Choose $(a_{1},a_{2})\in\mu$ with $a_{1}\neq a_{2}$. Again by (B1), there exist $b_{1},b_{2}\in B$ with $(a_{1},a_{2},b_{1},b_{2})\in T$ and hence $(a_{1},b_{1})\stackrel{{\scriptstyle\tau}}{{\equiv}}(a_{2},b_{2})$, which proves $\tau\nleq\delta_{1}$. A similar argument shows $\tau\nleq\delta_{2}$. This proves $\tau\vee\delta_{i}=\delta^{+}_{i}$ for $i=1,2$.
[…]
Here is the main theorem of this section. The equivalence of items (1)–(3) was proved for congruence modular varieties by Freese [5, Theorems 2.1 and 5.1]; see also [6, Theorems 10.8 and 10.11].
Theorem 7.7.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible. The following are equivalent:
  1. (1)
    ${\mathbf{A}}\sim{\mathbf{B}}$.
  2. (2)
[…]
  3. (3)
    There exist an algebra ${\mathbf{C}}\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and a congruence $\tau\in\operatorname{Con}({\mathbf{C}})$ which, with the projection homomorphisms $f_{1}=\operatorname{pr}_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}=\operatorname{pr}_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$ and $\psi:=0_{C}$, satisfy the conclusion of item (2).
  4. (4)
    There exists a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}.
Proof.
Throughout this proof, let $\mu,\kappa$ denote the monoliths of ${\mathbf{A}},{\mathbf{B}}$ respectively. The implication (3) $\Rightarrow$ (2) is clear, and (4) $\Rightarrow$ (3) holds by Lemma 7.6. We will prove (2) $\Rightarrow$ (1) $\Rightarrow$ (4) first in the case when one of $\mu$ or $\kappa$ is nonabelian, and then in the case when both $\mu$ and $\kappa$ are abelian.
Case 1: $\mu$ or $\kappa$ is nonabelian.
(1) $\Rightarrow$ (4). Since ${\mathbf{A}}\sim{\mathbf{B}}$, it follows that both $\mu$ and $\kappa$ are nonabelian and there exists an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$. Then the set $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}.
[…]
For each $i=1,2$ we have $\delta_{i}\wedge\tau=\psi$ and hence $C(\delta_{i},\tau;\psi)$; see e.g. [16, Theorem 2.19(8)]; thus $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ by semidistributivity of the centralizer relation in the first variable; see e.g. [16, Theorem 2.19(5)]. Suppose $\delta_{1}\neq\delta_{2}$; then without loss of generality we can assume $\delta_{1}\vee\delta_{2}>\delta_{2}$. Hence $\delta_{1}\vee\delta_{2}\geq\delta^{+}_{2}$ (as $\delta_{2}$ is completely meet-irreducible) and thus $\delta_{1}\vee\delta_{2}\geq\tau$. Then $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ implies $C(\tau,\tau;\psi)$ by monotonicity, which contradicts the fact that $\tau/\psi$ is nonabelian. This proves $\delta_{1}=\delta_{2}$ and hence ${\mathbf{A}}\cong{\mathbf{C}}/\delta_{1}={\mathbf{C}}/\delta_{2}\cong{\mathbf{B}}$. Hence ${\mathbf{A}}\sim{\mathbf{B}}$.
Case 2: $\mu$ and $\kappa$ are abelian.
Thus $D({\mathbf{A}})=D({\mathbf{A}},\mu)$ and $D({\mathbf{B}})=D({\mathbf{B}},\kappa)$. For the remainder of this proof, let $\alpha=(0:\mu)$, $\Delta_{{\mathbf{A}}}=\Delta_{\mu,\alpha}$, and $\varphi_{{\mathbf{A}}}=\overline{\alpha}_{{\mathbf{A}}}/\Delta_{{\mathbf{A}}}$. Also let $D^{o}_{{\mathbf{A}}}\leq D({\mathbf{A}})$ as given in the proof of Corollary 4.8 for $({\mathbf{A}},\mu)$. Similarly define $\beta:=(0:\kappa)$, $\Delta_{{\mathbf{B}}}$, $\varphi_{{\mathbf{B}}}$, and $D^{o}_{{\mathbf{B}}}$ for $({\mathbf{B}},\kappa)$.
(2) $\Rightarrow$ (1) We have $D({\mathbf{A}})\cong D({\mathbf{C}}/\delta_{1},\delta_{1}^{+}/\delta_{1})$ and $D({\mathbf{B}})\cong D({\mathbf{C}}/\delta_{2},\delta_{2}^{+}/\delta_{2})$ since ${\mathbf{A}}\cong{\mathbf{C}}/\delta_{1}$ and ${\mathbf{B}}\cong{\mathbf{C}}/\delta_{2}$. We can replace $\tau$ with any $\tau^{\prime}\in\operatorname{Con}({\mathbf{C}})$ satisfying $\psi\prec\tau^{\prime}\leq\tau$ and the hypotheses of (2) still hold; thus we may assume $\psi\prec\tau$. We have that $\tau/\psi$ is abelian, since $(\psi,\tau)\nearrow(\delta_{1},\delta_{1}^{+})$ and $\delta_{1}^{+}/\delta_{1}$ is abelian. Then by Theorem 6.10 we get $D({\mathbf{C}}/\delta_{1},\delta_{1}^{+}/\delta_{1})\cong D({\mathbf{C}}/\psi,\tau/\psi)\cong D({\mathbf{C}}/\delta_{2},\delta_{2}^{+}/\delta_{2})$. Hence $D({\mathbf{A}})\cong D({\mathbf{B}})$, so ${\mathbf{A}}\sim{\mathbf{B}}$.
(1) $\Rightarrow$ (4). Assume $D({\mathbf{A}})\cong D({\mathbf{B}})$. As in the proof of Proposition 6.9, there exists an isomorphism $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$ satisfying $\lambda(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$ and $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$. Define
$T=\{(a_{1},a_{2},b_{1},b_{2})\in A\times A\times B\times B$ $:~\mbox{$a_{1}\stackrel{{\scriptstyle\mu}}{{\equiv}}a_{2}$ and $b_{1}\stackrel{{\scriptstyle\kappa}}{{\equiv}}b_{2}$ and}$
[…]
Because $\lambda$ is an isomorphism, we get $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ with $\operatorname{pr}_{1,2}(T)=\mu$ and $\operatorname{pr}_{3,4}(T)=\kappa$, proving (B1). Let $(a_{1},a_{2},b_{1},b_{2})\in T$ and assume $a_{1}=a_{2}$. Then $(a_{1},a_{2})/\Delta_{{\mathbf{A}}}\in D^{o}_{{\mathbf{A}}}$, As $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$, we get $(b_{1},b_{2})/\Delta_{{\mathbf{B}}}=\lambda((a_{1},a_{2})/\Delta_{{\mathbf{A}}})\in D^{o}_{{\mathbf{B}}}$, so $b_{1}=b_{2}$ (see Theorem 4.6(1)). A similar argument shows that if $b_{1}=b_{2}$, then $a_{1}=a_{2}$. This proves (B2).
Again suppose $(a_{1},a_{2},b_{1},b_{2})\in T$ and let $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}$. Observe that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{A}}}}}{{\equiv}}(a_{1},a_{2})/\Delta_{{\mathbf{A}}}$ and $(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}(b_{1},b_{2})/\Delta_{{\mathbf{B}}}$. Then
$r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ $\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}\lambda((a_{1},a_{2})/\Delta_{{\mathbf{A}}})$ as $\lambda(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$
$=(b_{1},b_{2})/\Delta_{{\mathbf{B}}}$ as $(a_{1},a_{2},b_{1},b_{2})\in T$
$\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}(b_{1},b_{1})/\Delta_{{\mathbf{B}}}=s,$
i.e., $(r,s)\in\varphi_{{\mathbf{B}}}$. Also recall that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\in D^{o}_{{\mathbf{A}}}$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\in D^{o}_{{\mathbf{B}}}$, and since $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$ we also get $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})\in D^{o}_{{\mathbf{B}}}$. So $r,s\in D^{o}_{{\mathbf{B}}}$ and $(r,s)\in\varphi_{{\mathbf{B}}}$. But $D^{o}_{{\mathbf{B}}}$ is a transversal for $\varphi_{{\mathbf{B}}}$. These facts imply $r=s$, which implies $(a_{1},a_{1},b_{1},b_{1})\in T$. A similar argument gives $(a_{2},a_{2},b_{2},b_{2})\in T$. This proves (B3). Thus $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}. ∎
Item (1) in the next result extends the first claim of [6, Theorem 10.10] from congruence modular varieties to varieties with a weak difference term. (The remainder of [6, Theorem 10.10] was extended by Corollary 4.8.) Item (3) extends a fact established, in the congruence modular setting, within the proof of [6, Theorem 10.11].
Corollary 7.8.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}}\in\mathcal{V}$ is subdirectly irreducible with abelian monolith $\mu$.
  1. (1)
    ${\mathbf{A}}\sim D({\mathbf{A}})$.
  2. (2)
    Setting $\alpha=(0:\mu)$ and $\Delta=\Delta_{\mu,\alpha}$, the set
    $T_{{\mathbf{A}}}^{D}:=\{(a,b,(a,e)/\Delta,(b,e)/\Delta)\,:\,\text{$a\stackrel{{\scriptstyle\mu}}{{\equiv}}b\stackrel{{\scriptstyle\mu}}{{\equiv}}e$}\}$
    is a similarity bridge from $\mathbf{A}$} to $D({\mathbf{A}})$.
  3. (3)
    If $(0:\mu)=\mu$ and there exists a subuniverse $A_{0}\leq{\mathbf{A}}$ which is a transversal for $\mu$, then ${\mathbf{A}}\cong D({\mathbf{A}})$.
Hence each similarity class in $\mathcal{V}$ consisting of subdirectly irreducible algebras with abelian monoliths contains a unique (up to isomorphism) subdirectly irreducible algebra $\mathbf{S}$} with abelian monolith $\varphi$ satisfying (i) $(0:\varphi)=\varphi$, and (ii) $\mathbf{S}$} has a subuniverse which is a transversal for $\varphi$.
Proof.
(1) Let $\eta_{1},\eta_{2}$ be the kernels of the two projection maps $\operatorname{pr}_{1},\operatorname{pr}_{2}:{\mathbf{A}}(\mu)\rightarrow{\mathbf{A}}$. By Theorem 4.6, if we set ${\mathbf{B}}=D({\mathbf{A}})$, ${\mathbf{C}}={\mathbf{A}}(\mu)$, $\delta_{1}=\eta_{1}$, $\delta_{2}=\Delta$, $\delta^{+}_{1}=\overline{\mu}$, $\delta^{+}_{2}=\overline{\alpha}$, $\psi=0$, and $\tau=\eta_{2}$, then the conditions of Theorem 7.7(2) are met, showing ${\mathbf{A}}\sim D({\mathbf{A}})$.
(2) This can be deduced from the above proof of item (1), following the proof of Theorem 7.7 (2) $\Rightarrow$ (4) in the abelian monolith case. Instead, we give a direct proof. Let $\varphi=\overline{\alpha}/\Delta$. Clearly $\operatorname{pr}_{1,2}(T_{{\mathbf{A}}}^{D})=\mu$, $\operatorname{pr}_{3,4}(T_{{\mathbf{A}}}^{D})\subseteq\varphi$, (B3) holds, and the ($\Rightarrow$) direction of (B2) holds. To prove the ($\Leftarrow$) direction of (B2), assume that $a,b,e$ belong to a common $\mu$ class and $(a,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e)$. Using the notation of Definition 5.1, we can express this as $\chi_{e}(a)=\chi_{e}(b)$. But $\chi_{e}$ is injective by Lemma 5.2, so $a=b$, completing the proof of (B2).
It remains to prove $\varphi\subseteq\operatorname{pr}_{3,4}(T_{{\mathbf{A}}}^{D})$. Let $((a_{1},a_{2})/\Delta,(b_{1},b_{2})/\Delta)\in\varphi$ be given; thus $(a_{1},a_{2}),(b_{1},b_{2})\in\mu$ and $a_{1},a_{2},b_{1},b_{2}$ belong to a common $\alpha$-class. By Lemma 5.7, there exists $e\in a_{1}/\alpha$ with $\operatorname{ran}(\chi_{a_{2}})\cup\operatorname{ran}(\chi_{b_{2}})\subseteq\operatorname{ran}(\chi_{e})$. As $(a_{1},a_{2})/\Delta\in\operatorname{ran}(\chi_{a_{2}})$ and $(b_{1},b_{2})/\Delta\in\operatorname{ran}(\chi_{b_{2}})$, this implies the existence of $a,b\in e/\mu$ with $\chi_{e}(a)=(a_{1},a_{2})/\Delta$ and $\chi_{e}(b)=(b_{1},b_{2})/\Delta$, which means
$((a_{1},a_{2})/\Delta,(b_{1},b_{2})/\Delta)=((a,e)/\Delta,(b,e)/\Delta)\in T_{{\mathbf{A}}}^{D}.$
(3) Since $(0:\mu)=\mu$, we have $D({\mathbf{A}})={\mathbf{A}}(\mu)/\Delta$ where $\Delta=\Delta_{\mu,\mu}$. Since $\mu$ is abelian, we get $\Delta=\left\{((a,b),(a^{\prime},b^{\prime})):{\scriptsize{\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}}}\in M(\mu,\mu)\right\}$ by Lemma 4.4(4). Let $\pi:{\mathbf{A}}\rightarrow{\mathbf{A}}_{0}$ be the homomorphism with $\pi{\upharpoonright}_{A_{0}}=\mathrm{id}_{A_{0}}$ and $\ker(\pi)=\mu$. Then define $\lambda:{\mathbf{A}}\rightarrow D({\mathbf{A}})$ by $\lambda(x)=(x,\pi(x))/\Delta$. Suppose $(a,b)\in\ker(\lambda)$. Then $((a,\pi(a),(b,\pi(b)))\in\Delta$, so $M:={\scriptsize{\begin{pmatrix}a&b\\ \pi(a)&\pi(b)\end{pmatrix}\in M(\mu,\mu)}}$. Also $(a,b)\in\mu$, so $\pi(a)=\pi(b)$. Thus the bottom entries of $M$ are equal. Since $\mu$ is abelian, it follows that the top entries of $M$ are also equal, i.e., $a=b$. Thus $\lambda$ is injective.
To prove that $\lambda$ is surjective, let $(a,b)\in\mu$ be given and define $e=\pi(a)=\pi(b)$ and $c=d(a,b,e)$ where $d(x,y,z)$ is a weak difference term for $\mathcal{V}$. Then $(a,b)/\Delta=(c,e)/\Delta=\lambda(c)$ as required. ∎
We end this section with two easy consequences of similarity in the abelian monolith case.
Proposition 7.9.
Suppose $\mathcal{V}$ is a variety with a weak difference term and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with abelian monoliths $\mu,\kappa$ respectively. Define $\alpha=(0:\mu)$ and $\beta=(0:\kappa)$, and let $\mathbb{F}_{\mu}$ and $\mathbb{F}_{\kappa}$ be the division rings associated to $({\mathbf{A}},\mu)$ and $({\mathbf{B}},\kappa)$ respectively. If ${\mathbf{A}}\sim{\mathbf{B}}$, then:
  1. (1)
    ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
  2. (2)
    $\mathbb{F}_{\mu}\cong\mathbb{F}_{\kappa}$.
Proof.
(1) Let $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$. Let $\varphi_{{\mathbf{A}}},\varphi_{{\mathbf{B}}}$ be the monoliths of $D({\mathbf{A}}),D({\mathbf{B}})$ respectively. Then $\lambda$ induces an isomorphism $\lambda^{\ast}:D({\mathbf{A}})/\varphi_{{\mathbf{A}}}\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$. We have ${\mathbf{A}}/\alpha\cong D({\mathbf{A}})/\varphi_{{\mathbf{A}}}$ and ${\mathbf{B}}/\beta\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ by Corollary 4.8, so ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
(2) follows from Proposition 6.9. ∎
8. Conclusion and questions
In this paper we have extended Freese’s analysis of abelian minimal congruences, and his similarity relation between subdirectly irreducible algebras, from the congruence modular setting to the setting of varieties with a weak difference term. We also established a connection between similarity and certain constructs which we have called “similarity bridges.” In the companion paper [31], we will compare similarity bridges to Zhuk’s original bridges (which are superficially weaker) and explain some related elements of Zhuk’s work in terms of centrality and similarity. In a third paper [30], we will unite the analyses of “rectangular critical relations” by Zhuk [35, Lemma 8.21] (in the idempotent locally finite Taylor case) and by Kearnes and Szendrei [19] (in the congruence modular case), showing in particular the role of similarity, associated finite fields, and linear equations.
In the congruence modular setting, Freese and McKenzie [6] established some nontrivial results concerning the similarity relation, such as bounding the number of similarity classes of subdirectly irreducible algebras in finitely generated varieties. One can ask whether any of these properties persist in varieties with a weak difference term, or with a difference term. For example:
Problem 1.
If $\mathbf{A}$} is a finite algebra and $\mathsf{HSP}({\mathbf{A}})$ has a weak difference term, does it follow that there are only finitely many similarity classes of subdirectly irreducible algebras in $\mathsf{HSP}({\mathbf{A}})$ with abelian monolith?
Finite subdirectly irreducible algebras with abelian monoliths in idempotent locally finite Taylor varieties are of particular interest. We have seen one result (Lemma 5.9) showing how the assumption of idempotency can restrict the possible sizes of monolith classes. In what other ways does idempotency affect the sizes of monolith classes?
Problem 2.
Suppose $\mathbf{A}$} is a finite idempotent subdirectly irreducible algebra in a Taylor variety with abelian monolith $\mu$, $\alpha=(0:\mu)$, and $E$ is an $\alpha$-class. If $q$ is the maximal size of a $\mu$-class in $E$, does it follow that every $\mu$-class in $E$ has size $q$ or 1?
When $\mathbf{A}$} is a finite algebra in a variety with a weak difference term and $\theta$ is an abelian minimal congruence, tame congruence theory tells us that $\operatorname{typ}_{{\mathbf{A}}}(0,\theta)=\mathbf{2}$ and therefore the minimal algebras ${\mathbf{A}}|_{N}$ associated to $\langle 0,\theta\rangle$-traces are weakly isomorphic to a one-dimensional vector space over a finite field $\mathbb{F}$ determined by $\theta$.
Problem 3.
Is the finite field $\mathbb{F}$ mentioned above isomorphic to $\mathbb{F}_{\theta}$? If not, how are $\mathbb{F}$ and $\mathbb{F}_{\theta}$ related?
Appendix 1: One division ring
Let $\mathbf{A}$} be an algebra in a variety with a weak difference term, let $\theta$ be an abelian minimal congruence, and let $\mathbb{F}_{\theta}$ be the division ring constructed in Section 6. In this appendix we will show that $\mathbb{F}_{\theta}$ is isomorphic to another division ring $\mathbb{D}$ associated to $\theta$, which Freese [5, Section 4, page 150] described in the congruence modular setting and whose construction easily adapts to algebras having a weak difference term. We also show that when $\mathbf{A}$} is finite, $\mathbb{F}_{\theta}$ is isomorphic to the finite field $\mathbb{F}$ arising in the proof sketch of Theorem 3.9.
First we describe our adaptation of Freese’s construction. Assume only that $\mathbf{A}$} is an algebra having a weak difference term $d$, and $\theta$ is an abelian minimal congruence. Let the $\theta$-classes be indexed by a set $I$ via the notation $A/\theta=\{C_{i}\,:\,\text{$i\in I$}\}$. Let $T$ be a transversal for $\theta$ and for each $i\in I$ let $e_{i}$ be the unique element of $T\cap C_{i}$. For $i,j\in I$ define
$R_{ij}=\{f{\upharpoonright}_{C_{j}}\,:\,\text{$f\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $f(e_{j})=e_{i}$}\}.$
(Freese denoted this set by $\operatorname{Hom}(\theta,e_{j},e_{i})$.) Also define $O_{ij}\in R_{ij}$ to be the constant map $C_{j}\rightarrow\{e_{i}\}$. Let $\mathcal{R}$ be the set of $I\times I$-matrices $(r_{ij})_{i,j\in I}$ where $r_{ij}\in R_{ij}$ for all $i,j\in I$, and for each $j\in I$, $r_{ij}=O_{ij}$ for all but finitely many $i\in I$. Let $V$ denote the direct sum of the abelian groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$, $i\in I$, realized as the subgroup of $\prod_{i\in I}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ consisting of those $(a_{i})_{i\in I}\in\prod_{i\in I}C_{i}$ for which $a_{j}=e_{j}$ for all but finitely many $j\in I$. Observe that each $f\in R_{ij}$ is a group homomorphism from $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{j})$ to $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$, by Proposition 3.8; that $R_{ij}$ is closed under the pointwise addition operation in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$; and that $f\in R_{ij}$, $g\in R_{jk}$ imply $f\circ g\in R_{ik}$.
Thus $\mathcal{R}$ acts in the usual “matrix” way on $V$: given $M=(r_{ij})_{i,j\in I}\in\mathcal{R}$, define $\mathsf{L}_{M}:V\rightarrow V$ so that $\mathsf{L}_{M}((a_{i})_{i\in I})=(b_{i})_{i\in I}$ where
$b_{i}=\sum_{j\in I,\,a_{j}\neq e_{j}}r_{ij}(a_{j})\quad\mbox{with the sum computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$}.$
One can show that each $\mathsf{L}_{M}\in\operatorname{End}(V)$, the map $M\mapsto\mathsf{L}_{M}$ is injective, and its range $\mathsf{L}_{\mathcal{R}}:=\{\mathsf{L}_{M}\,:\,\text{$M\in\mathcal{R}$}\}$ is a unital subring of $\mathbf{End}(V)$. Via the identification of $\mathcal{R}$ with $\mathsf{L}_{\mathcal{R}}$, $\mathcal{R}$ inherits the operations of a unital ring, and we obtain the faithful left $\mathcal{R}$-module $_{\mathcal{R}}V$. The minimality of $\theta$ implies that $_{\mathcal{R}}V$ is a simple module. Freese [5, p. 150] explained this in the congruence modular case; here is a proof that does not require congruence modularity. Let $\overline{0}=(e_{i})_{i\in I}\in V$. The claim of simplicity reduces to proving the following: if $\mathbf{a}=(a_{i})_{i\in I}\in V$ with $\mathbf{a}\neq\overline{0}$, and $\mathbf{b}=(b_{i})_{i\in I}\in V$ is such that there exists $k\in I$ with $b_{i}=e_{i}$ for all $i\in I\setminus\{k\}$, then $\mathbf{b}\in\mathcal{R}\mathbf{a}$. To prove it, pick $\ell\in I$ with $a_{\ell}\neq e_{\ell}$. By minimality of $\theta$ we have $(b_{k},e_{k})\in\operatorname{Cg}(a_{\ell},e_{\ell})$. Thus by Corollary 3.2, there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(a_{\ell})=b_{k}$ and $f(e_{\ell})=e_{k}$. Let $M=(r_{ij})_{i,j\in I}\in\mathcal{R}$ be the matrix given by $r_{k\ell}=f{\upharpoonright}_{C_{\ell}}$ and $r_{ij}=O_{ij}$ for $(i,j)\neq(k,\ell)$. Then $M\mathbf{a}=\mathsf{L}_{M}(\mathbf{a})=\mathbf{b}$ as required. It follows by Schur’s Lemma for modules that the endomorphism ring $\mathbf{End}({}_{\mathcal{R}}V)$ of the module $_{\mathcal{R}}V$ is a division ring $\mathbb{D}$. This is the division ring Freese associated to $\theta$ and the transversal $T$ (in the congruence modular case).
Proposition A1.1.
Suppose $\mathbf{A}$} is an algebra in a variety with a weak difference term and $\theta$ is an abelian minimal congruence of $\mathbf{A}$}. Let $\mathbb{D}$ be the division ring associated to $\theta$ and some transversal $T$ in the previous two paragraphs. Then $\mathbb{D}\cong\mathbb{F}_{\theta}$.
Proof sketch.
Let $\alpha=(0:\theta)$ and $\Delta=\Delta_{\theta,\alpha}$. Recall that the universe of $\mathbb{F}_{\theta}$ is the set of all endomorphisms $\lambda$ of the difference algebra $D({\mathbf{A}},\theta)={\mathbf{A}}(\theta)/\Delta$ which preserve $\varphi:=\overline{\alpha}/\Delta$ and fix the elements of $D^{o}=\{0_{E}\,:\,\text{$E\in A/\alpha$}\}$ pointwise. Let $V$ be the abelian group defined in the paragraphs preceding Proposition A1.1. Define the ring ${\mathbf{P}}=\prod_{i\in I}\mathbf{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i}))$ and let $\Psi:{\mathbf{P}}\rightarrow\mathbf{End}(V)$ be the obvious (coordinatewise) ring embedding. Given $\lambda\in\mathbb{F}_{\theta}$, define $\Phi(\lambda)=(\lambda_{i})_{i\in I}$ where for each $i\in I$, $\lambda_{i}:C_{i}\rightarrow C_{i}$ is the map given as follows: given $a\in C_{i}$, $\lambda_{i}(a)$ is the unique $b\in C_{i}$ such that $\lambda((a,e_{i})/\Delta)=(b,e_{i})/\Delta$ (see Corollary 6.7). It is tedious but straightforward to check that $\Phi$ is an injective ring homomorphism from $\mathbb{F}_{\theta}$ into $\mathbf{P}$}.
It remains only to show that the range of $\Psi\circ\Phi$ is precisely $\operatorname{End}({}_{\mathcal{R}}V)$, for then $\Psi\circ\Phi$ will be an isomorphism from $\mathbb{F}_{\theta}$ to $\mathbb{D}$. We can establish this by proving the following:
  1. (1)
    $\operatorname{End}({}_{\mathcal{R}}V)\subseteq\Psi(P)$.
  2. (2)
    For all $(\lambda_{i})_{i\in I}\in P$, the following are equivalent:
    1. (a)
      $\Psi((\lambda_{i})_{i\in I})\in\operatorname{End}({}_{\mathcal{R}}V)$.
    2. (b)
      For all $i,j\in I$ and all $r\in R_{ij}$, $\lambda_{i}\circ r=r\circ\lambda_{j}$.
    3. (c)
      $(\lambda_{i})_{i\in I}\in\Phi(\mathbb{F}_{\theta})$.
The first item can be deduced by applying the law $h\circ\mathsf{L}_{M}=\mathsf{L}_{M}\circ h$ for $h\in\operatorname{End}({}_{\mathcal{R}}V)$ and $M\in\mathcal{R}$, in particular for $M=(m_{ij})_{i,j\in I}$ where for some $k$ we have $m_{kk}=\mathrm{id}_{C_{k}}$ and $m_{ij}=O_{ij}$ for all $(i,j)\neq(k,k)$. A similar argument establishes the equivalence of (2a) and (2b).
It remains to prove (2b) $\Leftrightarrow$ (2c). First we record a helpful observation.
Claim 1. For all $i,j\in I$ and $r\in R_{ij}$, there exists a term $t(x,y_{1},\ldots,y_{n})$ and $\alpha$-classes $E_{1},\ldots,E_{n}$ such that for all $a\in C_{j}$,
$(r(a),e_{i})/\Delta=t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}}).$
To prove Claim 1, choose $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(e_{j})=e_{i}$ and $r=f{\upharpoonright}_{C_{j}}$, and pick a term $t(x,y_{1},\ldots,y_{n})$ and $\mathbf{u}\in A^{n}$ so that $f(x)=t(x,\mathbf{u})$. Then for each $\ell\in[n]$ let $E_{\ell}=u_{\ell}/\alpha$. Then for any $a\in C_{j}$,
$(r(a),e_{i})/\Delta$ $=(t(a,\mathbf{u}),t(e_{j},\mathbf{u}))/\Delta$
$=t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,(u_{1},u_{1})/\Delta,\ldots,(u_{n},u_{n})/\Delta)$
$=t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}}),$
which proves Claim 1.
Now assume that $(\lambda_{i})_{i\in I}=\Phi(\lambda)$ with $\lambda\in\mathbb{F}_{\theta}$. Fix $i,j\in I$ and $r\in R_{ij}$ and let $a\in C_{j}$. We must show $\lambda_{i}(r(a))=r(\lambda_{j}(a))$. Choose $t(x,\mathbf{y})$ and $E_{1},\ldots,E_{n}\in A/\alpha$ for $i,j,r$ as in Claim 1. By definition, $\lambda_{i}(r(a))$ is the unique $b\in C_{i}$ such that $\lambda((r(a),e_{i})/\Delta)=(b,e_{i})/\Delta$. Well,
$\lambda((r(a),e_{i})/\Delta)$ $=\lambda(t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}}))$
$=t^{D({\mathbf{A}},\theta)}(\lambda((a,e_{j})/\Delta),0_{E_{1}},\ldots,0_{E_{n}})$ as $\lambda\in\mathbb{F}_{\theta}$
$=t^{D({\mathbf{A}},\theta)}((\lambda_{j}(a),e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}})$
$=(r(\lambda_{j}(a)),e_{i})/\Delta,$
which implies $\lambda_{i}(r(a))=r(\lambda_{j}(a))$, proving (2c) $\Rightarrow$ (2b).
For the opposite implication, assume that $(\lambda_{i})_{i\in I}$ satisfies the condition of (2b). We first claim that for all $i,j\in I$, all $a\in C_{i}$, and all $b\in C_{j}$,
(8.1) $(a,e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e_{j})~~\mbox{implies}~~(\lambda_{i}(a),e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(\lambda_{j}(b),e_{j}).$
Indeed, assume $(a,e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e_{j})$. Choose a $\theta$-class $C_{k}$ so that $C_{i},C_{j}\rightarrow C_{k}$ (see Lemma 5.7). Let $f_{i},f_{j}\in\operatorname{Pol}_{1}({\mathbf{A}})$ be witnesses to $C_{i}\rightarrow C_{k}$ and $C_{j}\rightarrow C_{k}$ respectively. By Lemma 5.7, we can assume $f_{i},f_{j}$ are chosen so that $f_{\ell}(e_{\ell})=e_{k}$ for $\ell=i,j$. Then $(f_{i}(a),e_{k})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e_{j})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(f_{j}(b),e_{k})$ by Lemma 5.7(1) and the assumption, which implies $f_{i}(a)=f_{j}(b)$ by Lemma 5.2(1). Let $r=f_{i}{\upharpoonright}_{\theta_{i}}$ and $s=f_{j}{\upharpoonright}_{\theta_{j}}$ and note that $r\in R_{ki}$ and $s\in R_{kj}$. Condition (2b) then gives $\lambda_{k}\circ r=r\circ\lambda_{i}$ and $\lambda_{k}\circ s=s\circ\lambda_{j}$. Then
$(\lambda_{i}(a),e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r(\lambda_{i}(a)),e_{k})$ $=(\lambda_{k}(r(a)),e_{k})$
$=(\lambda_{k}(s(b)),e_{k})=(s(\lambda_{j}(b)),e_{k})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(\lambda_{j}(b),e_{j})$
which proves (8.1).
Now let $\lambda:D({\mathbf{A}},\theta)\rightarrow D({\mathbf{A}},\theta)$ be the unique map satisfying
$\lambda((x,e_{i})/\Delta)=(\lambda_{i}(x),e_{i})/\Delta$
for all $i\in I$ and all $x\in C_{i}$. ($\lambda$ is well-defined by (8.1).) We will show that $\lambda\in\mathbb{F}_{\theta}$, which will complete the proof of Proposition A1.1. Clearly $\lambda(\varphi)\subseteq\varphi$ and $\lambda(0_{E})=0_{E}$ for all $E\in A/\alpha$. It remains to show that $\lambda\in\operatorname{End}(D({\mathbf{A}},\theta))$. The next Claim is an immediate consequence of Proposition 3.8.
Claim 2. For all $n\geq 1$, all $f\in\operatorname{Pol}_{n}({\mathbf{A}})$, and all $\theta$-classes $C_{i_{1}},\ldots,C_{i_{n}}$, if $C_{\ell}$ is the $\theta$-class containing $f(e_{i_{1}},\ldots,e_{i_{n}})$, then there exist $r_{t}\in R_{\ell i_{t}}$ for $t\in[n]$ such that the following holds: for all $(a_{1},\ldots,a_{n})\in C_{i_{1}}\times\cdots\times C_{i_{n}}$,
[…]
Now let $F$ be an $n$-ary basic operation. Choose $n$ elements of $D({\mathbf{A}},\theta)$, which may be written $(a_{1},e_{i_{1}})/\Delta,\ldots,(a_{n},e_{i_{n}})/\Delta$ for some $\theta$-classes $\theta_{i_{t}}$ and $a_{t}\in C_{i_{t}}$. We must show
(8.2) $\lambda(F((a_{1},e_{i_{1}})/\Delta,\ldots,(a_{n},e_{i_{n}})/\Delta))=F(\lambda((a_{1},e_{i_{1}})/\Delta),\ldots,\lambda((a_{n},e_{i_{n}})/\Delta)).$
Let $C_{\ell}$ be the $\theta$-class containing $F(\mathbf{a})$ and choose $r_{t}\in R_{\ell i_{t}}$ for $t\in[n]$ as in Claim 2 for $F$ and $C_{i_{1}},\ldots,C_{i_{n}}$. We have $(F(\mathbf{a}),F(\mathbf{e}))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}),e_{\ell})$ by Lemma 5.3(1). Thus the left-hand side of (8.2) can be written
$\lambda(F((a_{1},e_{i_{1}})/\Delta,\ldots,(a_{n},e_{i_{n}})/\Delta))$ $=\lambda((F(\mathbf{a}),F(\mathbf{e}))/\Delta)$
$=\lambda((d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}))/\Delta)$
$=(\lambda_{\ell}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell})),e_{\ell})/\Delta.$
The right-hand side of (8.2) can be written
$F(\lambda((a_{1},e_{i_{1}})/\Delta),\ldots,\lambda((a_{n},e_{i_{n}})/\Delta))$ $=F((\lambda_{i_{1}}(a_{1}),e_{i_{1}})/\Delta,\ldots,(\lambda_{i_{n}}(a_{n}),e_{i_{n}})/\Delta)$
$=(F(\lambda_{i_{1}}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}))/\Delta$
$=(d(F(\lambda_{i-1}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}),e_{\ell}),e_{\ell})/\Delta.$
Thus to prove (8.2), it will suffice to show
(8.3) $\lambda_{\ell}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}))=d(F(\lambda_{i_{1}}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}),e_{\ell}).$
By Claim 2, there exist $r_{t}\in R_{\ell i_{t}}$ for $t\in[n]$ such that for all $(x_{1},\ldots,x_{n})\in C_{i_{1}}\times\cdots\times C_{i_{n}}$, $F(\mathbf{x})=\sum_{t=1}^{n}r_{t}(x_{t})+F(\mathbf{e})$. Recall that each $\lambda_{j}$ and $r_{t}$ is a group homomorphism between the appropriate groups from $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i_{1}}),\ldots,\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i_{n}}),\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{\ell})$. Thus computing in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{\ell})$,
$\lambda_{\ell}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}))$ $=\lambda_{\ell}(F(\mathbf{d}))-\lambda_{\ell}(F(\mathbf{e}))$
$=\lambda_{\ell}\left(\sum_{t=1}^{n}r_{t}(a_{t})+F(\mathbf{e})\right)-\lambda_{\ell}(F(\mathbf{e}))$
$=\sum_{t=1}^{n}\lambda_{\ell}(r_{t}(a_{t}))$ $(\dagger)$
while
$d(F(\lambda_{i_{1}}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}),e_{\ell})$ $=\left(\sum_{t=1}^{n}r_{t}(\lambda_{i_{t}}(a_{t}))+F(\mathbf{e})\right)-F(\mathbf{e})$
$=\sum_{t=1}^{n}r_{t}(\lambda_{i_{t}}(a_{t}).$ $(\ddagger)$
Since $\lambda_{\ell}\circ r_{t}=r_{t}\circ\lambda_{i_{t}}$ for each $t\in[n]$ by assumption (2b), we have established (8.3) and so have finished the proof of Proposition A1.1. ∎
Finally, we show that $\mathbb{F}_{\theta}$ coincides with the finite field $\mathbb{F}$ constructed in the proof of Theorem 3.9 when $\mathbf{A}$} is finite.
Proposition A1.2.
Suppose $\mathcal{V}$ is a variety with a weak difference term $d$, $\mathbf{A}$} is a finite algebra in $\mathcal{V}$, and $\theta$ is an abelian minimal congruence of $\mathbf{A}$}. Let $\mathbb{F}$ be the finite field associated to $\theta$ in the proof of Theorem 3.9. That is, let $C_{1},\ldots,C_{k}$ be a list of the $\theta$-classes, let $\overline{C}:=C_{1}\times\cdots\times C_{k}$, choose $\mathbf{e}=(e_{1},\ldots,e_{k})\in\overline{C}$, and define ++ on $\overline{C}$ by $x+y:=d(x,\mathbf{e},y)$ so that ${\mathbf{C}}:=(\overline{C},+,\mathbf{e})$ is an abelian group. Let $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ be defined as in the proof of Theorem 3.9, let $\mathbf{R}$} be the subring of $\operatorname{End}({\mathbf{C}})$ such that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is polynomially equivalent to $_{{\mathbf{R}}}\hskip-1.0pt{\mathbf{C}}$, and let $\mathbb{F}=\operatorname{End}({}_{{\mathbf{R}}}\hskip-1.0pt{\mathbf{C}})$. Then $\mathbb{F}\cong\mathbb{F}_{\theta}$.
Proof sketch.
First recall the construction of Freese’s division ring $\mathbb{D}$ associated to $\theta$ and the transversal $T:=\{e_{1},\ldots,e_{k}\}$ described at the beginning of this section. By finiteness, the abelian group $V$ in Freese’s construction coincides with the group $\mathbf{C}$}. Also recall the unital subring $\mathsf{L}_{\mathcal{R}}$ of $\mathbf{End}({\mathbf{C}})$ from the construction of $\mathbb{D}$. If we can prove $\mathsf{L}_{\mathcal{R}}={\mathbf{R}}$, then it will follow that $\mathbb{D}=\mathbb{F}$ and therefore $\mathbb{F}\cong\mathbb{F}_{\theta}$ by Proposition A1.1.
Following [8], the universe of $\mathbf{R}$} may be given by
$R=\{f(x)\in\operatorname{Pol}_{1}(\mathfrak{C}({\mathbf{A}},\nu_{\theta}))\,:\,\text{$f(\mathbf{e})=\mathbf{e}$}\}.$
By [23, Remark 2.8], the unary polynomials of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ are the functions $f:C\rightarrow C$ which can be described as follows: for some $k$-ary polynomials $f_{1},\ldots,f_{k}\in\operatorname{Pol}_{k}({\mathbf{A}})$ with $f_{i}(\mathbf{e})\in C_{i}$ for $i\in[k]$,
[…]
Assuming $f\in R$, i.e., $f(\mathbf{e})=\mathbf{e}$, we get $f_{i}(\mathbf{e})=e_{i}$ for each $i\in[k]$. Thus for each $i\in[k]$ there exist $r_{1},\ldots,r_{k}\in R_{ij}$ (see Freese’s construction) so that for all $\mathbf{c}=(c_{1},\ldots,c_{k})\in C$,
$f_{i}(\mathbf{c})=\sum_{j=1}^{k}r_{ij}(c_{j})\quad\mbox{computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$},$
by Proposition 3.8. That is, $f=\mathsf{L}_{M}$ where $M=(r_{ij})_{i,j\in[k]}\in\mathcal{R}$ (assuming $f\in R$). The logic is reversible and we have $R=\mathsf{L}_{\mathcal{R}}$ as desired. ∎
Appendix 2: Examples
In this appendix we illustrate some of our results from Sections 5 and 6 by presenting a recipe for constructing certain finite subdirectly irreducible algebras with abelian monoliths in varieties having weak difference terms. The algebras $\mathbf{A}$} can be constructed so that if $\mu$ is the monolith, $\alpha=(0:\mu)$, and $\varphi=\overline{\alpha}/\Delta_{\mu,\alpha}$, then
  1. (1)
    $\mathbb{F}_{\mu}$ can be any desired finite field $\mathbb{F}$;
  2. (2)
    $|A/\alpha|$ can be any positive integer.
  3. (3)
    The vector spaces $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{D({\mathbf{A}})}(\varphi,0_{E})$ as $E$ ranges over the $\alpha$-classes can have any desired (finite) dimensions.
  4. (4)
    The ranges of the $\mu$-classes within a given $\alpha$-class $E$ can be any collection of subspaces of $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{D({\mathbf{A}})}(\varphi,0_{E})$, as long as the collection includes $0_{E}/\varphi$ itself.
Our recipe is a variant of a known construction [3]. We first give a useful method for constructing varieties with a prescribed weak difference term. For this, we need the following result characterizing weak difference terms at the level of varieties. Following [18], given a variety $\mathcal{V}$, let ${\mathbf{F}}={\mathbf{F}}_{\mathcal{V}}(3)$ be the $\mathcal{V}$-free algebra freely generated by $\{x,y,z\}$, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ calculated in $\mathbf{F}$}, and define $\beta_{0}=\beta$, $\gamma_{0}=\gamma$, $\beta_{n+1}=\beta\vee(\alpha\wedge\gamma_{n})$, and $\gamma_{n+1}=\gamma\vee(\alpha\wedge\beta_{n})$ for $n\geq 0$.
Theorem A2.1.
There exists an infinite sequence $\Sigma_{0},\Sigma_{1},\Sigma_{2},\ldots$ of finite sets of identities involving variables $x,y$ and ternary operation symbols $\mathsf{d},\mathsf{f}_{1},\mathsf{f}_{2},\ldots$, such that for every variety $\mathcal{V}$ and every ternary term $d(x,y,z)$ in the signature of $\mathcal{V}$, the following are equivalent:
  1. (1)
    $d$ is a weak difference term for $\mathcal{V}$.
  2. (2)
    For some $n\geq 0$ there exist idempotent ternary terms $f_{1},f_{2},\ldots$ in the signature of $\mathcal{V}$ so that, when the symbols $\mathsf{d},\mathsf{f}_{1},\mathsf{f}_{2},\ldots$ are interpreted by $d,f_{1},f_{2},\ldots$ respectively, the identities in $\Sigma_{n}$ are true in $\mathcal{V}$.
  3. (3)
    Letting ${\mathbf{F}}={\mathbf{F}}_{\mathcal{V}}(3)$ and interpreting $d$ as an element of $\mathbf{F}$} in the usual way, i.e. as $d^{{\mathbf{F}}}(x,y,z)$, we have $(x,d)\in\gamma_{m}$ and $(d,z)\in\beta_{m}$ for some $m\geq 0$.
Proof sketch.
(1) $\Rightarrow$ (3) can be extracted from the proof of Theorem 4.8 (1) $\Rightarrow$ (2) $\Rightarrow$ (3) in [18]. Standard arguments with roots in [27, 32] can produce sets $\Sigma_{0},\Sigma_{1},\Sigma_{2},\ldots$ of identities witnessing (2) $\Leftrightarrow$ (3). To prove (3) $\Rightarrow$ (1), assume that $(x,d)\in\gamma_{m}$ and $(d,z)\in\beta_{m}$. Then by [33, Theorem 2.1 (3) $\Rightarrow$ (4)], there exist idempotent ternary terms $f_{1},\ldots,f_{n},g_{1},\ldots,g_{n}$ (for some $n$) such that $\mathcal{V}\models f_{i}(x,y,x)\approx g_{i}(x,y,x)$ for each $i\in[n]$, and for all ${\mathbf{A}}\in\mathcal{V}$ and all $a,b\in A$, if
[…]
then $d(a,a,b)=b=d(b,a,a)$ and $d(b,b,a)=a=d(a,b,b)$. Now one can easily adapt the proof of [13, Theorem 1.2 (3) $\Rightarrow$ (1)] to get that $d$ is a weak difference term for $\mathcal{V}$. ∎
Definition A2.2.
A semilattice-over-Maltsev operation is any operation $d:A^{3}\rightarrow A$ constructed in the following way. Let $(S,\wedge)$ be a (meet) semilattice. For each $s\in S$ let $V_{s}$ be a non-empty set, and assume that $V_{s}\cap V_{t}=\varnothing$ for all $s,t\in S$ with $s\neq t$. For each pair $s,t\in S$ with $s\geq t$, assign a function $f_{(s,t)}:V_{s}\rightarrow V_{t}$, so that $f_{(s,s)}=\mathrm{id}_{V_{s}}$. Finally, for each $s\in S$ let $m_{s}:(V_{s})^{3}\rightarrow V_{s}$ be a Maltsev operation.
From this data, define $A=\bigcup_{s\in S}V_{s}$ and $\nu:A\rightarrow S$ so that $a\in V_{\nu(a)}$ for all $a\in A$. Then define $d:A^{3}\rightarrow A$ by
$d(a,b,c)=m_{t}(f_{(\nu(a),t)}(a),f_{(\nu(b),t)}(b),f_{(\nu(c),t)}(c))\quad\mbox{where $t=\nu(a)\wedge\nu(b)\wedge\nu(c)$}.$
Lemma A2.3.
Suppose $\mathbf{A}$} is an algebra and $d$ is a ternary term operation of $\mathbf{A}$} which is semilattice-over-Maltsev. Then $d$ is a weak difference term for $\mathsf{HSP}({\mathbf{A}})$.
Proof.
Clearly $d$ is idempotent, and it can be easily checked that $\mathbf{A}$} satisfies the following identities:
(8.4) $d(d(x,y,y),x,x)$ $\approx d(x,y,y)\approx d(d(x,y,y),y,y)$
$d(x,x,d(x,x,y))$ $\approx d(x,x,y)\approx d(y,y,d(x,x,y)).$
Define ternary terms $f_{i},g_{i}$ for $i=0,1,2,3$ by
$f_{0}(x,y,z)$ $=x$ $g_{0}(x,y,z)$ $=d(x,z,z)$
$f_{1}(x,y,z)$ $=d(x,y,y)$ $g_{1}(x,y,z)$ $=d(d(x,y,y),z,z)$
$f_{2}(x,y,z)$ $=z$ $g_{2}(x,y,z)$ $=d(x,x,z)$
$f_{3}(x,y,z)$ $=d(y,y,z)$ $g_{3}(x,y,z)$ $=d(x,x,d(y,y,z)).$
From idempotency of $d$ and the identities (8.4), one can deduce that $\mathbf{A}$} satisfies the identities
(8.5) $f_{i}(x,y,x)\approx g_{i}(x,y,x)\quad\mbox{for all $i$}$
and
(8.6) $x$ $\approx f_{0}(x,y,y)$ $f_{2}(x,x,y)$ $\approx y$
$f_{1}(x,x,y)$ $\approx f_{0}(x,x,y)$ $f_{2}(x,y,y)$ $\approx f_{3}(x,y,y)$
$f_{1}(x,y,y)$ $\approx g_{1}(x,y,y)$ $g_{3}(x,x,y)$ $\approx f_{3}(x,x,y)$
$g_{0}(x,x,y)$ $\approx g_{1}(x,x,y)$ $g_{3}(x,y,y)$ $\approx g_{2}(x,y,y)$
$g_{0}(x,y,y)$ $\approx d(x,y,y)$ $d(x,x,y)$ $\approx g_{2}(x,x,y).$
The identities (8.5) and (8.6) therefore hold in $\mathsf{HSP}({\mathbf{A}})$. Let ${\mathbf{F}}={\mathbf{F}}_{\mathsf{HSP}({\mathbf{A}})}(3)$ be the free algebra in $\mathsf{HSP}({\mathbf{A}})$ freely generated by $\{x,y,z\}$. As in the discussion preceding Theorem A2.1, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ be calculated in $\mathbf{F}$}. Interpreting $d$ and $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ as elements of $\mathbf{F}$} in the usual way, we find that the identities (8.5) and (8.6) insure that $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ witness
$(x,d)$ $\in\gamma\circ(\alpha\cap(\beta\circ(\alpha\cap\gamma)\circ\beta))\circ\gamma\subseteq\gamma\vee(\alpha\wedge(\beta\vee(\alpha\wedge\gamma)))=\gamma_{2}$
$(d,z)$ $\in\beta\circ(\alpha\cap(\gamma\circ(\alpha\cap\beta)\circ\gamma))\circ\beta\subseteq\beta\vee(\alpha\wedge(\gamma\vee(\alpha\wedge\beta)))=\beta_{2}.$
Thus $d$ is a weak difference term for $\mathsf{HSP}({\mathbf{A}})$ by Theorem A2.1. ∎
Here is our recipe for constructing interesting finite subdirectly irreducible algebras with abelian monoliths.
Example A2.4.
Fix a finite field $\mathbb{F}$. Let $V_{(0,0)},V_{(1,0)},\ldots,V_{(m,0)}$ be finite-dimensional vector spaces over $\mathbb{F}$ with $m\geq 0$ and $\dim(V_{(0,0)})>0$, chosen so that $V_{(0,0)},\ldots,V_{(m,0)}$ are pairwise disjoint. For each $\ell\leq m$ let $W_{(\ell,0)},W_{(\ell,1)},\ldots,W_{(\ell,n_{\ell})}$ be a list of (not necessarily distinct) subspaces of $V_{(\ell,0)}$ with $W_{(\ell,0)}=V_{(\ell,0)}$.
Choose vector spaces $V_{(\ell,i)}$ over $\mathbb{F}$ and isomorphisms $\sigma_{(\ell,i)}:V_{(\ell,i)}\cong W_{(\ell,i)}$ for $\ell\leq m$ and $1\leq i\leq n_{\ell}$, so that the spaces $V_{(\ell,i)}$ ($0\leq\ell\leq m$, $0\leq i\leq n_{\ell}$) are pairwise disjoint. Also set $\sigma_{(\ell,0)}=\mathrm{id}_{V_{(\ell,0)}}$ for each $\ell\leq m$. Let $E_{\ell}$ be the (disjoint) union of $V_{(\ell,0)},V_{(\ell,1)},\ldots,V_{(\ell,n_{\ell})}$, and let $A$ be the (disjoint) union of $E_{0},\ldots,E_{m}$. Then let $\mu$ and $\alpha$ be the equivalence relations on $A$ whose equivalence classes are the spaces $V_{(\ell,i)}$ ($\ell\leq m$, $i\leq n_{\ell}$) and the sets $E_{0},\ldots,E_{m}$ respectively. Also let $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$ and let $\psi$ be the equivalence relation on $B$ whose equivalence classes are $V_{(0,0)},\ldots,V_{(m,0)}$. We will build an algebra $\mathbf{A}$} with universe $A$ such that:
  • $\mathbf{A}$} belongs to a variety with a weak difference term.
  • $\mathbf{A}$} is subdirectly irreducible with abelian monolith $\mu$. In fact, $\mu$ will be the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
  • $\alpha=(0:\mu)$.
  • Letting $\Delta=\Delta_{\mu,\alpha}$ and $\varphi=\overline{\alpha}/\Delta$, there is a bijection from $D({\mathbf{A}},\mu)$ to $B$ which sends $\varphi$ to $\psi$ and sends each $\varphi$-class $(\mu\cap(E_{\ell})^{2})/\Delta$ to $V_{(\ell,0)}$.
  • Modulo this bijection, for each $\alpha$-class $E_{\ell}$, the ranges (see Definition 5.4) of the $\mu$-classes $V_{(\ell,i)}$ contained in $E_{\ell}$ correspond precisely to the subspaces $W_{(\ell,i)}$ in $V_{(\ell,0)}$.
  • The division ring $\mathbb{F}_{\mu}$ associated to $\mu$ is $\mathbb{F}$ (up to isomorphism).
Begin by letting $S=\{(\ell,i)\,:\,\text{$\ell\leq m$ and $i\leq n_{\ell}$}\}$. Thus $S$ is the set of indices for the vector spaces $V_{(\ell,i)}$ which constitute $A$. Define a binary operation $\wedge$ on $S$ by
[…]
Then $(S,\wedge)$ is the meet semilattice with least element $(0,0)$ pictured in Figure 2.
[…]
Figure 2. The semilattice $(S,\wedge)$
Recall that for each $s=(\ell,i)\in S$ we have a corresponding vector space $V_{s}$ over $\mathbb{F}$. To each pair of elements $s,t\in S$ with $s\geq t$ in the semilattice ordering, we associate an $\mathbb{F}$-linear map $f_{(s,t)}:V_{s}\rightarrow V_{t}$ as follows.
  1. (1)
    $f_{(s,s)}=\mathrm{id}_{V_{s}}$ for all $s\in S$.
  2. (2)
    $f_{((\ell,i),(\ell,0))}=\sigma_{(\ell,i)}$ for $(\ell,i)\in S$ with $i\neq 0$.
  3. (3)
    $f_{((\ell,i),(0,0))}=$ the constant zero map, for $(\ell,i)\in S$ with $\ell\neq 0$.
Now let $d:A^{3}\rightarrow A$ be the semilattice-over-Maltsev operation (see Definition A2.2) defined by the data $(S,\wedge)$, $(V_{s}:s\in S)$, $(f_{(s,t)}:s,t\in S,\,s\geq t)$, and the family of Maltsev operations $m_{s}(x,y,z)=x-y+z$ evaluated in each vector space $V_{s}$. That is, for all $a\in V_{s_{1}}$, $b\in V_{s_{2}}$ and $c\in V_{s_{3}}$, letting $t=s_{1}\wedge s_{2}\wedge s_{3}$ we have
$\mbox{$d(a,b,c)=f_{(s_{1},t)}(a)-f_{(s_{2},t)}(b)+f_{(s_{3},t)}(c)$\quad calculated in $V_{t}$}.$
For each $\ell\leq m$ define $\sigma_{\ell}:E_{\ell}\rightarrow V_{(\ell,0)}$ so that $\sigma_{\ell}{\upharpoonright}_{V_{(\ell,i)}}=\sigma_{(\ell,i)}$ for all $i\leq n_{\ell}$. For each $s\in S$ let $0_{s}$ denote the zero element of $V_{s}$. We define some unary operations on $A$.
  1. (1)
    For each $(\ell,i)\in S$ and each nonconstant $\mathbb{F}$-linear map $g:V_{(\ell,0)}\rightarrow V_{(\ell,i)}$, define $F_{(\ell,i)}^{g}:A\rightarrow A$ by
    $F_{(\ell,i)}^{g}(a)=\left\{\begin{array}[]{cl}(g\circ\sigma_{\ell})(a)&\mbox{if $x\in E_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in E_{k}$ with $k\neq\ell$.}\end{array}\right$.
  2. (2)
    For each $\ell\in\{1,\ldots,m\}$ with $\dim(V_{(\ell,0)})>0$, choose nonconstant $\mathbb{F}$-linear maps $g_{\ell}:V_{(0,0)}\rightarrow V_{(\ell,0)}$ and $h_{\ell}:V_{(\ell,0)}\rightarrow V_{(0,0)}$ and define $G_{\ell},G^{\prime}_{\ell}:A\rightarrow A$ by
    $G_{\ell}(a)=\left\{\begin{array}[]{cl}(g_{\ell}\circ\sigma_{0})(a)&\mbox{if $a\in E_{0}$}\\ 0_{(k,0)}&\mbox{if $a\in E_{k}$ with $k\neq 0$}\end{array}\right$.
    and
    $G_{\ell}^{\prime}(a)=\left\{\begin{array}[]{cl}(h_{\ell}\circ\sigma_{\ell})(a)&\mbox{if $a\in E_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in E_{k}$ with $k\neq\ell$}\end{array}\right$.
We also need the following binary operations.
  1. (3)
    Fix $1\in V_{(0,0)}$ with $1\neq 0_{(0,0)}$. For every $s\in S$, define $H_{s}:A^{2}\rightarrow A$ by
    $H_{s}(a,b)=\left\{\begin{array}[]{cl}1&\mbox{if $a,b\in V_{s}$}\\ 0_{(0,0)}&\mbox{otherwise.}\end{array}\right$.
  2. (4)
    For each $\ell\in\{1,\ldots,m\}$, define $K_{\ell}:A^{2}\rightarrow A$ by
    $K_{\ell}(a,b)=\left\{\begin{array}[]{cl}b&\mbox{if $a\in E_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\not\in E_{\ell}$ and $b\in E_{k}$}.\end{array}\right$.
Let $\mathscr{F}$ be the set of all the operations $F_{(\ell,i)}^{g}$, let $\mathscr{G}$ be the set of all the operations $G_{\ell}$ and $G_{\ell}^{\prime}$, let $\mathscr{H}=\{H_{s}\,:\,\text{$s\in S$}\}$, and let $\mathscr{K}=\{K_{\ell}\,:\,\text{$1\leq\ell\leq m$}\}$. Finally, let
${\mathbf{A}}=(A,\{d\}\cup\mathscr{F}\cup\mathscr{G}\cup\mathscr{H}\cup\mathscr{K}).$
Claim A2.4.
  1. (1)
    $d$ is a weak difference term $\mathsf{HSP}({\mathbf{A}})$.
  2. (2)
    $\mathbf{A}$} is subdirectly irreducible with abelian monolith $\mu$. Moreover, $\mu$ is the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
  3. (3)
    $(0:\mu)=\alpha$.
  4. (4)
    $\Delta_{\mu,\alpha}$ is the set of all pairs $((a,b),(a^{\prime},b^{\prime}))$ such that for some $\ell\leq m$ and some $i,j\leq n_{\ell}$, $\{a,b\}\subseteq V_{(\ell,i)}$ while $\{a^{\prime},b^{\prime}\}\subseteq V_{(\ell,j)}$ and
    $\sigma_{(\ell,i)}(a)-\sigma_{(\ell,i)}(b)=\sigma_{(\ell,j)}(a^{\prime})-\sigma_{(\ell,j)}(b^{\prime})$ calculated in $V_{(\ell,0)}$.
Proof.
(1) follows from Lemma A2.3. It is easy to check that $\mu$ and $\alpha$ are congruences of $\mathbf{A}$}. Let $\theta\in\operatorname{Con}({\mathbf{A}})$ satisfy $0\neq\theta\subseteq\mu$. (We want to show $\theta=\mu$.) We can choose $(\ell,i)\in S$ so that $\theta\cap(V_{(\ell,i)})^{2}$ contains a pair $(c,c^{\prime})$ with $c\neq c^{\prime}$. Let $a=c-c^{\prime}$ calculated in $V_{(\ell,i)}$. Then $a\neq 0_{(\ell,i)}$ and $(a,0_{(\ell,i)})\in\theta$ because of $d(x,y,z)$. Now let $j\leq n_{\ell}$ and $b\in V_{(\ell,j)}$ with $b\neq 0_{(\ell,j)}$. Let $a^{\prime}=\sigma_{(\ell,i)}(a)$. Then $a^{\prime}\neq 0_{(\ell,0)}$. Hence there exists an $\mathbb{F}$-linear map $g:V_{(\ell,0)}\rightarrow V_{(n,j)}$ with $g(a^{\prime})=b$. Then $F_{(\ell,j)}^{g}(a)=b$ while $F_{(\ell,j)}^{g}(0_{(\ell,i)})=0_{(\ell,j)}$, proving $(b,0_{(\ell,j)})\in\theta$. As $j$ and $b$ were arbitrary, this proves $\mu\cap(E_{\ell})^{2}\subseteq\theta$.
If $\ell>0$, we can use the operation $G_{\ell}^{\prime}$ to show that $\theta\cap(V_{(0,0)})^{2}$ contains a nontrivial pair. Then the argument in the previous paragraph gives $\mu\cap(E_{0})^{2}\subseteq\theta$. Now let $k\in\{1,\ldots,m\}$ be such that some $\mu$-class $V_{(k,j)}$ contained in $E_{k}$ has more than one element. Then we can use the operation $G_{k}$ to show that $\theta\cap(V_{(k,0)})^{2}$ contains a nontrivial pair. Again by the argument in the previous paragraph, we get $\mu\cap(E_{k})^{2}\subseteq\theta$. This finishes the proof that $\theta=\mu$; hence $0_{A}\prec\mu$.
Next let $(c,c^{\prime})$ be any nontrivial pair in $A^{2}$ with $(c,c^{\prime})\not\in\mu$. Say $c\in V_{s}$ and $c^{\prime}\in V_{t}$ with $s\neq t$. Then the polynomial $f(x)=H_{s}(x,c)$ sends $(c,c^{\prime})$ to $(1,0_{(0,0)})$, so $\operatorname{Cg}(c,c^{\prime})\cap\mu\neq 0$. As we have shown $0_{A}\prec\mu$, this proves $\mu$ is the monolith of $\mathbf{A}$}. Combining the above argument with Corollary 3.2, we also get that for all $(a,b)\in\mu$ and all $(c,c^{\prime})\in A^{2}$ with $c\neq c^{\prime}$, there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(c)=a$ and $f(c^{\prime})=b$. This proves that $\mu$ is the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
Next we will show that $C(\alpha,\mu;0)$ holds. Define $R$ to be the set of all matrices $\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in A^{2\times 2}$ such that for some $\ell\leq m$ and some $i,j\leq n_{\ell}$,
  • $0\in\{i,j\}$ or $i=j$.
  • $a,b\in V_{(\ell,i)}$ and $a^{\prime},b^{\prime}\in V_{(\ell,j)}$.
  • $\sigma_{\ell}(a)-\sigma_{\ell}(b)=\sigma_{\ell}(a^{\prime})-\sigma_{\ell}(b^{\prime})$ calculated in $V_{(\ell,0)}$.
One can show that $R$ is a subuniverse of ${\mathbf{A}}^{2\times 2}$ (exercise). Clearly $X(\mu,\alpha)\subseteq R$ (see Definition 2.1). Conversely, suppose
$\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in R\quad\mbox{with $a,b\in V_{(\ell,i)}$ and $a^{\prime},b^{\prime}\in V_{(\ell,j)}$}.$
We can assume with no loss of generality that $j=0$ or $i=j$. Then
$\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}=d\left(\begin{pmatrix}a&a^{\prime}\\ a&a^{\prime}\end{pmatrix},\begin{pmatrix}a&a\\ a&a\end{pmatrix},\begin{pmatrix}a&a\\ b&b\end{pmatrix}\right)$,
proving $R\subseteq\operatorname{Sg}^{{\mathbf{A}}^{2\times 2}}(X(\mu,\alpha))$. These remarks prove $R=M(\mu,\alpha)$. Since every matrix in $R$ has the property that the first column is in $0_{A}$ if and only if the second column is in $0_{A}$, we have proved that $C(\alpha,\mu;0)$ holds and therefore $\mu$ is abelian and $(0:\mu)\geq\alpha$. To prove $(0:\mu)=\alpha$, suppose $\theta\in\operatorname{Con}({\mathbf{A}})$ and $\theta>\alpha$. Pick $(a,b)\in\theta\setminus\alpha$, say with $a\in E_{\ell}$ and $b\in E_{k}$ with $k\neq\ell$. Using the operation $K_{\ell}$ and the pair $(0_{(0,0)},1)\in\mu\setminus 0_{A}$, we have
$K_{\ell}(b,0_{(0,0)})=0_{(0,0)}=K_{\ell}(b,1)$
but
$K_{\ell}(a,0_{(0,0)})=0_{(0,0)}\neq 1=K_{\ell}(a,1),$
proving that $C(\theta,\mu;0)$ fails for every $\theta>\alpha$. Hence $(0:\mu)=\alpha$.
Recall from Lemma 4.4(1) that $\Delta_{\mathsf{h}}(\mu,\alpha)$ is the horizontal transitive closure of $M(\mu,\alpha)$. Thus $\Delta_{\mu,\alpha}$ is the transitive closure of
$\left\{((a,b),(a^{\prime},b^{\prime})):\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in R\right\}.$
The description of $\Delta_{\mu,\alpha}$ in (4) then follows. ∎
Recall that $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$. Let $\Delta=\Delta_{\mu,\alpha}$, so $D({\mathbf{A}},\mu)={\mathbf{A}}(\mu)/\Delta$. Observe that every element of $D({\mathbf{A}},\mu)$ can be uniquely written in the form $(a,0_{(\ell,0)})/\Delta$ for some $\ell\leq m$ and $a\in V_{(\ell,0)}$. Let $h:D({\mathbf{A}},\mu)\rightarrow B$ be the bijection sending $(a,0_{(\ell,0)})/\Delta\mapsto a$. Then by construction, for every $(\ell,i)\in S$ we have
$\operatorname{ran}(V_{(\ell,i)})$ $=\operatorname{ran}(\chi_{0_{(\ell,i)}})$ Definition 5.4
$=\{(a,0_{(\ell,i)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$
$=\{(\sigma_{(\ell,i)}(a),0_{(\ell,0)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$ Claim A2.4(4)
$=\{(a^{\prime},0_{(\ell,0)})/\Delta\,:\,\text{$a^{\prime}\in W_{(\ell,i)}$}\}$ as $\sigma_{(\ell,i)}(V_{(\ell,i)})=W_{(\ell,i)}$.
Hence $h(\operatorname{ran}(V_{(\ell,i)}))=W_{(\ell,i)}$ as desired.
Finally, let $T$ be the transversal for $\mu$ given by $T=\{0_{s}\,:\,\text{$s\in S$}\}$, and let $\mathbb{D}$ be the division ring associated to $\mu,T$ by Freese’s construction as described in Appendix 1. We will verify that $\mathbb{F}\cong\mathbb{D}$, which will imply $\mathbb{F}\cong\mathbb{F}_{\mu}$ by Proposition A1.1.
As in the proof of Proposition A1.1, let $V$ be the group $\prod_{s\in S}\operatorname{Grp}_{{\mathbf{A}}}(\mu,0_{s})$, let $\mathbf{P}$} be the ring $\prod_{s\in S}\mathbf{End}(\operatorname{Grp}_{{\mathbf{A}}}(\mu,0_{s}))$, let $\Psi:{\mathbf{P}}\rightarrow\mathbf{End}(V)$ be the obvious (coordinatewise) ring embedding, and recall that $\mathbb{D}$ is the subring of $\mathbf{End}(V)$ whose universe is
$\{\Psi((\lambda_{s})_{s\in S})\,:\,\text{$\lambda_{s}\circ f{\upharpoonright}_{V_{t}}=f\circ\lambda_{t}$ for all $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(0_{t})=0_{s}$}\}.$
For each $\zeta\in\mathbb{F}$ and $s\in S$ let $\zeta_{s}:V_{s}\rightarrow V_{s}$ be the “scalar multiplication by $\zeta$ in $V_{s}$” map and let $\Theta(\zeta)=(\zeta_{s})_{s\in S}$. Clearly $(\zeta_{s})_{s\in S}\in P$. Every $n$-ary basic operation of $\mathbf{A}$}, when restricted to an $n$-tuple of $\mu$-classes, is $\mathbb{F}$-affine, and this property is inherited by polynomial operations. It follows that $(\Psi\circ\Theta)(\zeta)\in\mathbb{D}$. It is not hard to check that $\Psi\circ\Theta$ is a ring embedding of $\mathbb{F}$ into $\mathbb{D}$.
To prove $\operatorname{ran}(\Psi\circ\Theta)=\mathbb{D}$, fix $\lambda\in\mathbb{D}$ and choose $(\lambda_{s})_{s\in S}\in P$ with $\Psi((\lambda_{s})_{s\in S})=\lambda$. Thus
($\ast$) $\lambda_{s}\circ f{\upharpoonright}_{V_{t}}=f\circ\lambda_{t}$ for all $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $s,t\in S$ with $f(0_{t})=0_{s}$.
[…]
$\zeta^{t}\cdot b=\zeta^{t}\cdot g_{\ell}(a)$ $=g_{\ell}(\zeta^{t}\cdot a)$ $\mathbb{F}$-linearity of $g_{\ell}$
$=g_{\ell}(\lambda_{t}(a))$
[…]
$=\zeta^{s}\cdot b.$
Since $b\neq 0_{s}$, we get $\zeta^{s}=\zeta^{t}$. A similar argument using operations in $\mathscr{F}$ gives $\zeta^{(\ell,i)}=\zeta^{(\ell,0)}$ whenever $\ell\leq m$ and $\dim(V_{(\ell,i)})>0$. This completes the proof that $\mathbb{F}_{u}\cong\mathbb{F}$.